A Complete Bibliography of Publications in IEEE Computer Architecture Letters

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 April 2024 Version 1.09

Title word cross-reference

3 [RMMLK16, ZBA^{+}20]. O(1) [LX08].

-D [RMMLK16].

128-Bit [DPP23].

2.0 [PZX15]. 2.5D [CWK^{+}22].

3D [HRF^{+}11, HLR21, XYMY16].

4T [JDK^{+}02].

Abstract [BEA^{+}13]. accel [VRFT24]. Accelerate [JLA16, NK22]. Accelerated [FFAMK15, JLKK23, KLR24, LZL^{+}20].

Accelerating [CPK^{+}23, KKJ^{+}22, KHS^{+}24, SK21, SPHS22, SAA^{+}23, VRS18, WMZY17, ZZW^{+}22]. Acceleration [GKK^{+}22, HCK^{+}21, JLRA18, KKL^{+}15, LYY^{+}21, LYL^{+}16, LHZ19, WKE12, YCH24]. Accelerator [BHY^{+}19, CMP^{+}14, DXSS15, GMPMC^{+}23, GIH^{+}24, KDS22, LAC14, LBB^{+}19, LWM20, LAM^{+}22, LHWB10, LWB13, MMY^{+}14, NBH13, RSRT19, VRFT24, XHG^{+}19, YHM17, YG18, ZL18a]. accelerator-based [LHWB10]. Accelerator-Rich [LBB^{+}19]. Accelerators [AW15, BSMB23, FSO^{+}22, JKK^{+}21, KPPK21, MNFI20, MMAAK21, OSH16, PGC22, RSO21, WLDN19, YSL^{+}21]. Access [Ano13h, Ano13i, CYAW20, DSVK12, KSB19, LGLK17, MSI18, PJ22, SCR^{+}17, WTSW21, XYMY16]. Access-Control [LGLK17]. accesses [Zha06]. Accounting [LJM^{+}14, LMC^{+}09]. Accumulate [GG17, JPC18]. Accuracy [DKD07, SHK15]. Accurate [BREM08, CAPS09, CVF^{+}24, JC17, KHB^{+}19, LYR^{+}20, RCBJ11, SJ22]. ACE [BREM08]. Achieve [WZLQ15]. Achieving [NB24, SCR^{+}17, WCZ^{+}12]. Across [WXZ^{+}21]. Active [BDJ06]. Ad [Ano09a]. Adapting [MNFI20]. Adaptive [GF16, KK21, LZLX15, MCKW10, SK21, SDTG04, SCF04, XYZ15, XYMY16, YKP^{+}22, ZWL15]. Address [IKW^{+}20, KNGK15, KJS^{+}19, MKP^{+}24, SfCL03, VD02, YWG17, AD06, LLLM06]. Addressable [VHN15, Yav24]. Addressing [ÇE14, MVJ17]. ADL [BVL09]. Adopting [LLLM06]. ADT [MDK^{+}23]. Advance [KMJ18]. Advanced [Ano16k]. Advancing [RAD^{+}23]. Advertisement [Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano10c, Ano10d, Ano10b, Ano10f, Ano10e, Ano12c, Ano12k, Ano10g, Ano14f]. Affinity [HLH16]. Against [LEBM20, ZNTJE23, OKS^{+}15, SKS^{+}15]. Aggressive [JTG23, MDK^{+}23]. Aging [SRH20]. Aging-Aware [SRH20]. AI [MGH^{+}22]. Algebraic [GMPMC^{+}23]. Algorithm [LX08, LAS22, XL07, YCH24]. Algorithm/Hardware [LAS22]. Algorithms [CLCG14]. AligneR [ZZJ18]. Alignment [HKO^{+}22, VRS18, ZZJ18]. Alleviating [ZZW^{+}23]. Allocation [LLJK23, LLM^{+}21, MJBD11, NPS21, PKKK23, ZWL15]. Allocator [KKLL22, LMK06]. Alternative [ÇTNL16, HBL^{+}10, KZL18, MAHK18]. Amdahl [CM08, VMS17]. Amoeba [MPA^{+}18]. Analysis [Ano14c, Ano14d, BY17, BREM08, CNHH15, GGS19, HLH16, HLR21, KCPG18, KKP^{+}18, SRS11, TOIS17, VP16]. Analytical [KZL18, SGBE18]. Analytics [Ano16k, HLR21, KKL20, LZL^{+}20]. Analyzing [NGS15, SQ23]. Annealers [AQ24]. Annual [Ano11a, Ano12a, Ano13a]. Application [CNHH15, CV15, GSG^{+}17, WCC14, ZCG18]. Application-Level [ZCG18]. Application-Specific [WCC14]. Applications [BGZT22, DVAE18, DSVK12, HMCP16, JLA16, KHS^{+}24, KPEC10, LPK16, MHM^{+}24, MLK15, MKD^{+}23, MGI14, MSE^{+}17, ODKK18, SAA^{+}23, VP16, WJA^{+}19, WLL^{+}22]. Approach [BGZT22, CV15, EGWM14, GMM^{+}19, HBW^{+}23, KZL18, LEBM20, PGC22, SBVB17]. Approaches [NGS15]. Approximate [LHPR23]. Approximation [CKZ^{+}20, KQD18, SRLM20]. Arbiter [ZAK^{+}17]. ARCE [RADZ19]. Architecting [SYC14, ZLS10]. Architectural [GD18, HPS23, KNQ15, MKP^{+}24, SMY15, Wu14, ZLM^{+}20, ZWT22]. Architecture [AWD^{+}18, Ano14a, Ano14b, Ano15b, Ano15d, Ano15c, Ano15e, Ano15a, Ano16a, Ano16b, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, ACG^{+}07, BDBS^{+}08, BVL09, DS09, DL20, DM06, Eec22, FFAMK15, Gau09, Jac16a, JPC18, JP13, KWL13, KLSD11, KLKK14, KL02, KR18, LCHL20, LKA15, LYL^{+}16, LJ18, LQYF23, MMR17, OKS^{+}15, PLL08, RADZ19, SRV^{+}19, SRS20, SKS^{+}15, SHK15, Ska13, SCR^{+}17, SJA^{+}17, SHJW21, SCB^{+}20, SM18, VRFT24, WLL^{+}22, YNS^{+}08, YKP^{+}22, ZL18b, ZTRA22, ZZJ18, AD06]. Architecture-Assisted [RADZ19]. Architectures [BRUS21, DXSS15, EOA^{+}23, IXS18, KFJ^{+}03, LLKS12, MTM18, NBH13, RAD^{+}23, RB14, SGBE18, SRT12, WCZ^{+}12, WLL17, XYMY16, XGH^{+}22, XWG^{+}14]. Area [FBN^{+}24, OKS^{+}15, SKS^{+}15]. Area-Efficient [OKS^{+}15, SKS^{+}15]. Argumented [YCH24]. Argus [NS15]. Argus-G [NS15]. Array [AS18, CTL^{+}20, KLCA21, LKKS15, LWM20, WLZZ23].

Arrays [APK^{+}18, SHW19]. ARSENAL [SM18]. Assertions [ZB19]. Assess [Eec22]. Assignment [EOA^{+}23]. Assisted [CST^{+}04, CKA20, DV13, KKL^{+}23, MPPS17, PPG^{+}17, RADZ19]. Associate [Eec13, Mar13a]. Associative [HCM10, KZL18, YKMG15]. Asymmetric [AA19, LBB^{+}19, MNU^{+}15, SCR^{+}17, MWK^{+}06]. Atomic [KLZ12]. atomicity [BLM06]. Attached [BSMB23]. Attack [ASSK21, KYP21, MPA^{+}18]. Attacks [BBZ^{+}19, SQ23]. Attention [LHPR23]. Authors [Ano14b, Ano14d, Ano15d, Ano15e, Ano08d, Ano09n, Ano09o, Ano10o, Ano10p]. Auto [CXS18]. Auto-Tuning [CXS18]. Automata [AS18, AWD^{+}18, SRV^{+}19]. Automata-Processing [AWD^{+}18]. Automated [WLN22]. Automatic [BVL09, LCW^{+}16, YSL^{+}21]. Automatically [MHM^{+}24]. Autonomous [APK^{+}21, KWB^{+}20, MPA^{+}18]. Available [KL18]. AVFs [BREM08]. Aware [AGJ18, APK^{+}21, AS14, CCWY17, DL20, EGWM14, FPA^{+}21, HCM10, JEAG^{+}19, JLS^{+}23, KPKK20, KQGS16, KKKH18, LZS^{+}08, LA16, LQYF23, MNU^{+}15, Mus09, NPS21, PBO^{+}15, SSVS21, SRH20, UKM02, Vol21, YC15, ZTS16, ZKF^{+}18, ZLAE17, IPS14]. Away [GBK^{+}09]. AYUSH [MV15].

B [PGJ12]. B-Fetch [PGJ12]. Back [Ano12j, Ano16p, Ano12d, Ano12e, Ano13c, Ano13d, Ano16c]. Backend [PDGV16]. Backup [MPA^{+}18]. Bad [MCM13]. Balanced [Ant09, FSO^{+}22, GVG^{+}08, SDTG04, Zha06]. Balancing [ILXY18a, ZNTJE23]. Bandwidth [AMW15, KL18, MA19, UTT^{+}24]. Bank [RMA^{+}20, XGH^{+}22]. Banked [RMA^{+}20]. Baobab [TWI^{+}24]. Barrier [CKZ^{+}20]. Basecalling [LJ18]. Basecalling-in-Memory [LJ18]. Based [APK^{+}18, BVL09, CNHH15, CPK^{+}23, DC18, FD08, FBN^{+}24, GLH^{+}20, Hos18, IKW^{+}20, KP21, KWL^{+}17, KL18, KJS^{+}19, KJK21, KKJ^{+}22, KL02, KNE^{+}14, LLKS12, LLSA18, LSJ^{+}19, LCKA23, LZLX15, LHZ19, LJ18, MPPS17, MM03, MAT17, Mus09, NSC20, NGS15, PJ22, PL10, RSRT19, SSVS21, SBVB17, SKTC05, SJM17, SRH20, SRLM20, VGMSLN^{+}18, Yav24, ZZW^{+}22, LAC14, LLLM06, LMK06, LHWB10, yPSS^{+}10, SYC14, HH22, MGH^{+}22, ZTRA22]. Batched [CPK^{+}23]. Bayesian [BHY^{+}19, KDL23, LLM^{+}21, NR21]. BayesTuner [NR21]. BDDs [PV06]. Be [TLG^{+}11]. Behavior [TV02]. Benchmark [ILG10, KL02, WLL17]. Benchmarking [MTM18, XHG^{+}19, ZWT22]. BENoC [WCK08]. Best [SKTC05]. Better [MCM13, YSL^{+}21]. Better-Than-Bad [MCM13]. Between [HSUS11, ILXY18b]. Beyond [Ant09, GVG^{+}08]. Bias [KK21, RZ06]. Bidirectional [LYY^{+}21]. Big [AG17, Ano16k, Jac16a, JLA16, MSE^{+}17]. Big-Data [MSE^{+}17]. BigData [LCHL20]. Bin [WLWZ19]. Binary [LAM^{+}22]. Birkhoff [DC18]. Bit [DPP23, ILXY18a, JAM17, WSVS22]. Bit-Level [ILXY18a]. Bit-Serial [JAM17, WSVS22]. Bitcoin [JLKK23]. Bitstream [KDL23]. Bitwise [SHB^{+}15]. Block [CCWY17, Jac16b, KG10, RB14, TMSA16, VD02, ZM07]. Block- [VD02]. Blockchain [JLKK23]. Blocks [MCM13]. Board [Ano08a, Ano09h, Ano09i, Ano10h, Ano10i, Ano14a, Ano14c, Ano15b, Ano15c]. Boomerang [FHL^{+}10]. Boost [VMS17]. Bootstrapping [KH18]. Bottleneck [AMW15, GGS19, KKP^{+}18, LLD^{+}18]. Bottlenecks [BHL^{+}18]. Bound [SCL13]. Bounded [RSO21]. Bounds [SD04]. Branch [CSSU20, EHdSH20, GAH^{+}23, MHAD15, PGJ12, ST20, SYC07]. BRAWL [LJ18]. Breaking [EHdSH20, LLD^{+}18, SQ23]. Browser [ZWT22]. Browsing [ZSLR14]. Brutus [BGS^{+}20]. BTB [AGK21]. BTB-X [AGK21]. Buffer [ASSK21, KLCA21, SD04, SRLP09]. Bufferless [DPC16, KKK13]. Buffers [LMJ12]. Building [Jac16b, MKD^{+}23, ZM07]. Bulk [SHB^{+}15]. Bursty [HMCP16]. Bus [WCK08]. Bus-Enhanced [WCK08]. Butterfly [KBD07]. BWM [VRS18]. By-Software [GAH^{+}23]. Byte [VHN15]. Byte-Addressable [VHN15].

C [ZAK^{+}17]. C-State [ZAK^{+}17]. Cabinet [Jac16a]. Cache [ALKSA19, AS14, BHL^{+}18, BS17, BGS^{+}20, BGP^{+}17, BSMB23, CWK^{+}22, CCWY17, CZYY11, FJ08, GRCV02, GKKW07, IPS14, JTG23, JP13, KLSD11, KG10, LLJK23, LKP^{+}23, MPPS17, MA19, MCY^{+}12, MCRV07, MKMJ23, NPBS23, OKS^{+}15, PPG11, SSSM18, SKS^{+}15, TV02, VGMSLN^{+}18, VMP^{+}16, WZLQ15, WKE12, XYMY16, YMG14, YPFP14, ZVYW03, ZLAE17, ZWL15, EPS06, Zha06]. Cache-Attached [BSMB23]. Cache-aware [IPS14]. Caches [BLKSA17, BS17, FJ08, JP13, KYP21, LKKS15, MV15, PHBC18, SSVS21, SLKD14, WMJM23, YSL^{+}21, ZS18, Zha06]. Caching [YJZ15]. Calculus [BS17]. call [LLLM06]. CAM [WSVS22]. Can [TLG^{+}11]. Capable [LYR^{+}20]. Capacity [HCK22, SMLS15]. Capsule [HA24]. CARB [ZAK^{+}17]. Carlo [SCL06]. Case [AA19, AS14, EE14, HBL^{+}10, Jac16b, KK21, KWL^{+}17, KKLL22, KR18, NMS14, Per21, PV06, ST20, SRT12, SCL13, Vol21, CMLV03, TD02, Zho06]. CasHMC [JC17]. CAT [LLJK23]. CAVA [CST^{+}04]. CEASER [BGS^{+}20]. Celebrates [Ano10b]. Cells [JDK^{+}02]. Cellular [AS18, CTL^{+}20]. Center [KPKK20]. Centralized [MCKW10]. Centric [HEDH21, KR18]. CF [CXS18]. CF-TUNE [CXS18]. Chaining [KLCA21, MJBD11]. Chains [AQ24]. Challenge [DK13]. Challenges [LG20, RCK21]. Chameleon [YNS^{+}08]. Change [Jun17, KJS^{+}19, KMJ18, Sez10]. Channels [KWKK18, MKMJ23, NAG17]. Characteristics [NBW^{+}23, ZSLR14]. Characterization [DS09, HS04, HLR21, SMY15, WXZ^{+}21, WLL^{+}22]. Characterizing [BKA^{+}09, HXL^{+}22, JSLW20, WYY^{+}23, YCD^{+}20, YZY^{+}22]. Checkpoint [CST^{+}04]. Checkpoint-Assisted [CST^{+}04]. Checkpointing [MAT17]. Chief [Eec13, Gau09, Mar13a, Ska10a, Ska11a, Ska13]. Chip [AGJ18, CGY^{+}14, DOM^{+}07, DOM^{+}08, GQLZ19, GGM^{+}16, GFAHSA24, GKKW07, HCM10, KBD07, KKK13, KDS22, KLZ12, LGLK17, LZS^{+}08, LMJ12, MJBD11, MTT12, PL15, PPG11, RMMLK16, SD02, WCK08, XL07, ZM07, ZNTJE23, ZZW^{+}23, ZKW12, MWK^{+}06, Zho06]. Chip-Multiprocessor [PPG11]. Chipkill [JSDK13]. Chiplet [CWK^{+}22]. Chopping [RSO21]. CIDR [OKS^{+}15, SKS^{+}15]. CIM [KKL^{+}07]. Circuit [JLP07, XJ09]. Circuit-level [XJ09]. Circuit-Switched [JLP07]. Circuits [EOA^{+}23, ZB19]. Claims [BGS^{+}20]. Class [KWKK18]. Classification [SRH20]. Classifications [KKL^{+}07]. Client [MLK15]. Clock [Mic20]. Closing [ILXY18b]. Cloud [DK16, GD18, HLR21, LAX^{+}20, PGR^{+}23, WLL17]. Clumsy [KKK13]. Cluster [DRGA12, MWK^{+}06]. Clustering [CVF^{+}24, SBQK21]. CMA [ZL18a]. CMP [Jac16b, KG10, LMC^{+}09, WCK08]. CMPs [MA19]. CNN [JLRA18, LWM20, SPHS22]. CNNs [WTSW21]. Co [DCG12, KWB^{+}20, LAS22]. Co-Design [KWB^{+}20]. Co-designed [DCG12]. Co-Optimization [LAS22]. Coarse [LYL^{+}16, ZM07]. Coarse-Grain [ZM07].

Coarse-Grained [LYL^{+}16]. Code [ALKSA19, GMMC15, KKL^{+}23, RADZ19]. Code-Pointer [KKL^{+}23]. Codesigned [MKM17]. Coding [CF24, YPFP14]. Cognitive [WL16]. Coherence [BGP^{+}17, BSMB23, CWK^{+}22, JLP07, KLSD11, SLC03, EPS06]. Coherency [BHY^{+}19, MAHK18]. Coherent [MAHK18]. Collaborative [ACG^{+}07, CXS18]. Collabratec [Ano16l, Ano16m]. collection [Ano12k]. Collective [RASW19]. Combinators [AYL22]. Combining [VD02]. Command [GLH^{+}20]. Commands [NBW^{+}23]. Comment [Ant09]. Commercial [GIH^{+}24]. Commit [DV13]. Commodity [TMNK19, Yav24]. Communication [BDJ06, GGM^{+}16, SPAP10, TASA13, LLLM06]. Communications [FJ08, RASW19]. Community [NSC20]. Compact [CGY^{+}14]. comparators [YE07]. Comparing [Man15, SCF04]. Competition [Ano10a]. Compilation [RAD^{+}23]. Compiler [DV13, UKM02, WLDN19]. Compiler-Assisted [DV13]. Compiler-Enabled [UKM02]. Complementary [SYC07]. Complex [ACG^{+}07, ZL18a]. Complexity [GG17, LX08]. Comprehensive [NS15]. Compressed [CEA18]. Compressing [PV06]. Compression [FPA^{+}21, JJP^{+}22, MM03, MVJ17, PBO^{+}15]. CompressPoints [CEA18]. Computation [ACSV02, MLA^{+}14, YHM17, ZB19]. Computational [SAA^{+}23]. Computations [BY17]. Compute [GIH^{+}24, JLRA18, LYL^{+}16, PL10]. Compute-in-SRAM [GIH^{+}24]. Compute-Intensive [LYL^{+}16]. Computer [AKK16, Ano08c, Ano09a, Ano09l, Ano09m, Ano10f, Ano10l, Ano10a, Ano10n, Ano10m, Ano11i, Ano12j, Ano13j, Ano14a, Ano14b, Ano14e, Ano14f, Ano15b, Ano15d, Ano15c, Ano15e, Ano15f, Ano15g, Ano15a, Ano16a, Ano16b, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, BVL09, Eec22, Gau09, KL02, Ska13, TXD^{+}23, Ano10c]. Computers [AG17, DL20, MTH11, Ano10b]. Computing [BSD^{+}19, BREM08, DL19, GJ21, JKK^{+}21, JAM17, KNG^{+}18, KDL23, LHPR23, LJM^{+}14, Man15, WLN22, Wu14, ZL17]. Concurrency [ZWL15]. Concurrent [ODKK18, SK21, ORS^{+}06]. Condition [XYZ15]. Conditions [KCPG18]. Conference [Ano15h, Ano10g, Ano12c]. Confidence [PL10]. Confidentiality [HH22]. Configurable [YLK21]. Configuration [NR21]. Configuring [MSA19]. conflict [Zha06]. Congestion [GF16]. Congestion-Insensitive [GF16]. Connected [Ano10f, Ano13j]. Conquer [CLCG14]. conscious [CMLV03]. Considering [MA19]. Consistency [SJM02, ZLS10]. Constrained [GO15, KPEC10]. Consumption [BKA^{+}09, FHL^{+}10]. Content [KWL^{+}17, Yav24]. Content-Based [KWL^{+}17]. Contention [ASSK21, SBVB17, TV02, WJFH11]. Contents [Ano14g, Ano14h, Ano15j, Ano15k, Ano16n, Ano16o, Ano12h, Ano16p]. Context [SRH20]. Continuous [SRT12]. Control [KKK13, KKL^{+}23, LGLK17, NHKR19]. Control-Flow [KKL^{+}23]. Controlled [ALSJ09, RCS15]. Controller [LLPC19, MGHP20, PDGV16]. conversion [RB14]. Convolutional [GG17, LHZ19, SW19, YKP^{+}22]. Cool [UKM02]. Cool-Fetch [UKM02]. Cooperative [CV15, YJZ15]. Coordinated [NHKR19]. Copies [EE16]. Coprocessor [DEC^{+}18, Jun17]. Copying [KLWJ21]. Core [BHL^{+}18, BEA^{+}13, CVP12, CXS18, DD18, EOA^{+}23, FJ08, GBK^{+}09, IXS18, Jun17, KFJ^{+}03, LMT^{+}09, LA16, MNU^{+}15, NPS21, NSF^{+}18, PKKK23, PHBC18, PL15, SW16, SSS^{+}21, SMY15, XYMY16, ZLAE17, SPAP10]. CoreNap [PKKK23]. Cores [NS15]. Corollaries [CM08]. Correct [JSDK13, KRB^{+}13]. Correction [EE16]. Correlating [GBS^{+}20]. Correlation [SfCL03, SW19]. Cost [DKD07, MAT17, NS15, ZNTJE23]. Count [VGMSLN^{+}18]. Counter [FBN^{+}24, KMJ18, LLSA18, SJM17, RZ06]. Counter-Based [FBN^{+}24, SJM17]. Countermeasure [BGS^{+}20]. Counters [WLWZ19]. counting [Rot08]. Cover [Ano08c, Ano11c, Ano11d, Ano11f, Ano11e, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano08b, Ano09j, Ano09k, Ano10j, Ano10k, Ano11g, Ano11h, Ano12d, Ano12e, Ano12h, Ano12i, Ano12j, Ano13c, Ano13d, Ano13f, Ano13g, Ano16c, Ano16p]. Cover2 [Ano08a, Ano09h, Ano09i, Ano10h, Ano10i, Ano12f]. Cover3 [Ano12g]. Cover4 [Ano09l, Ano09m, Ano10l, Ano10m]. Covert [KWKK18, NAG17]. CPI [EHDH18]. CPS [Ano10g, Ano12c]. CPU [CFM^{+}03, FLSZ17, HDAS18, LMC^{+}09, NMS14, PHO^{+}15]. CPUs [KCPG18]. Creating [MKMJ23]. Critical [BGZT22, GKK^{+}22, ODKK18, PKKK23, TOIS17, ZAK^{+}17]. Criticality [KP21]. Critique [MLA^{+}14]. Cross [LEBM20, SHK15]. Cross-Layer [SHK15]. Cross-Stack [LEBM20]. Crossbar [KZY^{+}19, ZL17]. Cryogenic [RCK21, UTT^{+}24]. Cryptojacking [LEBM20]. CSDP [Ano10d]. Cube [JC17, JPC18]. Customization [LZD^{+}23]. CXL [SAA^{+}23]. CXL-Memory [SAA^{+}23]. Cyber [Ano16d, KWB^{+}20]. Cyber-Physical [KWB^{+}20]. Cybersecurity [Ano15h]. Cycle [JC17, KHB^{+}19, LYR^{+}20, MJBD11, MMAAK21, RL17, RCBJ11]. Cycle-Accurate [JC17, KHB^{+}19, LYR^{+}20]. Cycle-Level [MMAAK21]. Cyclic [ÇTNL16].

D [RMMLK16, ZBA^{+}20]. DAEGEN [WLDN19]. Dagger [LAX^{+}20]. DAM [SSVS21]. DAMARU [KYP21]. Danger [SKTC05]. Dark [CMP^{+}14, DXSS15, TNC19]. Data [AG17, AD06, ASK^{+}21, Ano16k, BLKSA17, BBZ^{+}19, DK16, FPA^{+}21, GBS^{+}20, HCK^{+}21, HLH16, HH22, Jac16a, KP21, KPKK20, KWL^{+}17, KJK21, KLR24, KLZ12, LPK16, LZL^{+}20, MCM13, MAT17, MVJ17, MSE^{+}17, NSF^{+}18, RL17, RMA^{+}20, VMP^{+}16, YKP^{+}22, ZZW^{+}22]. Data-Aware [FPA^{+}21]. Data-Dependent [KWL^{+}17]. Database [CSSU20]. Datacenter [DSVK12, DK13, KQD18, LMT^{+}09, LLS^{+}15]. Datacenters [SG14]. Dataflow [KLCA21, WLL^{+}22]. Datatype [WKE12]. Day [RTKQ21]. DC [MDSG20]. DCC [KLSD11]. DDMR [GWR08]. Deadblock [SSVS21]. Deadlock [LX08, XL07, XYZ15]. Deadlock-Free [XYZ15]. Debugging [CVP12]. Decay [JDK^{+}02]. Decoder [YWG17]. Decomposition [LCKA23]. Decomposition-Based [LCKA23]. Decongest [WMZY17]. Decoupled [IXS18, KJK21, PTND24, WLDN19]. Decoupling [DSVK12, SLC03]. Deduplicating [SMLS15]. Deduplication [APK^{+}18]. Deep [GMM^{+}19, HLR21, JE22, JKK^{+}21, JAM17, KHS^{+}24, KR18, LLPC19, LTL23, NHKR19, SCB^{+}20]. Deeply [ILXY18a, ILXY18b]. Defending [LEBM20]. Defense [MPA^{+}18, WYY^{+}23]. Delay [Cit04, SD04]. Demand [MHAD15]. DeMM [PTND24]. Demotion [MDK^{+}23]. Demystifying [Mic13]. Denial [KYP21]. Denial-of-Service [KYP21]. Dense [WMZY17]. Dependable [KLSD11]. Dependence [GGS19, TOIS17]. Dependency [PS17]. Dependent [KWL^{+}17, MCM13]. Design [AS18, Ano10a, ACG^{+}07, CTL^{+}20, HRF^{+}11, HBW^{+}23, KNG^{+}18, KWB^{+}20, LKA15, LLPC19, PLK^{+}23, SKK22, TASA13, TDO16, VMP^{+}16, WL16, YHM17]. designed [DCG12]. Designs [KSO^{+}16, XHG^{+}19]. Detailed [XCW^{+}19, YLK21]. Detect [WLWZ19]. Detecting [LG20, YE07]. Detection [KWL^{+}17, KJS^{+}19, LX08, MMR17, NS15, XL07, ZL18a]. Detection-Based [KJS^{+}19]. Determining [BHY^{+}19]. Determinism [RSO21]. Deterministic [Man15, Mic20, ODKK18]. Development [ACG^{+}07]. Device [HSUS11]. Devices [GLH^{+}20, WLZZ23]. Die-Stacked [SFFG^{+}19]. Differential [BS17, GMPMC^{+}23]. Digital [Ano09a, Ano10c, SPHS22]. Dimensional [RL08]. DIMM [ALSJ09]. Direct [CF24, JLS^{+}23, NSF^{+}18, Zha06]. Direct-Coding [CF24]. direct-mapped [Zha06]. Directed [PGJ12, ZMC17]. Directional [LCHL20]. Directory [HR10]. Disaggregated [Vol21]. Discovering [BGZT22, NBW^{+}23]. Discrete [SRT12]. Discrete-Continuous [SRT12]. Disintermediated [BDJ06]. Disk [YNS^{+}08]. Distance [BY17]. Distinguish [Ano10d]. Distributed [AKK16, CZYY11, FD08, MGH^{+}22, SSS^{+}21, SLKD14, SB18, SRLP09, YJZ15, YP23]. Distribution [SK21]. Disturbance [MVJ17, WMZY17, WLWZ19]. Divergence [ZTS16]. Divergent [WJA^{+}19]. Diversity [TDO16]. Divide [CLCG14, ZKW12]. Divide-and-Conquer [CLCG14]. DMA [MAHK18]. DNA [CF24, HKO^{+}22]. DNN [HCK^{+}21, JKK^{+}21, KPPK21, MMAAK21, NR21, SBQK21, YP23]. DNNs [KKJ^{+}22, RAD^{+}23]. Domain [GGM^{+}16, ST20]. Domain-Specialized [ST20]. Dot [AS18]. Down [EGWM14]. DRACO [SMLS15]. DRAM [EHH21, ILXY18b, KWL^{+}17, KNQ15, KYM16, KKKH18, KKLL22, LLKS12, LKK19, LYR^{+}20, MCY^{+}12, MAT17, MGHP20, NBW^{+}23, OKS^{+}15, SSSM18, SKS^{+}15, SHB^{+}15, SJM17, SPHS22, SCR^{+}17, TMNK19, WLWZ19, WSVS22, XGH^{+}22, Yav24, YYK^{+}18, ZTRA22]. DRAM-CAM [WSVS22]. DRAM-NVM [KKLL22]. DRAMA [FFAMK15, Yav24]. DRAMs [ALSJ09]. DRAMSim2 [RCBJ11]. DRAMsim3 [LYR^{+}20]. Drive [SYC14]. Driven [MLM^{+}06]. Drives [JZA^{+}18, KKL20]. Dual [GWR08, MTT12]. Due [RCS15]. Duplication [KRB^{+}13, MVJ17]. DVFaaS [TMSX23]. DVFS [CLCG14, NHKR19, RCS15, TMSX23]. Dynamic [CFM^{+}03, GWR08, GMM^{+}19, HCM10, JMKP07, JMKP08, KK21, KCP^{+}19, KDS22, LLJK23, LMK06, MHAD15, MCRV07, RMMLK16, SPJ02, SCF04, SKD09, YC15, ZB19]. Dynamically [MSA19, WTSW21].

eADR [HH22]. eADR-Based [HH22]. Early [NBH13]. Early-Stage [NBH13]. EARtH [EGWM14]. Easy [MKD^{+}23, Ano12k]. ECC [RK22]. Ecosystem [AWD^{+}18]. Edge [DL19, KKL^{+}23, GGS19]. Edition [DK13]. Editor [Eec13, Gau09, Mar13a, Ska09a, Ska10a, Ska11a, Ska13]. Editor-in-Chief [Eec13, Gau09, Mar13a, Ska10a, Ska11a, Ska13]. Editorial [Ano08a, Ano09h, Ano09i, Ano10h, Ano10i, Ano14a, Ano14c, Ano15b, Ano15c, Mar13b, Ska10a, Ska11a]. Editors [Mar13a, Eec13]. eDKM [CVF^{+}24]. Effective [AGK21, HRF^{+}11]. Effects [MTT12]. Efficiency [HA24, IXS19, JSLW20, KCP^{+}19, KQD18, LLS^{+}15, SKTC05, VHN15, MWK^{+}06]. Efficient [AYL22, AG17, ALSJ09, BLKSA17, BDBS^{+}08, BGP^{+}17, CGY^{+}14, CVF^{+}24, CLCG14, CHK^{+}18, CXS18, DM06, GDF^{+}04, HMCP16, HR10, JSDK13, JJP^{+}22, KP21, KDL23, LAX^{+}20, LWM20, MCY^{+}12, MJBD11, OKS^{+}15, PKKK23, SRV^{+}19, SKS^{+}15, TLG^{+}11, TWI^{+}24, WCK08, YHM17, YP23, ZL18b, ZSLR14, SPJ02]. Efficiently [LJ04]. EH [SGBE18]. Electromagnetic [HDAS18]. Emanations [HDAS18]. Embedded [BDBS^{+}08, CLJ^{+}02, DS09, GGM^{+}16, GRCV02, ILG10, MAHK18, PPG^{+}17, RADZ19, SKA^{+}20, TLG^{+}11, YC15]. Embedding [KKJ^{+}22, LQYF23]. Embedding-Aware [LQYF23]. Embedding-Based [KKJ^{+}22]. Emerging [SQ23, WJA^{+}19]. Employing [LGLK17]. Enabled [KKL20, UKM02, ZL17]. Enabling [MCY^{+}12, MMAAK21, SRS20, SMZ18, WLN22]. Enclave [NK22]. Encrypted [LGLK17]. Encryption [KMJ18, RM18]. End [GF16, HXL^{+}22]. End-Point [GF16]. End-to-End [HXL^{+}22]. Endurance [PLK^{+}23, YPFP14]. Energy [ALSJ09, BKA^{+}09, BDBS^{+}08, CV15, CM08, CLCG14, CXS18, DL20, EGWM14, GJ21, GO15, HA24, JSLW20, JSDK13, KP21, KQGS16, KDL23, KKL^{+}15, KPEC10, LJM^{+}14, PKKK23, SGBE18, TLG^{+}11, VHN15, Wu14, ZVYW03, ZL18b, ZSLR14]. Energy-Constrained [KPEC10]. Energy-Efficiency [HA24, VHN15]. Energy-Efficient [BDBS^{+}08, KP21, KDL23, TLG^{+}11, ZSLR14]. Energy-Harvesting [DL20, GJ21, SGBE18]. Enforced [MS16]. Engine [LTL23, OK22, PTND24]. Engines [NK22]. Enhance [FJ08, SJM02, TMSA16]. Enhanced [KRB^{+}13, TOIS17, WCK08]. enhancement [Zho06]. Enhancing [AQ24, GLJ^{+}21, HBW^{+}23, VMP^{+}16]. Ensuring [HH22]. Entangling [RJ20]. Enterprise [LHCK22]. Entropy [Cit04]. Environment [ACG^{+}07, CVP12, TXD^{+}23]. Environments [KKH14]. Epoch [CNHH15]. Equality [YHY^{+}22]. Equations [BS17, GMPMC^{+}23]. Era [CMP^{+}14, SSS^{+}21]. Error [EE16, EUVG06, MMR17, NBW^{+}23, NS15, PL15, RTKQ21, Vol21, WLWZ19]. Errors [GSG^{+}17, KRB^{+}13, YE07]. Estimate [SW16]. Estimating [CFM^{+}03]. Estimation [FAR^{+}23]. Evaluate [EE14, KKL^{+}15]. Evaluating [KKL^{+}07, LJ04, WLL17]. Evaluation [CEA18, KSO^{+}16, SJA^{+}17]. Evasive [LG20]. Exact [WSVS22]. Example [GRCV02]. Exascale [Jac16b, Jac16a]. Exceeding [SfCL03]. Exchange [NSF^{+}18]. Executed [MKSP05, WB14]. Execution [AWD^{+}18, BBZ^{+}19, CSSU20, HMCP16, IXS19, KKL^{+}15, LLD^{+}18, MLK15, MKSP05, NFAE19, ODKK18, TXD^{+}23, ZTS16]. Existing [EE16]. Expander [HCK^{+}21]. Expectations [YMBA19]. Expected [VGMSLN^{+}18]. Experience [Ano16k, CZYY11]. Expert [PB16]. Explaining [MCRV07]. Explicit [BHD09]. Exploit [ZLAE17]. Exploiting [ÇE14, Cit04, EE16, EUVG06, GRCV02, GG11, KWKK18, LKK19, Mic20, yPSS^{+}10, SBQK21, XJ09, ZSLR14]. Exploration [LLPC19, SGBE18]. Explore [BSD^{+}19]. Exploring [BHL^{+}18, CSSU20, HSUS11, NPBS23, SHJW21, WLDN19]. Extending [JP13, MV15, VMS17]. Extensible [KYM16, MGHP20]. Extra [SMLS15]. EZ [ZL18b]. EZ-Pass [ZL18b].

FaaS [TMSX23]. Fabric [ZL17]. Facilitate [ZLS10]. Failures [KWL^{+}17, SG14]. Fairness [MA19, VS11]. Fast [KYM16, LGLK17, MKD^{+}23, SMZ18, SHB^{+}15, YP23]. FastDrain [ZKH^{+}20]. Fat [Ant09, GVG^{+}08]. Fat-tree [Ant09, GVG^{+}08]. Fault [GDF^{+}04, HRF^{+}11, ZKF^{+}18, Zho06]. Fault-Aware [ZKF^{+}18]. Fault-Tolerant [GDF^{+}04, HRF^{+}11]. Faults [OKS^{+}15, SKS^{+}15]. Feature [YSL^{+}21]. FESSD [LGLK17]. Fetch [MSA19, UKM02, AGJ18, PGJ12, UKM02]. Fighting [AMW15]. File [EE16, JEAG^{+}19]. Filter [GF16, HKO^{+}22]. Filtering [CXS18]. find [Ano12k]. Fine [BRUS21, MKM17, MCY^{+}12, MKP^{+}24, WYM^{+}16]. Fine-Grained [BRUS21, MKP^{+}24, WYM^{+}16]. Fine-Granularity [MCY^{+}12]. Firmware [BGL^{+}23]. First [CAPS09, Eec22]. First-Order [Eec22]. Fit [LWB13]. Fixed [GRCV02]. Flash [IKW^{+}20, KJK21, LKA15, LZLX15, yPSS^{+}10, SYC14, YNS^{+}08]. Flash-Based [LZLX15, yPSS^{+}10, SYC14]. Flash/FRAM [YNS^{+}08]. Flattened [KBD07]. Flexibility [KPPK21, TND^{+}21]. Flexible [LQYF23, LWB13, XCW^{+}19]. Flexion [KPPK21]. FlexScore [TND^{+}21]. Floating [ACSV02, DKD07, LTL23]. Floating-Point [DKD07, LTL23]. Flow [Hos18, KKK13, KKL^{+}23, MSE^{+}17]. Flow-Based [Hos18]. Footprint [SW16]. Foreword [GPS06]. Format [LZD^{+}23]. Forward [ASSK21, KKL^{+}23, NB24]. Forward-Edge [KKL^{+}23]. Forwarding [BHD09]. FPGA [FLSZ17, KLR24, LAC14, LZL^{+}20, MGH^{+}22, PP12]. FPGA-Accelerated [KLR24]. FPGA-Based [MGH^{+}22, LAC14]. FPGAs [LCHL20, RAD^{+}23, SKK22]. Fractal [ZLS10]. FRAM [YNS^{+}08]. Framework [BSD^{+}19, BVL09, CYAW20, JJP^{+}22, KLZ12, LHZ19, LWB13, MHM^{+}24, SKK22, TMNK19, LHWB10]. Free [GFAHSA24, PS17, XYZ15]. Frequency [ÇTNL16, MLM^{+}06, Mic20, YC15]. Friendly [LHPR23, LCKA23, PZX15]. Front [Ano08b, Ano09j, Ano09k, Ano10j, Ano10k, Ano11g, Ano11h, Ano12h, Ano12i, Ano13f, Ano13g]. FTL [SMLS15]. Fully [ZL17]. Function [LLKS12]. Functional [CAPS09, DCG12]. Functional-First [CAPS09]. Functions [TD02]. Fusion [LYY^{+}21]. Fuzzy [ACSV02].

G [NS15]. Gap [ILXY18b]. Gating [ÇTNL16, LMT^{+}09, ZL18b]. GCMS [WJFH11]. GCN [YSL^{+}21]. GCNs [LYY^{+}21, YCD^{+}20]. gem5 [RSRT19, AKK16, PHO^{+}15, VRFT24]. gem5-accel [VRFT24]. gem5-gpu [PHO^{+}15]. GEMM [LWM20]. General [DPP23, LZD^{+}23, WSVS22]. General-Purpose [WSVS22]. Generalized [AS18, CTL^{+}20, GO15, MMY^{+}14]. Generate [MHM^{+}24]. Generation [BVL09, GMMC15, JLKK23, SKK22]. Generative [CPK^{+}23]. Genome [LJ18]. Global [KK21, MPPS17, WJFH11]. Globally [SDTG04]. GNNs [WYY^{+}23]. Goal [TDO16]. GP [JJP^{+}22]. GP-GPU [JJP^{+}22]. GPGPU [CCWY17, LLKS12, NS15, SW16, ZCG18]. GPGPUs [NAG17, SSSM18, ZLAE17, ZWL15]. GPU [ABC^{+}19, IXS19, JSLW20, JEAG^{+}19, JJP^{+}22, KLKK14, KCP^{+}19, LSJ^{+}19, RASW19, WCYC09, WJA^{+}19, XWG^{+}14, YCD^{+}20, PHO^{+}15]. GPU-NEST [JSLW20]. GPUs [HA24, HCK^{+}21, HCK22, HXL^{+}22, NMS14, NSF^{+}18, PGR^{+}23, PBO^{+}15, WYM^{+}16, WYY^{+}23, YZY^{+}22, YC15, ZTS16, ZNTJE23, ZZW^{+}23]. Grain [MKM17, ZM07]. Grained [BRUS21, LYL^{+}16, MKP^{+}24, WYM^{+}16]. GraNDe [YKP^{+}22]. Granular [MNFI20, YJZ15]. Granularity [MCY^{+}12]. Graph [AYL22, BY17, BHL^{+}18, FSO^{+}22, GKK^{+}22, LKR21, NSC20, NGS15, ST20, SHJW21, TOIS17, YKP^{+}22, ZLM^{+}20, ZZW^{+}22]. Graph-Based [NGS15]. Graph-Processing [ST20]. Graphs [GGS19]. GraphSCC [NSC20]. Greedy [DC18]. GreenRouter [KWL13]. Guarantee [NB24]. Guard [MKMJ23]. Guessing [WTSW21]. Guiding [BY17].

HAD-TWL [KJS^{+}19]. Half [MTT12]. Half-Speed [MTT12]. Halt [EGWM14]. Halting [ZVYW03]. Hammer [LKP^{+}23, LHPR23]. Hammering [KNQ15, LLSA18, SJM17]. Hardware [AGJ18, AW15, CTJ^{+}17, CWK^{+}22, CV15, CKA20, DVAE18, DD18, GKK^{+}22, JLKK23, KP21, KKL^{+}23, KH18, LHPR23, LAS22, LMK06, LCW^{+}16, LYY^{+}21, LCKA23, LAM^{+}22, MLK15, MKM17, MS16, NB24, NGS15, OK22, PB16, SK21, SPHS22, WJFH11, WLL17, XL07, YCH24, ZS18]. Hardware-Assisted [CKA20, KKL^{+}23]. Hardware-Friendly [LHPR23, LCKA23]. Hardware-Software [CV15]. Hardware/Software [MKM17]. Harmonic [PL10]. Harnessing [GBS^{+}20]. Harvesting [DL20, GJ21, SGBE18, Wu14]. Hashing [SMZ18]. HBM3 [GLJ^{+}21]. HCI [VMP^{+}16]. Headers [JLKK23]. Heavy [SSTS17]. Heterogeneity [APK^{+}21, MTH11]. Heterogeneity-Aware [APK^{+}21]. Heterogeneous [AEJE17, BRUS21, DL20, FLSZ17, GO15, GMM^{+}19, HCK22, KFJ^{+}03, LLS^{+}15, MMY^{+}14, PHO^{+}15, TDO16, TMNK19, ZBA^{+}20, ZKW12]. Heterogeneous-ISA [BRUS21]. Heterogeneous-Reliability [TMNK19]. HeteroSim [FLSZ17]. Heuristics [MGI14]. HGNNs [YZY^{+}22]. Hiding [CST^{+}04]. Hierarchical [BSBD^{+}08, SKA^{+}20]. Hierarchy [BHL^{+}18, YMG14, ZM07]. High [CTL^{+}20, DPC16, JSDK13, KKK13, KL18, LTL23, PLK^{+}23, PP12, RMMLK16, RB14, SKK22, SD04, SYC14, SRLP09, SHJW21, TASA13, YPFP14, YNS^{+}08, ZVYW03, LHWB10]. High-Bandwidth [KL18]. High-Endurance [PLK^{+}23]. High-Level [PP12]. High-Performance [CTL^{+}20, DPC16, LTL23, PLK^{+}23, RMMLK16, SKK22, SHJW21, TASA13, ZVYW03, SYC14, LHWB10]. High-Throughput [KKK13, SRLP09]. Highly [KL18, RMA^{+}20]. Highly-Banked [RMA^{+}20]. HiLITE [SKA^{+}20]. Hit [VGMSLN^{+}18]. HLS [KDS22]. HMC [JPC18]. HMC-MAC [JPC18]. Holes [AEJE17]. Holistic [JZA^{+}18, KSO^{+}16]. Homogeneous [MTH11]. Horizontal [GG11]. Hot [KJS^{+}19, WMZY17]. HPC [KR18]. Hungarian [EOA^{+}23]. HW [APK^{+}18, DCG12]. HW-Based [APK^{+}18]. HW/SW [DCG12]. Hy [NPS21]. Hy-Sched [NPS21]. Hybrid [JC17, JPC18, JP13, KKLL22, KSB19, LMK06, MCY^{+}12, MV15, SSVS21, SRT12, YNS^{+}08, YYK^{+}18]. Hyperthreading [NPS21]. Hyperthreading-Aware [NPS21]. Hypervisor [PPG^{+}17]. hysteresis [RZ06].

I/O [JLS^{+}23, KLWJ21, LKA15, LKKS15, MAHK18, SYC14]. I/O-Intensive [JLS^{+}23]. IBM [LCW^{+}16]. Ideal [ALKSA19]. Ideas [JLA16]. IDIO [ASK^{+}21]. IEEE [Ano08c, Ano09a, Ano09l, Ano09m, Ano10f, Ano10l, Ano10a, Ano10n, Ano10m, Ano11i, Ano12j, Ano13h, Ano13i, Ano13j, Ano14e, Ano14f, Ano15f, Ano15g, Ano16l, Ano16m, Ano10b, Ano13e, Ano14a, Ano14b, Ano14c, Ano14d, Ano15b, Ano15d, Ano15c, Ano15e, Ano15a, Ano16d, Ano16a, Ano16b, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, Gau09, Ska13]. IF [RB14]. IF-conversion [RB14]. IK [YCH24]. IMEC [ZL17]. Imitation [SKA^{+}20]. Immediate [EHH21]. Immediate-Response [EHH21]. Impact [FHL^{+}10, GSG^{+}17]. Impacts [WKE12]. Implementation [LAM^{+}22, WLL^{+}22]. Implementing [JDK^{+}02, TMNK19]. Implication [LKR21]. Implications [DK16, GD18, OSH16, ZLM^{+}20]. Improve [KH18, KQD18, MMR17, WMJM23, XJ09]. Improved [DKD07, PGR^{+}23]. Improvement [MA19]. Improving [CCWY17, CZYY11, HA24, IXS19, ILXY18a, KCP^{+}19, LLS^{+}15, MSA19]. In-DRAM [MAT17, XGH^{+}22]. in-Hardware [SK21]. In-Line [LAC14]. In-Memory [CSSU20, CHK^{+}18, LAS22, SRV^{+}19, WLN22, ZL17]. In-network [EPS06]. In-Order [EHdSH20, HEDH21, PGJ12]. In-SRAM [SRS20]. Inbound [ASK^{+}21]. Including [DRGA12]. Increasing [ÇE14]. Incremental [MAT17]. Independent [DS09, LKKS15, WLZZ23]. Independently [ALSJ09]. Index [Ano11a, Ano12a, Ano13a, Ano15a, Ano16a, Ano16b, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, WMJM23]. Indirect [JMKP07, JMKP08]. Induced [DXSS15]. Industry [ILNS20]. Inference [CPK^{+}23, DL19, JSLW20, KDL23, KKL20, LWM20, MMAAK21, NR21]. Inference-Enabled [KKL20]. Infinity [WLN22]. Information [Ano08d, Ano09n, Ano09o, Ano10o, Ano10p, Ano14b, Ano14d, Ano15d, Ano15e, Ano11i, Ano11j]. Infrastructure [AKK16]. Initial [ACSV02]. Inline [APK^{+}18]. Innovating [KWL13]. Inputs [BEA^{+}13]. Insensitive [GF16]. Insertion [JTG23]. Inspired [OKS^{+}15, SKS^{+}15]. Instruction [ALKSA19, BSBD^{+}08, GIH^{+}24, ILNS20, KP21, MMR17, MSA19, RYSN04, RJ20, WCZ^{+}12, Zha06]. Instructions [ASSK21, MKSP05, WB14]. Integrated [NMS14, OK22]. Integrating [FAR^{+}23]. Integration [Jun17, ZBA^{+}20]. Integrity [KKL^{+}23, RADZ19, SB18]. Intel [CLCG14, MDSG20]. Intelligence [Ano14c, Ano14d]. Intelligent [BGL^{+}23]. Intensive [JLS^{+}23, LYL^{+}16, SAA^{+}23]. Inter [GGM^{+}16, LA16, NSF^{+}18, RASW19, RMA^{+}20, SPAP10, UTT^{+}24, ZTS16, ZLAE17]. Inter-Bank [RMA^{+}20]. Inter-Core [LA16, NSF^{+}18, ZLAE17]. Inter-Domain [GGM^{+}16]. Inter-GPU [RASW19]. Inter-Socket [SPAP10]. Inter-Temperature [UTT^{+}24]. Inter-Warp [ZTS16]. Interaction [HSUS11]. Interconnect [CGY^{+}14, KG10, SRV^{+}19]. Interconnection [Ant09, GVG^{+}08, SPJ02, SD04, GD06].

Interface [BHY^{+}19]. Interference

[ASSK21]. Interleaving [LLJK23, VD02]. Intermediate [LZD^{+}23, WXZ^{+}21]. Internal [NBW^{+}23, yPSS^{+}10]. Interpreter [MSI18]. Interval [SKTC05]. Interval-Based [SKTC05]. Intervals [GWR08, PL10]. Intra [SPAP10]. Intra-Socket [SPAP10]. Intrinsic [MMR17]. Introducing [Ano16l, Ano16m, Gau09, Ska13]. Introduction [Eec13, Mar13a]. Intrusion [ZL18a]. Intrusive [PDGV16]. Invariant [ASSK21]. Inverse [YCH24]. IP [KL18]. IPC [EE14]. Irregular [CLCG14]. ISA [BRUS21, KFJ^{+}03, MNU^{+}15, WXZ^{+}21]. Isolating [BBZ^{+}19]. Isolation [ODKK18]. Issue [MVJ17, RYSN04]. Issuing [NBW^{+}23].

Jacobian [YCH24]. JANM [YCH24]. JANM-IK [YCH24]. Jaseci [MKD^{+}23]. Java [DS09]. JavaScript [VP16]. Jobs [Ano10n]. Journaling [BGL^{+}23]. Jumps [JMKP07, JMKP08].

Kernel [NMS14]. Key [KKLL22, PLK^{+}23]. Key-Value [PLK^{+}23]. Kinematics [YCH24]. Kobold [BSMB23]. KSM [ZCG18]. KV [PJ22]. kW [Jac16a].

L1 [BLKSA17, PHBC18, VMP^{+}16]. L2 [CST^{+}04]. L3 [FJ08]. LA-LLC [ZLAE17]. LADIO [JLS^{+}23]. Language [CVF^{+}24, LZD^{+}23]. Large [CVF^{+}24, CPK^{+}23, DRGA12, DSVK12, HCM10, JLA16, LKR21, PLK^{+}23, SG14]. Large-Key [PLK^{+}23]. Large-Scale [DRGA12, DSVK12, LKR21]. Last [JTG23, KYP21, LKP^{+}23, YPFP14, ZLAE17]. Last-Level [JTG23, KYP21, LKP^{+}23, ZLAE17]. Latency [KJS^{+}19, KLWJ21, LTL23, NPBS23, PKKK23, SCR^{+}17, ZAK^{+}17, ZZW^{+}23].

Latency-Critical [PKKK23, ZAK^{+}17]. Latency-Versatile [LTL23]. Law [CM08, VMS17]. Layer [KSO^{+}16, SHK15]. Layout [ALKSA19]. LazyPIM [BGP^{+}17]. Leakage [FAR^{+}23, JLS^{+}23]. Leakage-Aware [JLS^{+}23]. Learned [LKL^{+}21]. Learning [GMM^{+}19, HLR21, JE22, KHS^{+}24, LLPC19, SKA^{+}20, SCB^{+}20, WMJM23, YG18, ZZW^{+}22]. Learning-Based [ZZW^{+}22]. LEO [RM18]. Letter [Ska09a, Ska10a, Ska11a]. Letters [Ano14a, Ano15b, Ano15c, Ano14b, Ano15d, Ano15e, Ano15a, Ano16a, Ano16b, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, Gau09, Ska13]. Level [ILXY18a, JTG23, KHS^{+}24, KYP21, LKK19, LKP^{+}23, LMJ12, MGI14, MMAAK21, PJ22, PP12, TV02, TMSA16, VE18, YPFP14, ZLAE17, ZCG18, LLLM06, XJ09]. Leveling [KJS^{+}19, LZLX15, ZKF^{+}18]. Leveraging [DD18, KG10, KQD18, LMJ12, LLS^{+}15, MXS19, NR21, TMSX23, WZLQ15, ZS18]. Library [ACG^{+}07, Ano09a, Ano10c]. Life [RTKQ21]. Lifetime [BSD^{+}19, HSUS11, JP13, MV15, SMY15]. Light [IXS19]. Light-Weight [IXS19]. Lightweight [CYAW20, LAM^{+}22, SKA^{+}20, ZZW^{+}22]. like [WCZ^{+}12]. Limit [KWB^{+}20]. Limited [AEJE17]. Limits [CTJ^{+}17, SfCL03]. Line [LAC14]. Linearization [LHPR23]. Link [HRF^{+}11, SCF04]. Links [SPJ02]. List [Ano11b, Ano12b, Ano13b]. LLC [KKH14, ZLAE17]. LLVM [RSRT19]. LLVM-Based [RSRT19]. LMT [SJ22]. Load [Ant09, GVG^{+}08, HR10, ILXY18a, SDTG04]. Load-Balanced [Ant09, SDTG04]. Load-Load [HR10]. Locality [BY17, CCWY17, EF07, GG11, JEAG^{+}19, LA16, SfCL03, SRLM20, XJ09, ZLAE17]. Locality-Aware [JEAG^{+}19, ZLAE17]. Lock [MNU^{+}15]. loft [IPS14]. LogCA [AW15]. Logic [FD08, TNC19].

Logic-Based [FD08]. Long [ZZW^{+}23].

Longer [AQ24]. LOOG [IXS19]. Lookaside [LMJ12]. Lookup [KL18]. Lookups [CSSU20]. Loop [GRCV02]. Loops [GAH^{+}23]. Low [CLJ^{+}02, Cit04, DKD07, GG17, GFAHSA24, KJS^{+}19, KLWJ21, MAT17, NS15, PHBC18, RM18, SRS20, YPFP14, ZVYW03, LHWB10, MTT12]. Low-Cost [DKD07, MAT17, NS15]. Low-Energy [ZVYW03]. Low-Latency [KLWJ21]. Low-Overhead [SRS20]. Low-Power [GFAHSA24, PHBC18, LHWB10]. LSM [PJ22]. LSTM [MXS19]. LT [ZTRA22]. LT-PIM [ZTRA22]. LUT [ZTRA22]. LUT-Based [ZTRA22]. LV [LTL23]. LWE [LAM^{+}22].

MAC [JPC18]. Machine

[Ano14c, Ano14d, DL19, YG18]. Machines [GBK^{+}09, KWB^{+}20, UTT^{+}24]. Main [Sez10, YYK^{+}18]. Making [YSL^{+}21]. Manage [MCM13]. Managed [GMMC15]. Management [CFM^{+}03, EGWM14, GMM^{+}19, KPKK20, KWKK18, LKK19, LKP^{+}23, LMT^{+}09, MPPS17, MCY^{+}12, MAHK18, MKP^{+}24, PGR^{+}23, RADZ19, SKA^{+}20, SSS^{+}21, SMY15, TMSA16, WJFH11, ZAK^{+}17]. Managing [DOM^{+}07, DOM^{+}08, JE22]. Many [BHY^{+}19, CXS18, DXSS15, GBK^{+}09, KDS22, NBH13, PHBC18, SSS^{+}21, SMY15, XYMY16, ZLAE17, ZZW^{+}23]. Many-Accelerator [BHY^{+}19, DXSS15, KDS22, NBH13]. Many-Chip-Module [ZZW^{+}23]. Many-Core [CXS18, GBK^{+}09, PHBC18, SSS^{+}21, SMY15, XYMY16]. Many-Thread [GBK^{+}09]. Many-to-Many [ZLAE17]. mapped [Zha06]. Mapping [EOA^{+}23, HLH16, LHZ19, YKP^{+}22]. MapReduce [IXS18, LYL^{+}16]. Massive [Mus09, SMZ18]. Massively [ADS^{+}19]. match [YE07]. Matching [PLL08, WSVS22, ZL18a]. Matrix [LCKA23, MNFI20, PTND24, YKP^{+}22]. MCsim [MGHP20]. Mead [YCH24]. Mean [LHPR23, PL10]. Mean-Redistribution [LHPR23]. Measuring [GSG^{+}17]. Mechanism [BGP^{+}17]. Mechanisms [RCS15, TVB^{+}13, XYMY16]. Memcached [LAC14]. Memoization [ZS18]. Memories [KHB^{+}19, KNQ15, KDS22, KZY^{+}19, MCY^{+}12, RM18, RMA^{+}20, SM18, WCC14]. Memory [ALSJ09, AA19, AMW15, BKA^{+}09, BGP^{+}17, CSSU20, CYAW20, CKA20, CEA18, CHK^{+}18, CMP^{+}14, DXSS15, DD18, EHH21, FFAMK15, GSG^{+}17, GLH^{+}20, HCK^{+}21, HKO^{+}22, HCK22, IXS18, JC17, JPC18, JSDK13, JJP^{+}22, JDK^{+}02, JLA16, Jun17, JLRA18, KQGS16, KWL^{+}17, KNG^{+}18, KLKK14, KL18, KJS^{+}19, KPL^{+}21, KKLL22, KKJ^{+}22, KMJ18, KR18, LAX^{+}20, LGLK17, LKK19, LKR21, LLJK23, LAS22, LMK06, LA16, LZLX15, LLPC19, LJ18, LQYF23, MDSG20, MGHP20, MDK^{+}23, NBW^{+}23, OK22, PS17, PZX15, RMA^{+}20, RCBJ11, SRV^{+}19, Sez10, SPHS22, SB18, SAA^{+}23, SFFG^{+}19, SCB^{+}20, TWI^{+}24, VE18, VHN15, Vol21, WJFH11, WYL^{+}15, WJA^{+}19, WLN22, WTSW21, WLZZ23, XWG^{+}14, XCW^{+}19, Yav24, YLK21, YYK^{+}18, ZM07, ZL17, ZLS10, ZZJ18, BLM06, SAA^{+}23]. Memory-Centric [KR18]. Memory-Divergent [WJA^{+}19]. Memory-Error [Vol21]. Memory-Induced [DXSS15]. Memory-Intensive [SAA^{+}23]. Memory-Level [VE18]. Memory-Unaware [KLKK14]. Memristor [KNE^{+}14]. Memristor-Based [KNE^{+}14]. Merkle [JLKK23, TWI^{+}24]. Mesh [RL08, SCL13, XYZ15]. Meshes [GDF^{+}04]. Message [Eec13, GGM^{+}16, Mar13a]. Metadata [GBS^{+}20, MKP^{+}24, RADZ19]. Method [LPK16]. Methodology [CEA18, DM06, GDF^{+}04, Hos18, WL16].

Methods [WYY^{+}23]. Metric

[KKL^{+}07, KPPK21]. Metrics [EE14, Mic13, NBH13, PL10, SKTC05, VS11, YMBA19]. Microarchitectural [BGZT22, DKD07, MSI18, MMAAK21]. Microarchitecture [CNHH15, FAR^{+}23, KCB^{+}20, WXZ^{+}21]. Microarchitecture-Based [CNHH15]. Microarchitectures [DOM^{+}07, DOM^{+}08, DPP23]. Microservices [GD18, LAX^{+}20, LLM^{+}21]. Migration [SSVS21, SD02, SLKD14]. MIMD [WCZ^{+}12]. MIMD-like [WCZ^{+}12]. Mind [AEJE17]. Minimal [FHL^{+}10, RL08]. Mining [DK16, SHJW21]. MinneSPEC [KL02]. Mirage [SQ23, SQ23]. Misprediction [SYC07]. Miss [EHdSH20]. Misses [CST^{+}04, Zha06]. Mitigate [VMP^{+}16]. Mitigating [DXSS15, KNQ15, MTT12, SBVB17]. Mitigation [CKZ^{+}20, KWL^{+}17, PHBC18, SJM17]. Mitigations [FBN^{+}24]. MNCaRT [AWD^{+}18]. Mobile [LWM20, TLG^{+}11, ZSLR14]. Modal [HXL^{+}22]. Mode [KPKK20]. Mode-Aware [KPKK20]. Model [AW15, BEA^{+}13, DL20, Eec22, EHH21, FSO^{+}22, KKL^{+}15, KWB^{+}20, LKL^{+}21, LHE^{+}21, PP12, PZX15, SGBE18, TOIS17, IPS14]. Modeling [ABC^{+}19, BS17, EHH21, FAR^{+}23, GJ21, GGS19, HEDH21, JZA^{+}18, KKP^{+}18, NBH13, PPG11, RSRT19, SQ23, VE18, WJA^{+}19, SCL06]. Models [BREM08, CVF^{+}24, CPK^{+}23, DRGA12, GO15, HBL^{+}10, LCKA23, SW16, SPHS22, SJM02]. Modern [CWK^{+}22, HA24, SFFG^{+}19, TOIS17]. Modular [GWR08, WLDN19]. Module [ALSJ09, ZNTJE23, ZZW^{+}23]. Monitoring [DEC^{+}18, GMMC15]. Monte [SCL06]. Morphable [ZL17]. Morphing [YSL^{+}21]. Most [Ano16k]. Movement [KJK21]. MPSoC [PP12]. MPSoCs [KLZ12]. MPU [VRS18, XGH^{+}22]. MPU-Sim [XGH^{+}22].

MQSim [LHCK22]. MQSim-E [LHCK22]. MRAM [ILXY18a, ILXY18b, ZBA^{+}20]. MTB [AGJ18]. MTB-Fetch [AGJ18]. Mth [MKM17]. Multi [AWD^{+}18, CVP12, EOA^{+}23, EHDH18, FJ08, HXL^{+}22, IXS18, JSLW20, JJP^{+}22, Jun17, KSO^{+}16, KFJ^{+}03, LSJ^{+}19, MNU^{+}15, MMY^{+}14, PLL08, PJ22, PL15, RL17, SMZ18, SKK22, SPAP10, VS11, YLK21, ZNTJE23, Zho06]. Multi-Accelerator [MMY^{+}14]. Multi-Architecture [AWD^{+}18]. Multi-Chip-Module [ZNTJE23]. Multi-Core [CVP12, EOA^{+}23, FJ08, IXS18, Jun17, KFJ^{+}03, MNU^{+}15, PL15, SPAP10]. Multi-Cycle [RL17]. Multi-Die [SKK22]. Multi-GPU [JSLW20, LSJ^{+}19]. Multi-Layer [KSO^{+}16]. Multi-Level [PJ22]. Multi-Modal [HXL^{+}22]. Multi-Prediction [JJP^{+}22]. multi-processors [Zho06]. Multi-Stack [YLK21]. Multi-Stage [EHDH18]. Multi-String [PLL08]. Multi-Tenant [LSJ^{+}19]. Multi-Threaded [VS11]. Multi-Threading [SMZ18]. MultiAmdahl [MMY^{+}14, ZKW12]. Multicore [ALSJ09, BEA^{+}13, CAPS09, DVAE18, DM06, KCPG18, KLSD11, Mic13, Mus09, NK22, ODKK18, SRH20, SHK15]. Multicores [AEJE17, SK21, VMS17]. Multidimensional [JSDK13]. Multikernel [WYM^{+}16]. Multilevel [CF24, PPG11]. Multimedia [ACSV02]. MultiPIM [YLK21]. Multiplication [LCKA23, PTND24]. Multiply [GG17, JPC18]. Multiprocessor [ILG10, PPG11, SLC03, XL07]. Multiprocessors [AGJ18, GKKW07, HCM10, LMJ12, MTT12, SD02, MWK^{+}06]. Multiprogram [EE14]. Multistage [Ant09, GVG^{+}08]. Multitasking [KCP^{+}19, ZCG18]. Multithreaded [BVL09, DVAE18]. Multithreading [AGJ18, KNE^{+}14, SHW19]. My [ZKW12].

Nahalal [GKKW07]. NAND [KJK21]. NAND-Based [KJK21]. Nanopore [LJ18]. Narrow [EUVG06, KRB^{+}13]. Native [MLK15]. Near [ALKSA19, FFAMK15, HCK^{+}21, HKO^{+}22, JKK^{+}21, KPEC10, LAX^{+}20, LZL^{+}20, LQYF23, XGH^{+}22, YKP^{+}22]. Near-Bank [XGH^{+}22]. Near-Data [HCK^{+}21, YKP^{+}22]. Near-Ideal [ALKSA19]. Near-Memory [LAX^{+}20]. Near-Storage [JKK^{+}21, LZL^{+}20]. Near-Threshold [KPEC10]. NearZero [Jun17]. Neda [NSF^{+}18]. Need [CVP12]. Neighbor [NSF^{+}18]. Nelder [YCH24]. NEST [JSLW20]. Nested [HBL^{+}10]. Netflix [DK13]. Network [ASK^{+}21, Ant09, CGY^{+}14, GFAHSA24, GVG^{+}08, JAM17, KPKK20, LKR21, LZS^{+}08, LHZ19, MHM^{+}24, PL15, RMA^{+}20, SCL13, WCK08, XHG^{+}19, ZL18a, EPS06, TASA13]. Network-on-Chip [CGY^{+}14, GFAHSA24, LZS^{+}08, PL15]. Network-on-Memory [RMA^{+}20]. Network-on-SSD [TASA13]. Networks [FPA^{+}21, GG17, GKK^{+}22, HA24, HXL^{+}22, KBD07, KKK13, KR18, LTL23, MXS19, MJBD11, NHKR19, RL08, RL09, RMMLK16, SPJ02, SW19, SD04, XYZ15, YHM17, YKP^{+}22, ZLM^{+}20, GD06]. Networks-on-Chip [RMMLK16]. Neumann [DC18]. Neural [FPA^{+}21, GG17, GKK^{+}22, HXL^{+}22, JAM17, KR18, LKR21, LHZ19, LTL23, MXS19, NHKR19, SW19, XHG^{+}19, YHM17, ZLM^{+}20]. Neuromorphic [BSD^{+}19]. Newest [Ano16k]. Newsletter [Ano13e]. Next [GMMC15]. NICs [LAX^{+}20, MGH^{+}22]. Nile [DEC^{+}18]. NMTSim [GLH^{+}20]. NNBench [XHG^{+}19]. NNBench-X [XHG^{+}19]. NoC [SRLP09, WL16]. NoCs [DPC16, FHL^{+}10, FD08, MCKW10, ZL18b]. NoHammer [LKP^{+}23]. Noise [CKZ^{+}20, HDAS18]. Noisy [MKMJ23]. NoM [RMA^{+}20]. Non [LKR21, PZX15, PDGV16, RM18, RSO21, SM18, VHN15, WZLQ15]. Non- [PZX15]. Non-Determinism [RSO21]. Non-Intrusive [PDGV16]. Non-Volatile [LKR21, RM18, SM18, VHN15, WZLQ15]. Novel [XL07]. NUMA [SJA^{+}17]. NVM [CYAW20, HH22, KKLL22, MV15, PDGV16]. NVMain [PZX15]. NVMe [ZKH^{+}20]. NVMM [OK22].

O [JLS^{+}23, KLWJ21, LKA15, LKKS15, MAHK18, SYC14]. O-Intensive [JLS^{+}23]. Obfuscation [CYAW20]. Oblivious [SCL13, TD02]. Odd [SCL13]. ODIN [SRLM20]. Off [GKK^{+}22, RSO21]. Offer [Ano10e]. Offlining [LKK19]. offs [BSD^{+}19]. On-Chip [GGM^{+}16, KBD07, KKK13, KDS22, KLZ12, LGLK17, MJBD11, ZM07, WCK08]. On-Demand [MHAD15]. Once [MSE^{+}17]. Online [ZCG18]. Open [AWD^{+}18, Ano13h, Ano13i, ACG^{+}07, ILG10, OK22, SKK22]. Open-Source [AWD^{+}18, ILG10, OK22, SKK22]. OpenMDS [SKK22]. Operand [BHD09, MSI18]. Operating [AEJE17]. Operation [KCPG18, RAD^{+}23]. Operations [JPC18]. Opportunities [RCK21, TNC19, Wu14]. Opportunity [MTH11]. Optane [MDSG20]. Optical [CGY^{+}14]. Optimal [BHY^{+}19, CFM^{+}03, NMS14]. Optimization [ALKSA19, BHY^{+}19, CNHH15, GO15, KDS22, LAS22, LLM^{+}21, MMY^{+}14, NR21, WCC14, YMG14, GD06]. Optimizations [BY17, WZLQ15, ZM07]. Optimized [EOA^{+}23]. Optimizing [MSE^{+}17]. ORAM [RM18]. Orbital [DL19]. Orchestrating [ASK^{+}21]. Order [CTJ^{+}17, DV13, Eec22, EHdSH20, HEDH21, IXS19, PGJ12, TOIS17, CMLV03]. Ordering [HR10]. Organization [AGK21, BSBD^{+}08, GKKW07]. OS-Level [LKK19]. Our [FAR^{+}23, Ano12k].

Out-Of-Order [DV13, IXS19, CTJ^{+}17, HEDH21, TOIS17, CMLV03]. Outcome [CSSU20]. Outputs [WTSW21]. Overall [LX08]. Overcoming [HCK22]. Overhead [BGL^{+}23, FBN^{+}24, RM18, SRS20]. Overheads [KQGS16, SHK15, ZKH^{+}20]. Overview [FUWT12].

Packet [KPKK20, MJBD11]. Packets [FHL^{+}10]. Page [LMK06, TV02, WMZY17, ZKH^{+}20]. Page-based [LMK06]. Page-Level [TV02]. Pages [JLA16]. Paging [HBL^{+}10]. Pairwise [GBS^{+}20]. Pairwise-Correlating [GBS^{+}20]. Paradigm [MKD^{+}23, TASA13]. Parallel [ADS^{+}19, AKK16, CLCG14, EHH21, KLZ12, KPEC10, LX08, MPPS17, XL07, AD06]. Parallel/Distributed [AKK16]. Parallelism [CF24, KHS^{+}24, yPSS^{+}10, TMSA16, VE18]. Parallelization [DM06]. Parity [JSDK13]. ParMiBench [ILG10]. Partially [RL08]. Partially-Minimal [RL08]. Partitioned [JKK^{+}21]. Partitioning [JLRA18, MCRV07]. Party [OSH16]. Pass [ZL18b]. Passing [GGM^{+}16]. PATer [LCW^{+}16]. Path [GKK^{+}22, TOIS17]. Paths [RL17]. Pattern [Ano14c, Ano14d, CYAW20, SRS20, SHJW21, WSVS22]. Patterns [LPK16, WTSW21]. PCIe [LSJ^{+}19]. PCM [KL18, WMZY17, YYK^{+}18]. PCM-Based [KL18]. pd [AKK16]. pd-gem5 [AKK16]. Per-Core [LMT^{+}09, SW16]. Per-task [LJM^{+}14]. Performance [AW15, ABC^{+}19, BSD^{+}19, BREM08, CCWY17, CZYY11, CLCG14, CTL^{+}20, CFM^{+}03, DPC16, DVAE18, EE14, FHL^{+}10, GMMC15, GGS19, GF16, GSG^{+}17, JSDK13, KKL^{+}15, KKP^{+}18, KH18, KWB^{+}20, LKL^{+}21, LTL23, MA19, MTH11, MDSG20, MWK^{+}06, NK22, PLK^{+}23, PL10, RMMLK16, RCS15, RB14, SKK22, SJM02, SJA^{+}17, SCL06, SHJW21, TASA13, VP16, WCZ^{+}12, WMJM23, YMBA19, YNS^{+}08, YP23, ZVYW03, ZNTJE23, ZCG18, ZL18b, ZWT22, LHWB10, SYC14, Zho06]. Performance-Efficient [ZL18b]. Performance-Energy [KKL^{+}15]. Periodic [GJ21]. Peripheral [AMW15]. Permanent [OKS^{+}15, SKS^{+}15]. Persistence [KQGS16, MAT17, PDGV16]. Persistent [KQGS16, MDSG20, WYL^{+}15]. Personalized [KLR24]. Perspective [ILNS20, ZWT22]. Petabyte [Jac16a]. PetaFLOP [Jac16a]. Phase [Jun17, KJS^{+}19, KHS^{+}24, KMJ18, KKL^{+}07, Sez10]. Phase-Change [KJS^{+}19]. Phase-Level [KHS^{+}24]. Physical [KWB^{+}20, Rot08]. Picture [WXZ^{+}21]. PID [RCS15]. PID-Controlled [RCS15]. PIM [CPK^{+}23, NSC20, RCK21, SHJW21, ZTRA22]. PIM-Based [NSC20]. PIM-GraphSCC [NSC20]. PIMSim [XCW^{+}19]. Pipeline [AS18, CTL^{+}20, MSA19, PL15]. Pipelined [PLL08]. Pipelining [FUWT12]. Pitfalls [SQ23]. Placement [CKA20, HCM10, LLPC19]. Plane [TMSA16]. Plane-Level [TMSA16]. Platform [EGWM14]. Platforms [GO15, KDS22]. Point [ACSV02, DKD07, GF16, HLR21, LTL23]. Pointer [KKL^{+}23, MAT17, RADZ19]. Pointer-Based [MAT17]. Points [AEJE17]. Policies [NPBS23]. Policy [JTG23, LLKS12, TMSA16, VGMSLN^{+}18]. Portable [LJ18]. Post [KCB^{+}20]. Post-Silicon [KCB^{+}20]. Potential [CPK^{+}23, LLKS12]. Power [AEJE17, ÇTNL16, CVP12, CGY^{+}14, CLJ^{+}02, DRGA12, DL20, FHL^{+}10, GFAHSA24, KWL13, KPKK20, KWKK18, KG10, KFJ^{+}03, LKK19, LMT^{+}09, LLS^{+}15, PGR^{+}23, PHBC18, PP12, SKA^{+}20, SBVB17, SKTC05, SW16, SSS^{+}21, SPJ02, SCF04, SFFG^{+}19, TVB^{+}13, UKM02, WCK08, YHM17, YPFP14, ZAK^{+}17, ZL18b, LHWB10, MWK^{+}06]. Power-Aware [DL20, UKM02]. Power-Efficient [YHM17, SPJ02]. Power-Gating [ÇTNL16, ZL18b]. Power-Limited [AEJE17]. POWER8 [LCW^{+}16]. pPIM [SCB^{+}20]. PPT [ABC^{+}19]. PQC [LAM^{+}22]. Practical [DPP23, DM06]. PRAM [JP13]. Pre [BGZT22, CSSU20, HKO^{+}22, MKSP05, VRFT24, WB14]. Pre-Alignment [HKO^{+}22]. Pre-Executed [MKSP05, WB14]. Pre-Execution [CSSU20]. Pre-RTL [VRFT24]. Pre-Silicon [BGZT22]. Precise [NFAE19]. Precision [NHKR19, SCB^{+}20]. Precision-Scaling [SCB^{+}20]. Predication [JMKP07, JMKP08]. Predictability [MXS19]. Predicting [PB16]. Prediction [CST^{+}04, DVAE18, EF07, GAH^{+}23, JJP^{+}22, MHAD15, PGJ12, PB16, PS17, SJ22, SYC07, SLKD14, SW19, YHY^{+}22, YP23, ZCG18]. Predictive [WCYC09]. predictor [RZ06]. Predictors [ST20, SYC07]. Prefetch [PB16]. Prefetcher [BLKSA17, RJ20, YYK^{+}18]. Prefetchers [JE22, PB16]. Prefetching [AGJ18, CSSU20, GBS^{+}20, ILNS20, JTG23, KP21, LCW^{+}16, PGJ12, TLG^{+}11, ZMC17]. PreGNN [GKK^{+}22]. Preprocessing [GKK^{+}22, KLR24, YYK^{+}18]. Presence [JTG23]. Preserving [MTM18]. Pressure [HCM10]. Pressure-Aware [HCM10]. Prevent [BBZ^{+}19, LLSA18]. Preventing [LKP^{+}23]. Primate [MHM^{+}24]. Priority [LSJ^{+}19, SK21]. Priority-Based [LSJ^{+}19]. Privacy [MS16, MTM18]. Privacy-Preserving [MTM18]. Proactive [FJ08]. Probabilistic [EF07, RZ06]. Probability [IKW^{+}20]. Probability-Based [IKW^{+}20]. Problem [HS04]. Proceedings [Ano10g]. Process [DOM^{+}07, DOM^{+}08, MTT12, Mus09, ZZJ18]. Process-in-Memory [ZZJ18]. Process-Variation [Mus09]. Processing [AG17, AA19, AWD^{+}18, BHL^{+}18, BGP^{+}17, CTJ^{+}17, CHK^{+}18, FSO^{+}22, FFAMK15, HCK^{+}21, HKO^{+}22, JPC18, KPKK20, KZL18, LQYF23, NSC20, SRV^{+}19, SRS20, ST20, SPHS22, XGH^{+}22, XCW^{+}19, YLK21, YKP^{+}22, ZZW^{+}22, ZTRA22]. Processing-in [JPC18]. Processing-in-DRAM [ZTRA22]. Processing-In-Memory [YLK21, BGP^{+}17, SPHS22, XCW^{+}19]. Processor [BDBS^{+}08, CZYY11, KPEC10, KFJ^{+}03, LCW^{+}16, LJ04, MKSP05, SCB^{+}20, VE18, YKMG15]. Processor-in-Memory [SCB^{+}20]. Processors [ADS^{+}19, ASK^{+}21, ACSV02, CXS18, FJ08, GGS19, GMM^{+}19, HEDH21, LLPC19, LMC^{+}09, MHM^{+}24, Mus09, NK22, PGJ12, RADZ19, RYSN04, SRH20, SMY15, TOIS17, TDO16, VS11, WCYC09, WB14, CMLV03, Zho06]. Production [MKD^{+}23]. Profiles [CNHH15]. Profiling [CV15, GMMC15, SFFG^{+}19]. Program [KKL^{+}07, NGS15, SSTS17, SHK15]. Programmable [DCG12, DEC^{+}18, SCB^{+}20]. Programming [KLKK14, MKD^{+}23]. Programs [GRCV02, MPPS17, ORS^{+}06]. Progress [NB24]. Progressive [AG17]. Promotion [JTG23, MDK^{+}23]. Proposed [BGS^{+}20]. Protection [OK22, Vol21]. Protocol [KSB19]. Providing [KKH14]. PRR [SKD09]. Pruned [WTSW21]. Pruning [AQ24]. Publication [Ano11j]. Publishing [Ano12c, Ano13h, Ano13i]. Pulley [LAS22]. Purpose [DPP23, WSVS22].

Q [GMM^{+}19]. Q-Learning [GMM^{+}19]. QAOA [UTT^{+}24]. Quality [YC15]. Quantifying [TND^{+}21]. Quantitative [KPPK21, LPK16]. Quantization [SBQK21]. Quantum [AS18, AQ24, EOA^{+}23, RTKQ21, RK22, TXD^{+}23, ZB19]. Quantum-Dot [AS18]. Quasi [JDK^{+}02]. Quasi-Static [JDK^{+}02]. Qubit [EOA^{+}23]. quick [Ano12k].

Race [EGWM14]. Racetrack

[HKO^{+}22, KHB^{+}19]. Radar [WLL^{+}22].

Radix [SD04, SCL13]. RAM [JP13, MVJ17, SSVS21, YPFP14]. RAMBO [LLM^{+}21]. Ramulator [KYM16]. Random [RL09]. Randomization [BGS^{+}20]. Randomized [KYP21, RL08]. Ransomware [MPA^{+}18]. Rapid [DVAE18, SRS11]. RAS [GLJ^{+}21, RCS15]. Rate [PL10]. Rate-Based [PL10]. ray [NBW^{+}23]. Re [RASW19]. Re-Routing [RASW19]. Reach [AQ24]. Read [MVJ17, MSE^{+}17, ZZJ18]. Read-Disturbance [MVJ17]. Read-Once [MSE^{+}17]. Real [PPG^{+}17, RSO21]. Real-Time [PPG^{+}17, RSO21]. Rebasing [ILNS20]. Rebuttal [BREM08]. Recommendation [KLR24, LQYF23]. Reconfigurable [LAX^{+}20, LLD^{+}18, LYL^{+}16, SSSM18, TNC19, WLL^{+}22, ZL18a]. Recovery [EHdSH20, MPA^{+}18, MAT17]. Redistribution [LHPR23]. ReDRAM [SSSM18]. Reduce [Cit04, KG10]. Reducing [FHL^{+}10, FBN^{+}24, KWL13, KQGS16, KJK21, Zha06]. Reduction [AYL22, HLH16, KKKH18, KPL^{+}21, KFJ^{+}03, Per21, SCF04, UTT^{+}24]. Redundancy [GWR08]. Redundant [WLZZ23]. Refactored [LKA15]. reference [Rot08]. Refresh [KKKH18, LLSA18]. Refuting [BGS^{+}20]. Regional [YJZ15]. Register [BSBD^{+}08, EE16, JEAG^{+}19, Rot08]. Registers [BHD09]. Regression [YYK^{+}18]. Regulation [HPS23]. Reinforcement [JE22, KHS^{+}24]. Relaxed [PTND24]. Reliability [AQ24, ÇE14, DD18, HSUS11, SMY15, TMNK19]. Reliable [KMJ18, KKL^{+}07]. Relocation [SKD09]. Remapping [WMZY17]. Remote [KSB19]. Removing [ZKH^{+}20]. Reorder [ASSK21]. Reordering [MNFI20, SJM02, ZZW^{+}22].

Replacement [NPBS23, VGMSLN^{+}18]. Replication [Vol21]. Replication-Aware [Vol21]. Reporting [SRS20]. Representation [NGS15, WXZ^{+}21]. Request [SJM02]. ReRAM [LHZ19]. ReRAM-Based [LHZ19]. ReRAMs [ZZJ18]. Resampling [PL10]. Research [AWD^{+}18, KL02]. Reservation [LZS^{+}08]. Resilience [GLJ^{+}21, LBB^{+}19, OKS^{+}15, SKS^{+}15, SHK15]. Resiliency [LLS^{+}15]. Resilient [ODKK18]. Resistive [MLA^{+}14, YKMG15, YWG17, ZL17]. Resource [KCP^{+}19, KQD18, LZS^{+}08, LLM^{+}21, ODKK18, RMMLK16, SJ22, CMLV03]. resource-conscious [CMLV03]. Resource-Scalable [SJ22]. Response [EHH21, FHL^{+}10]. Restating [EE14]. Results [ACSV02, MKSP05, WB14]. RETROFIT [ZKF^{+}18]. Reusable [JLKK23]. Reuse [BY17, CMP^{+}14, LPK16, YHM17]. Reusing [MKSP05]. Revenues [DOM^{+}07, DOM^{+}08]. Reviewers [Ano11b, Ano12b, Ano13b]. Revisiting [WB14, ZWT22]. Rich [LBB^{+}19]. Ring [LAM^{+}22]. Ring-LWE [LAM^{+}22]. RIO [HEDH21]. RISC [ZBA^{+}20]. RISC-V [ZBA^{+}20]. ROB [HEDH21]. ROB-Centric [HEDH21]. Robotics [KHS^{+}24]. Rock [Ano15h, Ano15i]. Rollback [MAT17]. Rollback-Recovery [MAT17]. Roofline [IPS14]. Router [KWL13, PL15, SRLP09, ZL18b]. RouteReplies [ZZW^{+}23]. Routing [FD08, GDF^{+}04, GF16, KK21, KL18, MCKW10, RL08, RL09, RASW19, SDTG04, SCF04, SCL13, TD02, XYZ15]. Row [KNQ15, KLCA21, LLSA18, LKP^{+}23, SJM17]. Row-Hammering [LLSA18]. Row-Streaming [KLCA21]. Rowhammer [FBN^{+}24, ZTRA22]. RPCs [LAX^{+}20]. RPPM [DVAE18]. RTL [VRFT24]. RTSim [KHB^{+}19]. Run [KNGK15, LX08, RADZ19]. Run-Time [KNGK15, RADZ19, LX08]. Runahead [GBS^{+}20, MKSP05, NFAE19, WB14]. Runtime [GMMC15, MPPS17, MKD^{+}23, MXS19, SPHS22, ZB19]. Runtime-Assisted [MPPS17]. Rusty [MXS19].

s [Jac16a]. SA [SHW19]. Safe

[MLK15, MKP^{+}24]. Safer [FAR^{+}23]. Safety [ODKK18]. Safety-Critical [ODKK18]. SALAD [SCR^{+}17]. Sampled [LJ04]. Scalability [VP16, MWK^{+}06]. Scalable [APK^{+}18, ABC^{+}19, FSO^{+}22, GWR08, JJP^{+}22, KKL20, KKJ^{+}22, MGH^{+}22, MCY^{+}12, RSRT19, SRV^{+}19, SJ22, SSS^{+}21, TASA13, ZL18b]. Scale [AG17, DRGA12, DSVK12, GLJ^{+}21, HCM10, LKR21, LHE^{+}21, MTH11, MKD^{+}23]. Scale-Model [LHE^{+}21]. Scale-Out [MKD^{+}23]. Scaled [ILXY18a, ILXY18b, KCPG18]. ScaleGPU [KLKK14]. Scaling [ÇTNL16, DL20, GO15, MLM^{+}06, MKP^{+}24, SPJ02, SCF04, SCB^{+}20, YC15]. SCC [CLCG14]. SCEPTER [DPC16]. Sched [NPS21]. Scheduling [APK^{+}21, BRUS21, CCWY17, DK16, DC18, LLKS12, LKKS15, LA16, LSJ^{+}19, MNU^{+}15, SBVB17, SK21]. Scheme [CLCG14, MMR17, SLC03, WJFH11]. Scientists [GMPMC^{+}23]. Second [LMJ12]. Section [MNU^{+}15]. Section-Aware [MNU^{+}15]. Secure [KZY^{+}19, NK22, ODKK18, Sez10, SM18, TWI^{+}24]. Security [BGZT22, BGS^{+}20, DK16, HPS23, HSUS11, HBW^{+}23, KZY^{+}19, NK22, OSH16, PGC22, Ano16d]. Selection [NR21]. Selective [DV13, MVJ17]. Self [LHPR23, ZTRA22]. Self-Attention [LHPR23]. Self-Tracking [ZTRA22]. semantics [BLM06]. Semi [MAHK18]. Semi-Coherent [MAHK18]. SEMS [KKJ^{+}22]. Sensing [HDAS18]. Sensitive [RYSN04]. Sensitivity [NPBS23].

Sequence [VRS18]. Sequencing [LJ18]. sequential [ORS^{+}06]. Serial [JAM17, WSVS22]. Server [ADS^{+}19, ASK^{+}21, TMNK19, WLL17]. Servers [JSLW20, KPKK20, PHBC18]. Service [KYP21, NK22, YC15]. Service-Aware [YC15]. Services [Ano10g, Ano12c]. Set [GIH^{+}24, MMR17, RAD^{+}23, YJZ15]. Set-Granular [YJZ15]. Shader [WCYC09]. Shaping [JLRA18]. Shared [CZYY11, FJ08, IXS18, SLKD14, SRLP09]. Shared-Buffer [SRLP09]. Shared-Memory [IXS18]. Sharing [GG17, KCP^{+}19, LMJ12, RMMLK16, WYM^{+}16]. Shell [SKK22]. Shifting [TVB^{+}13]. Shimmer [TMNK19]. Shootdown [PHBC18]. Short [GWR08, ZZJ18]. Should [ZKW12]. Shrink [LWB13]. Shrink-Fit [LWB13]. Shuffle [WCZ^{+}12]. Shutdown [WCYC09]. Side [MKMJ23]. Side-Channels [MKMJ23]. Sifting [AEJE17]. Silicon [BGZT22, CMP^{+}14, DXSS15, FBN^{+}24, KCB^{+}20]. Sim [XGH^{+}22]. SIMD [WCZ^{+}12]. Simple [NPS21]. SimpleSSD [JZA^{+}18]. Simplified [BSMB23]. SIMT [LPK16]. Simulating [FAR^{+}23]. Simulation [AKK16, ACG^{+}07, DM06, FAR^{+}23, Hos18, JZA^{+}18, KL02, LHZ19, LHE^{+}21, LJ04, MMAAK21, VRFT24, SCL06]. Simulation-Based [KL02]. Simulator [Ano10a, FLSZ17, GLH^{+}20, JC17, KHB^{+}19, KYM16, LHCK22, LYR^{+}20, MGHP20, OK22, PZX15, PHO^{+}15, RCBJ11, XGH^{+}22, XCW^{+}19, YLK21]. Simulators [BVL09, CAPS09, EHH21]. Simultaneous [SHW19, WYM^{+}16]. Single [BEA^{+}13, KKL^{+}15, KH18, KFJ^{+}03, MNU^{+}15, MJBD11, SD02]. Single-Cycle [MJBD11]. Single-ISA [KFJ^{+}03, MNU^{+}15]. Single-Thread [KH18]. Singular [LCKA23]. Singular-Value [LCKA23]. Situ [MNFI20]. Size [NMS14]. Sizing [LWB13]. Sky [KWB^{+}20]. Slicing [YSL^{+}21]. Slowdown [SJ22, ZCG18]. Small [JLA16, NB24]. SmaQ [SBQK21]. Smart [MGH^{+}22, SBQK21]. SmartIndex [WMJM23]. SmartSSD [LZL^{+}20]. SMT [HR10, KH18, RYSN04, SHW19, TVB^{+}13]. SMT-Directory [HR10]. SMT-SA [SHW19]. SoC [HBW^{+}23, MMY^{+}14]. Society [Ano09a, Ano10c, Ano11i, Ano08c, Ano09l, Ano09m, Ano10f, Ano10l, Ano10a, Ano10n, Ano10m, Ano11i, Ano12j, Ano13j, Ano14e, Ano14f, Ano15f, Ano15g]. Socket [PGC22, SPAP10]. SoCs [APK^{+}21, BHY^{+}19, SKA^{+}20]. SoCurity [HBW^{+}23]. Soft [EE16, EUVG06, KRB^{+}13, MHM^{+}24, PL15, SG14, YE07]. Software [BKA^{+}09, CTJ^{+}17, CV15, FAR^{+}23, GAH^{+}23, LMK06, MKM17, TVB^{+}13, XWG^{+}14]. Solid [JZA^{+}18, KKL20, SYC14, YNS^{+}08]. Solid-State [KKL20, SYC14]. Solution [SAA^{+}23]. Sorting [LCHL20, LAS22]. Source [AWD^{+}18, ILG10, OK22, SKK22]. Space [DL19, LLPC19]. Sparing [MCM13]. Sparse [LWM20, LZD^{+}23, YG18]. Sparsity [MNFI20, PTND24]. Spatial [SW19, WLDN19, ZCG18]. SPEC [KL02]. Special [Ano10e]. Specialization [NGS15]. Specialized [ST20]. Specific [BSD^{+}19, WCC14]. Spectre [LG20]. Speculation [ASSK21, MGI14, RL17, XJ09, YHY^{+}22]. Speculative [ASSK21, BBZ^{+}19, GQLZ19, PJ22, Per21, SLC03]. Speed [MTT12, MCRV07, ZL18a]. Speeding [RASW19]. Speedup [LJ04]. Spintronics [CHK^{+}18, LJ18]. Spintronics-Based [LJ18]. SPMD [GG11]. SRAM [GIH^{+}24, MV15, SRS20]. SSD [APK^{+}18, BGL^{+}23, KKP^{+}18, KJK21, LGLK17, LZL^{+}20, LKL^{+}21, LHCK22, MPA^{+}18, PLK^{+}23, TASA13]. SSDs [IKW^{+}20, KLWJ21, LKKS15, yPSS^{+}10, TASA13, TMSA16]. SSE [NK22]. Stack [HCK22, KLWJ21, LEBM20, MKD^{+}23, YLK21, ZKH^{+}20]. Stacked [SFFG^{+}19]. Stacking [HRF^{+}11]. Stacks [EHDH18, KNGK15]. Stage [EHDH18, NBH13]. Stars [Ano15h, Ano15i]. State [JZA^{+}18, KKL20, SYC14, YNS^{+}08, ZAK^{+}17]. Static [FUWT12, JDK^{+}02]. Statistical [JLRA18, KCPG18, MS16]. Stay [Ano10f, Ano13j, GBK^{+}09]. Steroids [JLA16]. Stochastic [KNG^{+}18, Man15]. STONNE [MMAAK21]. Storage [AGK21, DSVK12, JKK^{+}21, LKA15, LZL^{+}20, PS17, WZLQ15, ZKH^{+}20]. Storage-Effective [AGK21]. Storage-Free [PS17]. Store [LHWB10, PJ22]. Stores [HS04, KKLL22]. Strategy [HCM10, NPS21]. Stream [KPEC10, MM03, WLN22]. Stream-Based [MM03]. Streaming [CSSU20, KLCA21, KLZ12]. Strength [Per21, RK22]. Strict [SJM02]. Stride [YHY^{+}22]. String [PLL08]. Stripes [JAM17]. Structure [CLJ^{+}02, KLCA21, NBW^{+}23, SSTS17, SJM17]. Structured [AYL22, LWM20, PTND24]. Structured-Sparse [LWM20]. Structures [NSC20]. STT [ILXY18a, ILXY18b, JP13, MVJ17, SSVS21, YPFP14, ZBA^{+}20]. STT-MRAM [ILXY18a, ILXY18b, ZBA^{+}20]. STT-RAM [YPFP14]. STT-RAM-Based [SSVS21]. Student [Ano10e]. Studies [DSVK12]. Study [CKA20, KKLL22, NMS14]. Studying [ZTS16]. Sub [GGM^{+}16]. Sub-System [GGM^{+}16]. subarray [Zha06]. Subtleties [BLM06, KNGK15]. Sufficient [XYZ15]. Suite [MTM18, WLL17]. Super [WMZY17]. Super-Dense [WMZY17]. Superscalar [VE18]. Supply [KLZ12, MTT12]. Support [DKD07, GMMC15, KNQ15, MLK15, MKM17, MSI18, MKP^{+}24, SPHS22]. Supporting [GIH^{+}24, NSF^{+}18, PTND24, SRS11].

Survive [MAT17]. Sustainability

[Eec22, ZNTJE23]. Sustainable [Wu14]. SVSoC [GQLZ19]. SW [DCG12]. Switch [DC18, TVB^{+}13]. Switched [JLP07]. Switching [SRH20]. Symmetric [SCR^{+}17]. Synchronization [LLD^{+}18, SLC03]. Synchronous [LKKS15]. Synctium [KPEC10]. System [GGM^{+}16, Jac16a, JZA^{+}18, KNG^{+}18, KL18, KKJ^{+}22, KSO^{+}16, KLZ12, KR18, LQYF23, MXS19, RCBJ11, SJA^{+}17, WLL^{+}22, XL07, ZLS10, ZBA^{+}20, LLLM06]. System-on-a-Chip [XL07]. Systems [AKK16, BDBS^{+}08, CWK^{+}22, CKA20, CLJ^{+}02, CEA18, GQLZ19, GJ21, GRCV02, HBL^{+}10, HH22, ILG10, KJS^{+}19, KLR24, LLJK23, LBB^{+}19, LSJ^{+}19, LJM^{+}14, MGH^{+}22, MAHK18, PPG^{+}17, PL15, PZX15, PPG11, SLC03, SPAP10, TLG^{+}11, ZVYW03, LHWB10]. Systems-on-a-Chip [GQLZ19]. Systolic [KLCA21, LWM20, SHW19].

T [LLJK23]. T-CAT [LLJK23]. Table [Ano14g, Ano14h, Ano15j, Ano15k, Ano16p, Ano16n, Ano16o, KL18, Ano12h]. Tackling [RCS15]. tag [YE07]. tag-match [YE07]. Tagging [KKL^{+}23]. Tail [RSO21]. Tails [SSTS17]. Take [GKK^{+}22]. Tale [FAR^{+}23]. Tangible [SMLS15]. Targeting [KDS22]. Task [KLZ12, MPPS17, SK21, LJM^{+}14]. Task-Based [MPPS17]. Task-Parallel [KLZ12]. TB [Jac16a]. TB/s [Jac16a]. TBM [TMSA16]. Technique [AMW15, ILXY18a, KRB^{+}13, MV15, Mus09, WCYC09]. Techniques [JDK^{+}02, PL10, SSVS21]. Technology [GLH^{+}20]. Temperature [UTT^{+}24]. Temporal [BLKSA17, EF07]. Tenant [LSJ^{+}19]. Tensor [KPL^{+}21, LWM20]. TERMinator [MTM18]. There [Ano12k]. Thermal [CFM^{+}03, LYR^{+}20, Mic20, SRS11, Wu14]. Thermal-Capable [LYR^{+}20]. Thermally [XYMY16]. Thinking [Ano16k]. Third [OSH16]. Third-Party [OSH16]. Thread [CCWY17, GBK^{+}09, KKL^{+}15, KH18, MNU^{+}15, MGI14, MKP^{+}24, NPS21, RYSN04, SLKD14]. Thread-Level [MGI14]. Thread-Safe [MKP^{+}24]. Thread-Sensitive [RYSN04]. Threaded [VS11]. Threading [SMZ18]. Threads [HLH16, MKM17, ORS^{+}06]. Threats [CWK^{+}22]. Three [RL08]. Three-Dimensional [RL08]. Threshold [KPEC10]. Throttling [UKM02]. Throughput [ILXY18a, ILXY18b, KKK13, LLPC19, MSA19, Mic13, SRLP09, SCL13]. Tiered [CKA20, LLJK23, MDK^{+}23]. Tile [Mus09, YSL^{+}21, CZYY11]. Tile-Based [Mus09]. Tiled [LCKA23]. Tilera [CZYY11]. Time [CVF^{+}24, KNGK15, LLSA18, PPG^{+}17, RADZ19, RSO21, LX08]. Timing [BGS^{+}20, EHH21, RL17, XJ09]. TLB [CLJ^{+}02, PHBC18]. Toggle [PBO^{+}15]. Toggle-Aware [PBO^{+}15]. TokenSmart [SSS^{+}21]. Tolerance [EUVG06, Zho06]. Tolerant [GDF^{+}04, HRF^{+}11, PL15]. Toolchain [VRFT24]. Topology [GD06, KBD07]. Tori [GDF^{+}04, SDTG04]. Torus [RL09]. Trace [MM03]. Traces [PV06]. Tracking [ZTRA22]. Trade [BSD^{+}19]. Trade-offs [BSD^{+}19]. Tradeoff [SHK15]. Traffic [HLH16, JLRA18, TD02, ZLAE17]. Train [CVF^{+}24]. Train-Time [CVF^{+}24]. Training [JKK^{+}21, KR18, LKR21, LHZ19, MGH^{+}22, SBQK21, YP23]. Transaction [GLH^{+}20, LZS^{+}08]. Transaction-Aware [LZS^{+}08]. Transaction-Command [GLH^{+}20]. Transactional [BKA^{+}09, DD18, LLD^{+}18, WJFH11, WYL^{+}15, XWG^{+}14, BLM06]. Transactions [Ano10b, Ano14c, Ano14d, NB24, Ano12k, Ano13e]. Transcending [CTJ^{+}17]. Transfer [RMA^{+}20]. Transformation [KKKH18, VD02]. Transformer [CPK^{+}23, LCKA23].

Transformer-Based [CPK^{+}23, LCKA23]. Transient [BBZ^{+}19]. Transients [Mic20]. Translation [IKW^{+}20, LMJ12, LLLM06]. Translation-Lookaside [LMJ12]. Transparent [KKH14, WLN22]. Tree [JLKK23, PJ22, SJM17, TWI^{+}24, Ant09, GVG^{+}08]. Tree-Based [PJ22]. Trees [SB18]. TRiM [KPL^{+}21]. Trojan [CWK^{+}22]. Trusted [TXD^{+}23]. TrustZone [PPG^{+}17]. TrustZone-Assisted [PPG^{+}17]. Tulip [GFAHSA24]. TUNE [CXS18]. Tuner [LCW^{+}16]. Tuning [CXS18, YMBA19]. Turbo [Mic20, VMS17]. Turn [GFAHSA24]. Turn-Free [GFAHSA24]. TWiCe [LLSA18]. Twin [TMSA16]. TWL [KJS^{+}19]. Two [LCHL20]. Two-Directional [LCHL20].

Ultra [MTT12]. Ultra-low [MTT12]. Unaware [KLKK14]. Understanding [HXL^{+}22, LKR21, WYY^{+}23, XHG^{+}19, YCD^{+}20, YZY^{+}22]. Unexpected [MDSG20]. Unfairness [SJA^{+}17]. Unidirectional [Ant09, GVG^{+}08]. Unification [RB14]. Unified [LHZ19]. UNISIM [ACG^{+}07]. Unit [DCG12, GG17]. Units [GMMC15, JLRA18, MTT12]. Unleashing [CPK^{+}23]. Untitled [Ska09b, Ska10b, Ska11b]. Unused [KG10]. updates [RZ06]. Upgrading [IPS14]. Ups [MCRV07]. Use [FJ08, YSL^{+}21]. Useful [GMPMC^{+}23]. Usefulness [PB16]. User [MLM^{+}06, PZX15, LLLM06]. User-Driven [MLM^{+}06]. User-Friendly [PZX15]. user-level [LLLM06]. Using [AG17, BHY^{+}19, CHK^{+}18, GGS19, GO15, HKO^{+}22, KDL23, KKP^{+}18, KCP^{+}19, KLCA21, KH18, LMT^{+}09, LLM^{+}21, LJ04, MCY^{+}12, NSC20, PL10, RADZ19, RAD^{+}23, SK21, WLWZ19, WB14, WTSW21, YE07, YHM17, BREM08, JDK^{+}02, MTT12, SLC03, SCL06, Zho06]. Utilization [MA19, TMSA16].

V [ZBA^{+}20]. Validation [GWR08, VRFT24]. Valley [GBK^{+}09]. Value [AS14, CST^{+}04, KKKH18, KKLL22, LCKA23, PLK^{+}23, SBQK21, SW19, SRLM20, YHY^{+}22]. Value-Aware [AS14]. Values [EUVG06, KRB^{+}13]. Variability [DRGA12, RCS15]. Variable [RK22]. Variation [MTT12, Mus09]. Variations [DOM^{+}07, DOM^{+}08]. Variety [AG17]. vCache [KKH14]. Vector [GIH^{+}24]. Vehicles [APK^{+}21]. Verification [ZLS10]. Versatile [LTL23, WZLQ15]. Vertical [HRF^{+}11, ILXY18b]. Via [KFJ^{+}03, BY17, CCWY17, KHS^{+}24, LYY^{+}21, YMBA19]. Victimization [ZKH^{+}20]. View [KKH14]. Virtual [ÇE14, GIH^{+}24, KNGK15, PHBC18]. Virtualization [SYC14]. Virtualized [HBL^{+}10, KKH14]. Vision [GQLZ19]. Visual [KWB^{+}20]. VLIW [Jac16b]. VMOR [MSI18]. Vol [Ano15a, Ano16a, Ano16b, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22]. Volatile [LKR21, PZX15, RM18, SM18, VHN15, WZLQ15]. Voltage [ÇTNL16, CKZ^{+}20, HDAS18, KCPG18, MTT12, SPJ02, SCF04, YC15]. Voltages [MTT12]. vs [GBK^{+}09]. Vulnerabilities [BGZT22, HSUS11, KWKK18, KZY^{+}19].

Wall [HCK22]. Warehouse [AG17, MTH11]. Warehouse-Scale [AG17, MTH11]. Warp [ZTS16]. Way [FAR^{+}23, ZVYW03, Ano12k]. Way-Halting [ZVYW03]. Wear [KJS^{+}19, LZLX15, ZKF^{+}18]. Wear-Leveling [LZLX15]. Wearables [Ano15i]. Web [MGI14, VP16, ZSLR14]. Webpage [ZSLR14]. Weight [CVF^{+}24, GG17, IXS19]. Weight-Sharing [GG17]. Weighted [EE14, RL09]. Weighted-IPC [EE14]. Whole [WXZ^{+}21]. Whole-Picture [WXZ^{+}21]. Window [LLSA18]. Wire [Cit04, ZL18a]. Wire-Speed [ZL18a]. Wires [TNC19].

Word [VD02]. Word-Interleaving [VD02]. Words [KG10]. Workers [VP16]. Workflows [TMSX23]. Worklist [ZMC17]. Worklist-Directed [ZMC17]. Workload [BSD^{+}19, EE14, KL02, SRH20, WXZ^{+}21]. Workload-Specific [BSD^{+}19]. Workloads [BHL^{+}18, DS09, JLS^{+}23, LMT^{+}09, PKKK23, PLK^{+}23, XHG^{+}19, ZAK^{+}17]. Worst [SKTC05, SCL13, TD02]. Worst-Case [SCL13, TD02]. WPC [WXZ^{+}21]. Write [ILXY18a, LKKS15, WMZY17]. Writes [ILXY18b].

X [AGK21, NBW^{+}23, XHG^{+}19]. X-ray [NBW^{+}23]. Xilinx [SKK22]. XML [BVL09]. XML-Based [BVL09].

Years [Ano10b]. Yourself [Ano10d].

Zebra [KKKH18]. Zero [BGL^{+}23, KKKH18, KLWJ21]. Zero-Aware [KKKH18]. Zero-Copying [KLWJ21]. Zero-Overhead [BGL^{+}23].

References

Akin:2019:CAP

[AA19] Berkin Akin and Alaa R. Alameldeen. A case for asymmetric processing in memory. IEEE Computer Architecture Letters, 18(1):22–25, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Arafa:2019:PGS

[ABC^{+}19] Yehia Arafa, Abdel-Hameed A. Badawy, Gopinath Chennupati, Nandakishore Santhi, and Stephan Eidenbenz. PPT–GPU: Scalable GPU performance modeling. IEEE Computer Architecture Letters, 18(1):55–58, January/ June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

August:2007:UOS

[ACG^{+}07] D. August, J. Chang, S. Girbal, D. Gracia-Perez, G. Mouchard, D. A. Penry, O. Temam, and N. Vachharajani. UNISIM: an open simulation environment and library for complex architecture design and collaborative development. IEEE Computer Architecture Letters, 6(2):45–48, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Alvarez:2002:IRF

[ACSV02] C. Alvarez, J. Corbal, E. Salami, and M. Valero. Initial results on fuzzy floating point computation for multimedia processors. IEEE Computer Architecture Letters, 1(1):1, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ahn:2006:DPA

[AD06] Jung Ho Ahn and W. J. Dally.

Data parallel address architecture. IEEE Computer Architecture Letters, 5(1):30–33, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Agrawal:2019:MPS

[ADS^{+}19] V. Agrawal, M. A. Dinani, Y. Shui, M. Ferdman, and N. Honarmand. Massively parallel server processors. IEEE Computer Architecture Letters, 18(1):75–78, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Adileh:2017:MPH

[AEJE17] Almutaz Adileh, Stijn Eyerman, Aamer Jaleel, and Lieven Eeckhout. Mind the power holes: Sifting operating points in power-limited heterogeneous multicores. IEEE Computer Architecture Letters, 16(1):56–59, January/ June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ahmadvand:2017:UDV

[AG17] Hossein Ahmadvand and Maziar Goudarzi. Using data variety for efficient progressive big data processing in warehouse-scale computers. IEEE Computer Architecture Letters, 16(2):166–169, July/ December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

AlBarakat:2018:MFM

[AGJ18] Laith M. AlBarakat, V. Paul Gratz, and Daniel A. Jiménez. MTB-Fetch: Multithreading aware hardware prefetching for chip multiprocessors.

IEEE Computer Architecture Letters, 17(2):175–178, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Asheim:2021:BXS

[AGK21] Truls Asheim, Boris Grot, and Rakesh Kumar. BTB-X: a storage-effective BTB organization. IEEE Computer Architecture Letters, 20 (2):134–137, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Alian:2016:PGS

[AKK16] Mohammad Alian, Daehoon Kim, and Nam Sung Kim. pdgem5: Simulation infrastructure for parallel/distributed computer systems. IEEE Computer Architecture Letters, 15(1):41–44, January/ June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ansari:2019:CLO

[ALKSA19] Ali Ansari, Pejman Lotfi-Kamran, and Hamid Sarbazi-Azad. Code layout optimization for near-ideal instruction cache. IEEE Computer Architecture Letters, 18(2):124– 127, July 2019. ISSN 1556-6064.

Ahn:2009:MDE

[ALSJ09] Jung Ho Ahn, Jacob Leverich, Robert S. Schreiber, and Norman P. Jouppi. Multicore DIMM: an energy efficient memory module with independently controlled DRAMs. IEEE Computer Architecture Letters, 8(1):5–8, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Azriel:2015:PMT

[AMW15] Leonid Azriel, Avi Mendelson, and Uri Weiser. Peripheral memory: a technique for fighting memory bandwidth bottleneck. IEEE Computer Architecture Letters, 14(1): 54–57, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2008:EBC

[Ano08a] Anonymous. Editorial board [cover2]. IEEE Computer Architecture Letters, 7(2):c2, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2008:FC

[Ano08b] Anonymous. [Front cover]. IEEE Computer Architecture Letters, 7(2):c1, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2008:ICS

[Ano08c] Anonymous. IEEE Computer Society [cover 4]. IEEE Computer Architecture Letters, 7 (2):c4, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2008:IA

[Ano08d] Anonymous. Information for authors. IEEE Computer Architecture Letters, 7(2):c3, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:AIC

[Ano09a] Anonymous. Ad — IEEE Computer Society Digital Library. IEEE Computer Architecture Letters, 8(1):36, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:Aa

[Ano09b] Anonymous. [Advertisement].

IEEE Computer Architecture Letters, 8(1):35, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:Ab

[Ano09c] Anonymous. [Advertisement].

IEEE Computer Architecture Letters, 8(2):68, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:Ac

[Ano09d] Anonymous. [Advertisement].

IEEE Computer Architecture Letters, 8(2):69, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:Ad

[Ano09e] Anonymous. [Advertisement]. IEEE Computer Architecture Letters, 8(2):70, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:Ae

[Ano09f] Anonymous. [Advertisement]. IEEE Computer Architecture Letters, 8(2):71, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:Af

[Ano09g] Anonymous. [Advertisement]. IEEE Computer Architecture Letters, 8(2):72, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:EBCa

[Ano09h] Anonymous. Editorial board [cover2]. IEEE Computer Architecture Letters, 8(1):c2, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:EBCb

[Ano09i] Anonymous. Editorial board [cover2]. IEEE Computer Architecture Letters, 8(2):c2, July/December 2009. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:FCa

[Ano09j] Anonymous. [Front cover].

IEEE Computer Architecture Letters, 8(1):c1, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:FCb

[Ano09k] Anonymous. [Front cover].

IEEE Computer Architecture Letters, 8(2):c1, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:ICSa

[Ano09l] Anonymous. IEEE Computer Society [cover4]. IEEE Computer Architecture Letters, 8 (1):c4, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:ICSb

[Ano09m] Anonymous. IEEE Computer Society [cover4]. IEEE Computer Architecture Letters, 8 (2):c4, July/December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:IAa

[Ano09n] Anonymous. Information for authors. IEEE Computer Architecture Letters, 8(1):c3, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2009:IAb

[Ano09o] Anonymous. Information for authors. IEEE Computer Architecture Letters, 8(2):c3, July/December 2009. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:ICSb

[Ano10a] Anonymous. 2011 IEEE Computer Society simulator design competition. IEEE Computer Architecture Letters, 9 (2):66, July/December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:AIT

[Ano10b]  Anonymous.  Advertise-

ment — IEEE Transactions on Computers celebrates 60 years. IEEE Computer Architecture Letters, 9(2):65, July/ December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:ACS

[Ano10c] Anonymous. Advertisement — Computer Society Digital Library. IEEE Computer Architecture Letters, 9(2):72, July/December 2010. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:ADY

[Ano10d] Anonymous. Advertisement — distinguish yourself with the CSDP. IEEE Computer Architecture Letters, 9(2):68, July/December 2010. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:ASS

[Ano10e] Anonymous. Advertisement — special student offer. IEEE Computer Architecture Letters, 9(2):67, July/December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:ASC

[Ano10f] Anonymous. Advertisement — stay connected to the IEEE Computer Society. IEEE Computer Architecture Letters, 9(2):71, July/December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:CPS

[Ano10g] Anonymous. Conference Proceedings Services (CPS) [advertisement]. IEEE Computer Architecture Letters, 9 (2):69, July/December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:EBCa

[Ano10h] Anonymous. Editorial board [cover2]. IEEE Computer Architecture Letters, 9(1):c2, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:EBCb

[Ano10i] Anonymous. Editorial board [cover2]. IEEE Computer Architecture Letters, 9(2):c2, July/December 2010. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:FCa

[Ano10j] Anonymous. [Front cover]. IEEE Computer Architecture Letters, 9(1):c1, January/ June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:FCb

[Ano10k] Anonymous. [Front cover]. IEEE Computer Architecture Letters, 9(2):c1, July/ December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:ICSa

[Ano10l] Anonymous. IEEE Computer Society [cover4]. IEEE Computer Architecture Letters, 9 (1):c4, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:ICSd

[Ano10m] Anonymous. IEEE Computer Society [cover4]. IEEE Computer Architecture Letters, 9 (2):c4, July/December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:ICSc

[Ano10n] Anonymous. IEEE Computer Society jobs. IEEE Computer Architecture Letters, 9 (2):70, July/December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:IAa

[Ano10o] Anonymous. Information for authors. IEEE Computer Architecture Letters, 9(1):c3, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2010:IAb

[Ano10p] Anonymous. Information for authors. IEEE Computer Architecture Letters, 9(2):c3, July/December 2010. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:AI

[Ano11a] Anonymous. 2010 annual index. IEEE Computer Architecture Letters, 10(1):??, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:RL

[Ano11b] Anonymous. 2010 reviewers list. IEEE Computer Architecture Letters, 10(1):28, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:Ca

[Ano11c] Anonymous. Cover 2. IEEE Computer Architecture Letters, 10(1):c2, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:Cb

[Ano11d] Anonymous. Cover 3. IEEE Computer Architecture Letters, 10(1):c3, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:Cd

[Ano11e] Anonymous. Cover 3. IEEE Computer Architecture Letters, 10(2):c3, July/December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:Cc

[Ano11f] Anonymous. Cover 4. IEEE Computer Architecture Letters, 10(1):c4, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:FCa

[Ano11g] Anonymous. [Front cover]. IEEE Computer Architecture Letters, 10(1):c1, January/ June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:FCb

[Ano11h] Anonymous. [Front cover].

IEEE Computer Architecture Letters, 10(2):c1, July/ December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:ICS

[Ano11i] Anonymous. IEEE Computer Society [society information]. IEEE Computer Architecture Letters, 10(2):c4, July/December 2011. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2011:PI

[Ano11j] Anonymous. Publication information. IEEE Computer Architecture Letters, 10(2):c2, July/December 2011. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:AI

[Ano12a] Anonymous. 2011 annual index. IEEE Computer Architecture Letters, 11(1):??, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:RL

[Ano12b] Anonymous. 2011 reviewers list. IEEE Computer Architecture Letters, 11(1):25–26, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:ACP

[Ano12c] Anonymous. Advertisement — Conference Publishing Services (CPS). IEEE Computer Architecture Letters, 11(1):28, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:BC

[Ano12d] Anonymous. [Back cover]. IEEE Computer Architecture Letters, 11(2):c4, July/ December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:BIC

[Ano12e] Anonymous. [Back inside cover]. IEEE Computer Architecture Letters, 11(2):c3, July/December 2012. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:Ca

[Ano12f] Anonymous. [Cover2]. IEEE Computer Architecture Letters, 11(1):c2, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:Cb

[Ano12g] Anonymous. [Cover3]. IEEE Computer Architecture Letters, 11(1):c3, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:FCT

[Ano12h] Anonymous. [Front cover and table of contents]. IEEE Computer Architecture Letters, 11(1):c1, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:FIC

[Ano12i] Anonymous. [Front inside cover]. IEEE Computer Architecture Letters, 11(2):c2, July/December 2012. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:ICS

[Ano12j] Anonymous. IEEE Computer Society [back cover]. IEEE Computer Architecture Letters, 11(1):c4, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2012:TNQ

[Ano12k] Anonymous. There now is a quick and easy way to find out about our collection of Transactions [advertisement]. IEEE Computer Architecture Letters, 11(1):26, January/ June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:AI

[Ano13a] Anonymous. 2012 annual index. IEEE Computer Architecture Letters, 12(1):1–4, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:RL

[Ano13b] Anonymous. 2012 reviewers list. IEEE Computer Architecture Letters, 12(1):33–34, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:BC

[Ano13c] Anonymous. [Back cover]. IEEE Computer Architecture Letters, 12(2):c4, July/ December 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:BIC

[Ano13d] Anonymous. [Back inside cover]. IEEE Computer Architecture Letters, 12(2):c3, July/December 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:ITN

[Ano13e] Anonymous. IEEE Transactions newsletter. IEEE Computer Architecture Letters, 12 (1):36, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:FC

[Ano13f] Anonymous. [Front cover]. IEEE Computer Architecture Letters, 12(2):c1, July/ December 2013. CODEN ????

ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:FIC

[Ano13g] Anonymous. [Front inside cover]. IEEE Computer Architecture Letters, 12(2):c2, July/December 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:IOAa

[Ano13h] Anonymous. IEEE open access publishing. IEEE Computer Architecture Letters, 12 (1):35, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:IOAb

[Ano13i] Anonymous. IEEE open access publishing. IEEE Computer Architecture Letters, 12 (2):71, July/December 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2013:SCI

[Ano13j] Anonymous. Stay connected to the IEEE Computer Society. IEEE Computer Architecture Letters, 12(2):72, July/December 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2014:ICAa

[Ano14a] Anonymous. IEEE Computer Architecture Letters Editorial Board. IEEE Computer Architecture Letters, 13(2):C2, July/December 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2014:ICAb

[Ano14b] Anonymous. IEEE Computer Architecture Letters information for authors. IEEE Computer Architecture Letters, 13 (2):C3, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2014:ITPa

[Ano14c] Anonymous. IEEE Transactions on Pattern Analysis and Machine Intelligence Editorial Board. IEEE Computer Architecture Letters, 13 (1):C2, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2014:ITPb

[Ano14d] Anonymous. IEEE Transactions on Pattern Analysis and Machine Intelligence information for authors. IEEE Computer Architecture Letters, 13 (1):C3, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2014:ICSa

[Ano14e] Anonymous. IEEE Computer Society. IEEE Computer Architecture Letters, 13(1):C4, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2014:ICSb

[Ano14f] Anonymous. IEEE Computer Society [advertisement]. IEEE Computer Architecture Letters, 13(2):C4, July/ December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2014:TCa

[Ano14g] Anonymous. Table of contents. IEEE Computer Architecture Letters, 13(1):C1–C4, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2014:TCb

[Ano14h] Anonymous. Table of contents. IEEE Computer Architecture Letters, 13(2):C1, July/December 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:IIC

[Ano15a] Anonymous. 2014 index IEEE Computer Architecture Letters vol. 13. IEEE Computer Architecture Letters, 14 (1):1–5, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:ICAa

[Ano15b] Anonymous. IEEE Computer Architecture Letters Editorial Board. IEEE Computer Architecture Letters, 14(1):C2, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:ICAc

[Ano15c] Anonymous. IEEE Computer Architecture Letters Editorial Board. IEEE Computer Architecture Letters, 14(2):C2, July/December 2015. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:ICAb

[Ano15d] Anonymous. IEEE Computer Architecture Letters information for authors. IEEE Computer Architecture Letters, 14 (1):C3, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:ICAd

[Ano15e] Anonymous. IEEE Computer Architecture Letters information for authors. IEEE Computer Architecture Letters, 14 (2):C3, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:ICSa

[Ano15f] Anonymous. IEEE Computer Society. IEEE Computer Architecture Letters, 14(1):C4, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:ICSb

[Ano15g] Anonymous. IEEE Computer Society. IEEE Computer Architecture Letters, 14(2):C4, July/December 2015. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:RSC

[Ano15h] Anonymous. Rock stars of cybersecurity 2015 conference. IEEE Computer Architecture Letters, 14(1):84, January/ June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:RSW

[Ano15i] Anonymous. Rock stars of wearables. IEEE Computer Architecture Letters, 14(1):83, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:TCa

[Ano15j] Anonymous. Table of contents. IEEE Computer Architecture Letters, 14(1):C1, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2015:TCb

[Ano15k] Anonymous. Table of contents. IEEE Computer Architecture Letters, 14(2):C1, July/December 2015. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:IICa

[Ano16a] Anonymous. 2015 index IEEE Computer Architecture Letters vol. 14. IEEE Computer Architecture Letters, 15 (1):1–6, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:IICb

[Ano16b] Anonymous. 2015 index IEEE Computer Architecture Letters vol. 14. IEEE Computer Architecture Letters, 15(1):1– 6, January/June 2016. ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:BC

[Ano16c] Anonymous. [Back cover]. IEEE Computer Architecture Letters, 15(1):C4, January/ June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:ICS

[Ano16d] Anonymous. IEEE Cyber Security. IEEE Computer Architecture Letters, 15(1):68, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:Ca

[Ano16e] Anonymous. Cover. IEEE Computer Architecture Letters, 15(1):C2, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:Cb

[Ano16f] Anonymous. Cover. IEEE Computer Architecture Letters, 15(1):C2, January/June 2016. ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:Cc

[Ano16g] Anonymous. Cover. IEEE Computer Architecture Letters, 15(1):C3, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:Cd

[Ano16h] Anonymous. Cover. IEEE Computer Architecture Letters, 15(1):C3, January/June 2016. ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:Ce

[Ano16i] Anonymous. Cover. IEEE Computer Architecture Letters, 15(2):C2, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:Cf

[Ano16j] Anonymous. Cover. IEEE Computer Architecture Letters, 15(2):C3, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:ENM

[Ano16k] Anonymous. Experience the newest and most advanced thinking in big data analytics. IEEE Computer Architecture Letters, 15(1):67, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:IICc

[Ano16l]  Anonymous.  Introducing

IEEE Collabratec. IEEE Computer Architecture Letters, 15(1):66, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:IICd

[Ano16m]  Anonymous.  Introducing

IEEE Collabratec. IEEE Computer Architecture Letters, 15(1):66, January/June 2016. ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:TCa

[Ano16n] Anonymous. Table of contents. IEEE Computer Architecture Letters, 15(1):C1, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:TCb

[Ano16o] Anonymous. Table of contents. IEEE Computer Architecture Letters, 15(2):C1, July/December 2016. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2016:TCBa

[Ano16p] Anonymous. Table of contents [back cover]. IEEE Computer Architecture Letters, 15(2):C4, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2017:IIC

[Ano17] Anonymous. 2016 index IEEE Computer Architecture Letters vol. 15. IEEE Computer Architecture Letters, 16 (1):1–6, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2018:IIC

[Ano18] Anonymous. 2017 index IEEE Computer Architecture Letters vol. 16. IEEE Computer Architecture Letters, 17 (1):1–6, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2019:IIC

[Ano19] Anonymous. 2018 index IEEE Computer Architecture Letters vol. 17. IEEE Computer Architecture Letters, 18 (1):1–8, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2020:IIC

[Ano20] Anonymous. 2019 index IEEE Computer Architecture Letters vol. 18. IEEE Computer Architecture Letters, 19(1):1– 8, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2021:IIC

[Ano21] Anonymous. 2020 index IEEE Computer Architecture Letters vol. 19. IEEE Computer Architecture Letters, 20(1):1– 7, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Anonymous:2022:IIC

[Ano22] Anonymous. 2021 index IEEE Computer Architecture Letters vol. 20. IEEE Computer Architecture Letters, 21(1):1– 8, January/June 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Antelo:2009:CBF

[Ant09] Elisardo Antelo. A comment on “Beyond Fat-tree: Unidirectional Load-Balanced Multistage Interconnection Network”. IEEE Computer Architecture Letters, 8(1):33–34, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic). See [GVG^{+}08].

Ajdari:2018:SHB

[APK^{+}18] Mohammadamin Ajdari, Pyeongsu Park, Dongup Kwon, Joonsung Kim, and Jangwoo Kim. A scalable HW-based inline deduplication for SSD arrays. IEEE Computer Architecture Letters, 17(1):47–50, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Amarnath:2021:HAS

[APK^{+}21] Aporva Amarnath, Subhankar Pal, Hiwot Tadese Kassa, Augusto Vega, Alper Buyuktosunoglu, Hubertus Franke, John-David Wellman, Ronald Dreslinski, and Pradip Bose. Heterogeneityaware scheduling on SoCs for autonomous vehicles. IEEE Computer Architecture Letters, 20(2):82–85, July/ December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Ayanzadeh:2024:ERR

[AQ24] Ramin Ayanzadeh and Moinuddin Qureshi. Enhancing the reach and reliability of quantum annealers by pruning longer chains. IEEE Computer Architecture Letters, 23(1):25–28, January/ June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Arelakis:2014:CVA

[AS14] Angelos Arelakis and Per Stenström. A case for a valueaware cache. IEEE Computer Architecture Letters, 13 (1):1–4, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Almatrood:2018:DGP

[AS18] Amjad F. Almatrood and Harpreet Singh. Design of generalized pipeline cellular array in quantum-dot cellular automata. IEEE Computer Architecture Letters, 17(1): 29–32, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Alian:2021:IOI

[ASK^{+}21] Mohammad Alian, Jongmin Shin, Ki-Dong Kang, Ren Wang, Alexandros Daglis, Daehoon Kim, and Nam Sung Kim. IDIO: Orchestrating inbound network data on server processors. IEEE Computer Architecture Letters, 20(1): 30–33, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Aimoniotis:2021:RBC

[ASSK21] Pavlos Aimoniotis, Christos Sakalis, Magnus Själander, and Stefanos Kaxiras. Reorder buffer contention: a forward speculative interference attack for speculation invariant instructions. IEEE Computer Architecture Letters, 20 (2):162–165, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Altaf:2015:LPM

[AW15] Muhammad Shoaib Bin Altaf and David A. Wood. LogCA: a performance model for hardware accelerators. IEEE Computer Architecture Letters, 14(2):132–135, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Angstadt:2018:MOS

[AWD^{+}18] Kevin Angstadt, Jack Wadden, Vinh Dang, Ted Xie, Dan Kramp, Westley Weimer, Mircea Stan, and Kevin Skadron. MNCaRT: an opensource, multi-architecture automata-processing research and execution ecosystem. IEEE Computer Architecture Letters, 17(1):84–87, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Accetti:2022:SCE

[AYL22] Cecil Accetti, Rendong Ying, and Peilin Liu. Structured combinators for efficient graph reduction. IEEE Computer Architecture Letters, 21(2):73–76, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Barber:2019:ISD

[BBZ^{+}19] Kristin Barber, Anys Bacha, Li Zhou, Yinqian Zhang, and Radu Teodorescu. Isolating speculative data to prevent transient execution attacks. IEEE Computer Architecture Letters, 18(2):178–181, July 2019. ISSN 1556-6064.

Balfour:2008:EEP

[BDBS^{+}08] J. Balfour, W. Dally, D. Black-Schaffer, V. Parikh, and J. Park. An energy-efficient processor architecture for embedded systems. IEEE Computer Architecture Letters, 7 (1):29–32, January 2008. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Bracy:2006:DAC

[BDJ06] A. Bracy, K. Doshi, and Q. Jacobson. Disintermediated active communication. IEEE Computer Architecture Letters, 5(2):15, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Blem:2013:MMA

[BEA^{+}13] Emily Blem, Hadi Esmaeilzadeh, Renee St Amant, Karthikeyan Sankaralingam, and Doug Burger. Multicore model from abstract single core inputs. IEEE Computer Architecture Letters, 12(2): 59–62, July/December 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Bae:2023:ISF

[BGL^{+}23] Hanyeoreum Bae, Donghyun Gouk, Seungjun Lee, Jiseon Kim, Sungjoon Koh, Jie Zhang, and Myoungsoo Jung. Intelligent SSD firmware for zero-overhead journaling.

IEEE Computer Architecture Letters, 22(1):25–28, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Boroumand:2017:LEC

[BGP^{+}17] Amirali Boroumand, Saugata Ghose, Minesh Patel, Hasan Hassan, Brandon Lucia, Kevin Hsieh, Krishna T. Malladi, Hongzhong Zheng, and Onur Mutlu. LazyPIM: an efficient cache coherence mechanism for processing-in-memory. IEEE Computer Architecture Letters, 16(1): 46–50, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Bodduna:2020:BRS

[BGS^{+}20] Rahul Bodduna, Vinod Ganesan, Patanjali SLPSK, Kamakoti Veezhinathan, and Chester Rebeiro. Brutus: Refuting the security claims of the cache timing randomization countermeasure proposed in CEASER. IEEE Computer Architecture Letters, 19(1):9– 12, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Barber:2022:PSA

[BGZT22] Kristin Barber, Moein Ghaniyoun, Yinqian Zhang, and Radu Teodorescu. A presilicon approach to discovering microarchitectural vulnerabilities in security critical applications. IEEE Computer Architecture Letters, 21(1):9– 12, January/June 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Balfour:2009:ORE

[BHD09] James Balfour, R. Curtis Harting, and William J. Dally. Operand registers and explicit operand forwarding. IEEE Computer Architecture Letters, 8(2):60–63, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Basak:2018:ECC

[BHL^{+}18] Abanti Basak, Xing Hu, Shuangchen Li, Sang Min Oh, and Yuan Xie. Exploring core and cache hierarchy bottlenecks in graph processing workloads. IEEE Computer Architecture Letters, 17 (2):197–200, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Bhardwaj:2019:DOC

[BHY^{+}19] K. Bhardwaj, M. Havasi, Y. Yao, D. M. Brooks, J. M. H. Lobato, and G. Wei. Determining optimal coherency interface for many-accelerator SoCs using Bayesian optimization. IEEE Computer Architecture Letters, 18(2):119–123, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Baldassin:2009:CEC

[BKA^{+}09] Alexandro Baldassin, Felipe Klein, Guido Araujo, Rodolfo Azevedo, and Paulo Centoducatte. Characterizing the energy consumption of software transactional memory. IEEE Computer Architecture Letters, 8(2):56–59, July/December 2009. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Bakhshalipour:2017:ETD

[BLKSA17] Mohammad Bakhshalipour, Pejman Lotfi-Kamran, and Hamid Sarbazi-Azad. An efficient temporal data prefetcher for L1 caches. IEEE Computer Architecture Letters, 16 (2):99–102, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Blundell:2006:STM

[BLM06] C. Blundell, E. C. Lewis, and M. M. K. Martin. Subtleties of transactional memory atomicity semantics. IEEE Computer Architecture Letters, 5(2):17, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Biswas:2008:CAA

[BREM08] A. Biswas, P. Racunas, J. Emer, and S. Mukherjee. Computing accurate AVFs using ACE analysis on performance models: a rebuttal.

IEEE Computer Architecture Letters, 7(1):21–24, January 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Boran:2021:FGS

[BRUS21] Nirmal Kumar Boran, Shubhankit Rathore, Meet Udeshi, and Virendra Singh. Finegrained scheduling in heterogeneous-ISA architectures. IEEE Computer Architecture Letters, 20(1):9–12, January/ June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Beckmann:2017:CCM

[BS17] Nathan Beckmann and Daniel Sanchez. Cache calculus: Modeling caches through differential equations. IEEE Computer Architecture Letters, 16(1):1–5, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Black-Schaffer:2008:HIR

[BSBD^{+}08] D. Black-Schaffer, J. Balfour, W. Dally, V. Parikh, and J. Park. Hierarchical instruction register organization. IEEE Computer Architecture Letters, 7(2):41–44, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Balaji:2019:FEW

[BSD^{+}19] Adarsha Balaji, Shihao Song, Anup Das, Nikil Dutt, Jeff Krichmar, Nagarajan Kandasamy, and Francky Catthoor. A framework to explore workload-specific performance and lifetime tradeoffs in neuromorphic computing. IEEE Computer Architecture Letters, 18(2):149– 152, July 2019. ISSN 1556-6064.

Brana:2023:KSC

[BSMB23] Jennifer Brana, Brian C.

Schwedock, Yatin A. Manerkar, and Nathan Beckmann. Kobold: Simplified cache coherence for cache-attached accelerators. IEEE Computer Architecture Letters, 22(1): 41–44, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Barnes:2009:XBA

[BVL09] Christopher Barnes, Pranav Vaidya, and Jaehwan John Lee. An XML-based ADL framework for automatic generation of multithreaded computer architecture simulators. IEEE Computer Architecture Letters, 8(1):13–16, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Badawy:2017:GLO

[BY17] Abdel-Hameed A. Badawy and Donald Yeung. Guiding locality optimizations for graph computations via reuse distance analysis. IEEE Computer Architecture Letters, 16 (2):119–122, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Chiou:2009:AFF

[CAPS09] Derek Chiou, Hari Angepat, Nikhil A. Patil, and Dam Sunwoo. Accurate functionalfirst multicore simulators. IEEE Computer Architecture Letters, 8(2):64–67, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Chen:2017:IGP

[CCWY17] Li-Jhan Chen, Hsiang-Yun Cheng, Po-Han Wang, and Chia-Lin Yang. Improving GPGPU performance via cache locality aware thread block scheduling. IEEE Computer Architecture Letters, 16 (2):127–131, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Cakmakci:2014:EVA

[ÇE14] Yaman Çakmakçi and Oğuz Ergin. Exploiting virtual addressing for increasing reliability. IEEE Computer Architecture Letters, 13(1):29–32, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Choukse:2018:CEM

[CEA18] Esha Choukse, Mattan Erez, and Alaa Alameldeen. CompressPoints: an evaluation methodology for compressed memory systems. IEEE Computer Architecture Letters, 17 (2):126–129, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Corontzos:2024:DCD

[CF24] Caden Corontzos and Eitan Frachtenberg. Direct-coding DNA with multilevel parallelism. IEEE Computer Architecture Letters, 23(1):21– 24, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Cohen:2003:EOP

[CFM^{+}03] A. Cohen, F. Finkelstein, A. Mendelson, R. Ronen, and D. Rudoy. On estimating optimal performance of CPU dynamic thermal management. IEEE Computer Architecture Letters, 2(1):6, January 2003. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Chen:2014:PEC

[CGY^{+}14] Zheng Chen, Huaxi Gu, Yintang Yang, Luying Bai, and Hui Li. A power efficient and compact optical interconnect for network-on-chip. IEEE Computer Architecture Letters, 13(1):5–8, January/ June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Chowdhury:2018:EMP

[CHK^{+}18] Zamshed Chowdhury, Jonathan D. Harms, S. Karen Khatamifard, Masoud Zabihi, Yang Lv, Andrew P. Lyle, Sachin S. Sapatnekar, Ulya R. Karpuzcu, and Jian-Ping Wang. Efficient in-memory processing using spintronics. IEEE Computer Architecture Letters, 17(1): 42–46, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Citron:2004:ELE

[Cit04] D. Citron. Exploiting low entropy to reduce wire delay. IEEE Computer Architecture Letters, 3(1):1, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Choe:2020:SMP

[CKA20] Wonkyo Choe, Jonghyeon Kim, and Jeongseob Ahn. A study of memory placement on hardware-assisted tiered memory systems. IEEE Computer Architecture Letters, 19 (2):122–125, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Chowdhury:2020:VNM

[CKZ^{+}20] Zamshed I. Chowdhury, S. Karen Khatamifard, Zhaoyong Zheng, Tali Moreshet, R. Iris Bahar, and Ulya R. Karpuzcu. Voltage noise mitigation with barrier approximation. IEEE Computer Architecture Letters, 19(2):155–158, July/ December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Chou:2014:EPE

[CLCG14] Yu-Liang Chou, Shaoshan Liu, Eui-Young Chung, and Jean-Luc Gaudiot. An energy and performance efficient DVFS scheme for irregular parallel divide-and-conquer algorithms on the Intel SCC. IEEE Computer Architecture Letters, 13(1): 13–16, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Choi:2002:LPT

[CLJ^{+}02] Jin-Hyuck Choi, Jung-Hoon Lee, Seh-Woong Jeong, Shin-Dug Kim, and C. Weems. A low power TLB structure for embedded systems. IEEE Computer Architecture Letters, 1(1):3, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Cho:2008:CAL

[CM08] S. Cho and R. Melhem. Corollaries to Amdahl’s Law for energy. IEEE Computer Architecture Letters, 7(1):25–28, January 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Cristal:2003:CRC

[CMLV03] A. Cristal, J. F. Martinez, J. Llosa, and M. Valero. A case for resource-conscious out-of-order processors. IEEE Computer Architecture Letters, 2(1):7, January 2003. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Cota:2014:AMR

[CMP^{+}14] Emilio G. Cota, Paolo Mantovani, Michele Petracca, Mario R. Casu, and Luca P. Carloni. Accelerator memory reuse in the dark silicon era. IEEE Computer Architecture Letters, 13(1):9–12, January/ June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Carlson:2015:EPM

[CNHH15] Trevor E. Carlson, Siddharth Nilakantan, Mark Hempstead, and Wim Heirman. Epoch profiles: Microarchitecturebased application analysis and optimization. IEEE Computer Architecture Letters, 14(1):30–33, January/ June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Choi:2023:UPP

[CPK^{+}23] Jaewan Choi, Jaehyun Park, Kwanhee Kyung, Nam Sung Kim, and Jung Ho Ahn. Unleashing the potential of PIM: Accelerating large batched inference of transformer-based generative models. IEEE Computer Architecture Letters, 22(2):113–116, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Cavus:2020:EPP

[CSSU20] Mustafa Cavus, Mohammed Shatnawi, Resit Sendag, and Augustus K. Uht. Exploring prefetching, pre-execution and branch outcome streaming for in-memory database lookups. IEEE Computer Architecture Letters, 19(1):5–8, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Ceze:2004:CHL

[CST^{+}04] L. Ceze, K. Strauss, J. Tuck, J. Renau, and J. Torrellas. CAVA: Hiding L2 misses with checkpoint-assisted value prediction. IEEE Computer Architecture Letters, 3(1):7, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Carlson:2017:THL

[CTJ^{+}17] Trevor E. Carlson, Kim-Anh Tran, Alexandra Jimborean, Konstantinos Koukos, Magnus Själander, and Stefanos Kaxiras. Transcending hardware limits with software outof-order processing. IEEE Computer Architecture Letters, 16(2):162–165, July/ December 2017. CODEN ????

ISSN 1556-6056 (print), 1556-6064 (electronic).

Chu:2020:HPD

[CTL^{+}20] Zhufei Chu, Huiming Tian, Zeqiang Li, Yinshui Xia, and Lunyao Wang. A highperformance design of generalized pipeline cellular array. IEEE Computer Architecture Letters, 19(1):47–50, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Cakmakci:2016:CPG

[ÇTNL16] Yaman Çakmakçi, Will Toms, Javier Navaridas, and Mikel Lujan. Cyclic power-gating as an alternative to voltage and frequency scaling. IEEE Computer Architecture Letters, 15(2):77–80, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Chen:2015:HSC

[CV15] Jie Chen and Guru Venkataramani. A hardware-software cooperative approach for application energy profiling. IEEE Computer Architecture Letters, 14(1):5–8, January/ June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Cho:2024:EEA

[CVF^{+}24] Minsik Cho, Keivan A. Vahid, Qichen Fu, Saurabh Adya, Carlo C. Del Mundo, Mohammad Rastegari, Devang Naik, and Peter Zatloukal. eDKM: an efficient and accurate train-time weight clustering for large language models. IEEE Computer Architecture Letters, 23(1):37–40, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Chen:2012:NPD

[CVP12] Jie Chen, Guru Venkataramani, and Gabriel Parmer. The need for power debugging in the multi-core environment. IEEE Computer Architecture Letters, 11(2):57– 60, July/December 2012. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Chacon:2022:HTT

[CWK^{+}22] Gino A. Chacon, Charles Williams, Johann Knechtel, Ozgur Sinanoglu, and Paul V. Gratz. Hardware Trojan threats to cache coherence in modern 2.5D chiplet systems. IEEE Computer Architecture Letters, 21(2):133–136, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Christoforidis:2018:CTC

[CXS18] Eleftherios-Iordanis Christoforidis, Sotirios Xydis, and Dimitrios Soudris. CF-TUNE: Collaborative filtering auto-tuning for energy efficient many-core processors. IEEE Computer Architecture Letters, 17(1):25–28, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Che:2020:LMA

[CYAW20] Yuezhi Che, Yuanzhou Yang, Amro Awad, and Rujia Wang. A lightweight memory access pattern obfuscation framework for NVM. IEEE Computer Architecture Letters, 19(2):163–166, July/ December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Choi:2011:EID

[CZYY11] Inseok Choi, Minshu Zhao, Xu Yang, and Donald Yeung. Experience with improving distributed shared cache performance on Tilera’s Tile processor. IEEE Computer Architecture Letters, 10(2): 45–48, July/December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Durkovic:2018:BNS

[DC18] Srdjan Durkovic and Zoran Cica. Birkhoff–von Neumann switch based on greedy scheduling. IEEE Computer Architecture Letters, 17(1): 13–16, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Deb:2012:HSC

[DCG12] Abhishek Deb, Josep Maria Codina, and Antonio Gonzalez. A HW/SW codesigned programmable functional unit. IEEE Computer Architecture Letters, 11(1):9– 12, January/June 2012. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Do:2018:CRL

[DD18] Sang Wook Stephen Do and Michel Dubois. Core reliability: Leveraging hardware transactional memory. IEEE Computer Architecture Letters, 17(2):105–108, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Delshadtehrani:2018:NPM

[DEC^{+}18] Leila Delshadtehrani, Schuyler Eldridge, Sadullah Canakci, Manuel Egele, and Ajay Joshi. Nile: a programmable monitoring coprocessor. IEEE Computer Architecture Letters, 17(1):92– 95, January/June 2018. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Delimitrou:2013:NCD

[DK13] Christina Delimitrou and Christos Kozyrakis. The Netflix challenge: Datacenter edition. IEEE Computer Architecture Letters, 12(1):29–32, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Delimitrou:2016:SID

[DK16] Christina Delimitrou and Christos Kozyrakis. Security implications of data mining in cloud scheduling. IEEE Computer Architecture Letters, 15 (2):109–112, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Dieter:2007:LCM

[DKD07] W. R. Dieter, A. Kaveti, and H. G. Dietz. Low-cost microarchitectural support for improved floating-point accuracy. IEEE Computer Architecture Letters, 6(1):13–16, January 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Denby:2019:OEC

[DL19] Bradley Denby and Brandon Lucia. Orbital edge computing: Machine inference in space. IEEE Computer Architecture Letters, 18(1):59–62, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Desai:2020:PAH

[DL20] Harsh Desai and Brandon Lucia. A power-aware heterogeneous architecture scaling model for energy-harvesting computers. IEEE Computer Architecture Letters, 19(1):

68–71, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Donald:2006:EPP

[DM06] J. Donald and M. Martonosi.

An efficient, practical parallelization methodology for multicore architecture simulation. IEEE Computer Architecture Letters, 5(2):14, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Das:2007:MMC

[DOM^{+}07] A. Das, S. Ozdemir, G. Memik, J. Zambreno, and A. Choudhary. Microarchitectures for managing chip revenues under process variations. IEEE Computer Architecture Letters, 6(2):29–32, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Das:2008:MMC

[DOM^{+}08] A. Das, S. Ozdemir, G. Memik, J. Zambreno, and A. Choudhary. Microarchitectures for managing chip revenues under process variations. IEEE Computer Architecture Letters, 7(1):5–8, January 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Daya:2016:THP

[DPC16] Bhavya K. Daya, Li-Shiuan Peh, and Anantha P. Chandrakasan. Towards highperformance bufferless NoCs with SCEPTER. IEEE Computer Architecture Letters, 15(1):62–65, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Deshpande:2023:TPB

[DPP23] Chandana S. Deshpande, Arthur Perais, and Frédéric Pétrot. Toward practical 128bit general purpose microarchitectures. IEEE Computer Architecture Letters, 22(2): 81–84, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Davis:2012:IVL

[DRGA12] John D. Davis, Suzanne Rivoire, Moises Goldszmidt, and Ehsan K. Ardestani. Including variability in largescale cluster power models. IEEE Computer Architecture Letters, 11(2):29–32, July/ December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Desai:2009:AIC

[DS09] Aniruddha Desai and Jugdutt Singh. Architecture independent characterization of embedded Java workloads. IEEE Computer Architecture Letters, 8(1):29–32, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Delimitrou:2012:DDS

[DSVK12] Christina Delimitrou, Sriram Sankar, Kushagra Vaid, and Christos Kozyrakis. Decoupling datacenter storage studies from access to large-scale applications. IEEE Computer Architecture Letters, 11(2): 53–56, July/December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Duong:2013:CAS

[DV13] Nam Duong and Alexander V. Veidenbaum. Compilerassisted, selective out-of-order commit. IEEE Computer Architecture Letters, 12(1):21– 24, January/June 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

DePestel:2018:RRP

[DVAE18] Sander De Pestel, Sam Van den Steen, Shoaib Akram, and Lieven Eeckhout. RPPM: Rapid performance prediction of multithreaded applications on multicore hardware. IEEE Computer Architecture Letters, 17(2):183–186, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Diamantopoulos:2015:MMI

[DXSS15] Dionysios Diamantopoulos, Sotirios Xydis, Kostas Siozios, and Dimitrios Soudris. Mitigating memory-induced dark silicon in many-accelerator architectures. IEEE Computer Architecture Letters, 14 (2):136–139, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Eyerman:2014:RCW

[EE14] Stijn Eyerman and Lieven Eeckhout. Restating the case for weighted-IPC metrics to evaluate multiprogram workload performance. IEEE Computer Architecture Letters, 13(2):93–96, July/ December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Eker:2016:EEC

[EE16] Abdulaziz Eker and Oğuz Ergin. Exploiting existing copies in register file for soft error correction. IEEE Computer Architecture Letters, 15(1): 17–20, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Eeckhout:2013:MNE

[Eec13] Lieven Eeckhout. A message from the new Editorin-Chief and introduction of new Associate Editors. IEEE Computer Architecture Letters, 12(1):2, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Eeckhout:2022:FOM

[Eec22] Lieven Eeckhout. A firstorder model to assess computer architecture sustainability. IEEE Computer Architecture Letters, 21(2):137– 140, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Etsion:2007:PPT

[EF07] Y. Etsion and D. G. Feitelson. Probabilistic prediction of temporal locality. IEEE Computer Architecture Letters, 6(1):17–20, January 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Efraim:2014:EAR

[EGWM14] Rotem Efraim, Ran Ginosar, C. Weiser, and Avi Mendelson. Energy aware race to halt: a down to EARtH approach for platform energy management. IEEE Computer Architecture Letters, 13(1):25–28, January/ June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Eyerman:2018:MSC

[EHDH18] Stijn Eyerman, Wim Heirman, Kristof Du Bois, and Ibrahim Hur. Multi-stage CPI stacks. IEEE Computer Architecture Letters, 17(1):55– 58, January/June 2018. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Eyerman:2020:BOB

[EHdSH20] Stijn Eyerman, Wim Heirman, Sam Van den Steen, and Ibrahim Hur. Breaking in-order branch miss recovery. IEEE Computer Architecture Letters, 19(1):30– 33, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Eyerman:2021:MDT

[EHH21] Stijn Eyerman, Wim Heirman, and Ibrahim Hur. Modeling DRAM timing in parallel simulators with immediate-response memory model. IEEE Computer Architecture Letters, 20(2): 90–93, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Escofet:2023:HQA

[EOA^{+}23] Pau Escofet, Anabel Ovide, Carmen G. Almudever, Eduard Alarcón, and Sergi Abadal. Hungarian qubit assignment for optimized mapping of quantum circuits on multi-core architectures. IEEE Computer Architecture Letters, 22(2):161–164, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Eisley:2006:NCC

[EPS06] N. Eisley, Li-Shiuan Peh, and Li Shang. In-network cache coherence. IEEE Computer Architecture Letters, 5(1):34– 37, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ergin:2006:ENV

[EUVG06] O. Ergin, O. Unsal, X. Vera, and A. Gonzalez. Exploiting narrow values for soft error tolerance. IEEE Computer Architecture Letters, 5(2):12, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Feng:2023:SOW

[FAR^{+}23] Justin Feng, Fatemeh Arkannezhad, Christopher Ryu, Enoch Huang, Siddhant Gupta, and Nader Sehatbakhsh. Simulating our way to safer software: a tale of integrating microarchitecture simulation and leakage estimation modeling. IEEE Computer Architecture Letters, 22 (2):109–112, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

France:2024:RSA

[FBN^{+}24] Loïc France, Florent Bruguier, David Novo, Maria Mushtaq, and Pascal Benoit. Reducing the silicon area overhead of counter-based rowhammer mitigations. IEEE Computer Architecture Letters, 23(1): 61–64, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Flich:2008:LBD

[FD08] J. Flich and J. Duato. Logicbased distributed routing for NoCs. IEEE Computer Architecture Letters, 7(1):13–16, January 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Farmahini-Farahani:2015:DAA

[FFAMK15] Amin Farmahini-Farahani, Jung Ho Ahn, Katherine Morrow, and Nam Sung Kim. DRAMA: an architecture for accelerated processing near memory. IEEE Computer Architecture Letters, 14(1): 26–29, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Fang:2010:BRP

[FHL^{+}10] Zhen Fang, Erik G. Hallnor, Bin Li, Michael Leddige, Donglai Dai, Seung Eun Lee, Srihari Makineni, and Ravi Iyer. Boomerang: Reducing power consumption of response packets in NoCs with minimal performance impact. IEEE Computer Architecture Letters, 9(2):49–52, July/ December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Fide:2008:PUS

[FJ08] S. Fide and S. Jenks. Proactive use of shared L3 caches to enhance cache communications in multi-core processors. IEEE Computer Architecture Letters, 7(2):57–60, July 2008. CODEN ????

ISSN 1556-6056 (print), 1556-6064 (electronic).

Feng:2017:HHC

[FLSZ17] Liang Feng, Hao Liang, Sharad Sinha, and Wei Zhang. HeteroSim: a heterogeneous CPU–FPGA simulator. IEEE Computer Architecture Letters, 16(1):38–41, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Falahati:2021:DAC

[FPA^{+}21] Hajar Falahati, Masoud Peyro, Hossein Amini, Mehran Taghian, Mohammad Sadrosadati, Pejman Lotfi-Kamran, and Hamid Sarbazi-Azad. Data-aware compression of neural networks. IEEE Computer Architecture Letters, 20(2):94–97, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Fariborz:2022:MSB

[FSO^{+}22] Marjan Fariborz, Mahyar Samani, Terry O’Neill, Jason Lowe-Power, S. J. Ben Yoo, and Venkatesh Akella. A model for scalable and balanced accelerators for graph processing. IEEE Computer Architecture Letters, 21 (2):149–152, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Finlayson:2012:OSP

[FUWT12] Ian Finlayson, Gang-Ryung Uh, David Whalley, and Gary Tyson. An overview of static pipelining. IEEE Computer Architecture Letters, 11(1): 17–20, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Goudarzi:2023:SBP

[GAH^{+}23] Maziar Goudarzi, Reza Azimi, Julian Humecki, Faizaan Rehman, Richard Zhang, Chirag Sethi, Tanishq Bomman, and Yuqi Yang. Bysoftware branch prediction in loops. IEEE Computer Architecture Letters, 22(2):129– 132, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Gaudiot:2009:INE

[Gau09] Jean-Luc Gaudiot. Introducing the new Editor-in-Chief of IEEE Computer Architecture Letters. IEEE Computer Architecture Letters, 8(2):37– 38, July/December 2009. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Guz:2009:MCV

[GBK^{+}09] Zvika Guz, Evgeny Bolotin, Idit Keidar, Avinoam Kolodny, Avi Mendelson, and Uri C. Weiser. Many-core vs. manythread machines: Stay away from the valley. IEEE Computer Architecture Letters, 8(1):25–28, January/ June 2009. CODEN ????

ISSN 1556-6056 (print), 1556-6064 (electronic).

Golshan:2020:HPC

[GBS^{+}20] Fatemeh Golshan, Mohammad Bakhshalipour, Mehran Shakerinava, Ali Ansari, Pejman Lotfi-Kamran, and Hamid Sarbazi-Azad. Harnessing pairwise-correlating data prefetching with runahead metadata. IEEE Computer Architecture Letters, 19 (2):130–133, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Gupta:2006:TOI

[GD06] A. K. Gupta and W. J.

Dally. Topology optimization of interconnection networks. IEEE Computer Architecture Letters, 5(1):10–13, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gan:2018:AIC

[GD18] Yu Gan and Christina Delimitrou. The architectural implications of cloud microservices. IEEE Computer Architecture Letters, 17(2):155–158, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gomez:2004:EFT

[GDF^{+}04] M. E. Gomez, J. Duato, J. Flich, P. Lopez, A. Robles, N. A. Nordbotten, O. Lysne, and T. Skeie. An efficient fault-tolerant routing methodology for meshes and tori. IEEE Computer Architecture Letters, 3(1):3, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gorgues:2016:EPC

[GF16] Miguel Gorgues and Jose Flich. End-point congestion filter for adaptive routing with congestion-insensitive performance. IEEE Computer Architecture Letters, 15 (1):9–12, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gheibi-Fetrat:2024:TTF

[GFAHSA24] Atiyeh Gheibi-Fetrat, Negar Akbarzadeh, Shaahin Hessabi, and Hamid Sarbazi-Azad. Tulip: Turn-free low-power network-on-chip. IEEE Computer Architecture Letters, 23 (1):5–8, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Gou:2011:ESH

[GG11] Chunyang Gou and Georgi N. Gaydadjiev. Exploiting SPMD horizontal locality. IEEE Computer Architecture Letters, 10(1):20–23, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Garland:2017:LCM

[GG17] James Garland and David Gregg. Low complexity multiply accumulate unit for weight-sharing convolutional neural networks. IEEE Computer Architecture Letters, 16 (2):132–135, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Garcia:2016:CMP

[GGM^{+}16] P. Garcia, T. Gomes, J. Monteiro, A. Tavares, and M. Ekpanyapong. On-chip message passing sub-system for embedded inter-domain communication. IEEE Computer Architecture Letters, 15(1):33– 36, January/June 2016. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Golestani:2019:PMB

[GGS19] H. Golestani, G. Gupta, and R. Sen. Performance modeling and bottleneck analysis of EDGE processors using dependence graphs. IEEE Computer Architecture Letters, 18(1):79–82, January/ June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Golden:2024:SVV

[GIH^{+}24] Courtney Golden, Dan Ilan, Caroline Huang, Niansong Zhang, Zhiru Zhang, and Christopher Batten. Supporting a virtual vector instruction set on a commercial compute-in-SRAM accelerator. IEEE Computer Architecture Letters, 23(1):29– 32, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Ghasemi:2021:MPE

[GJ21] Fatemeh Ghasemi and Magnus Jahre. Modeling periodic energy-harvesting computing systems. IEEE Computer Architecture Letters, 20 (2):142–145, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Gouk:2022:PHA

[GKK^{+}22] Donghyun Gouk, Seungkwan Kang, Miryeong Kwon, Junhyeok Jang, Hyunkyu Choi, Sangwon Lee, and Myoungsoo Jung. PreGNN: Hardware acceleration to take preprocessing off the critical path in graph neural networks. IEEE Computer Architecture Letters, 21(2):117–120, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Guz:2007:NCO

[GKKW07] Z. Guz, I. Keidar, A. Kolodny, and U. Weiser. Nahalal: Cache organization for chip multiprocessors. IEEE Computer Architecture Letters, 6 (1):21–24, January 2007. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gu:2020:NTC

[GLH^{+}20] Peng Gu, Benjamin S. Lim, Wenqin Huangfu, Krishan T. Malladi, Andrew Chang, and Yuan Xie. NMTSim: Transaction-command based simulator for new memory technology devices. IEEE Computer Architecture Letters, 19(1):76–79, January/ June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Gurumurthi:2021:HRE

[GLJ^{+}21] Sudhanva Gurumurthi, Kijun Lee, Munseon Jang, Vilas Sridharan, Aaron Nygren, Yesin Ryu, Kyomin Sohn, Taekyun Kim, and Hoeju Chung. HBM3 RAS: Enhancing resilience at scale. IEEE Computer Architecture Letters, 20(2):158–161, July/ December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Gupta:2019:DQL

[GMM^{+}19] Ujjwal Gupta, Sumit K. Mandal, Manqing Mao, Chaitali Chakrabarti, and Umit Y. Ogras. A deep Q-learning approach for dynamic management of heterogeneous processors. IEEE Computer Architecture Letters, 18(1):14–17, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gibert:2015:PSR

[GMMC15] Enric Gibert, Raul Martínez, Carlos Madriles, and Josep M. Codina. Profiling support for runtime managed code: Next generation performance monitoring units. IEEE Computer Architecture Letters, 14(1): 62–65, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Garcia-Mallen:2023:TAD

[GMPMC^{+}23] Jonathan Garcia-Mallen, Shuohao Ping, Alex Miralles-Cordal, Ian Martin, Mukund Ramakrishnan, and Yipeng Huang. Towards an accelerator for differential and algebraic equations useful to scientists. IEEE Computer Architecture Letters, 22(2):185– 188, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Gupta:2015:CEO

[GO15] Ujjwal Gupta and Umit Y. Ogras. Constrained energy optimization in heterogeneous platforms using generalized scaling models. IEEE Computer Architecture Letters, 14(1):21–25, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gaudiot:2006:F

[GPS06] J.-L. Gaudiot, Y. Patt, and K. Skadon. Foreword. IEEE Computer Architecture Letters, 5(1):11, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gan:2019:SSV

[GQLZ19] Yiming Gan, Yuxian Qiu, Jingwen Leng, and Yuhao Zhu. SVSoC: Speculative vision systems-on-a-chip. IEEE Computer Architecture Letters, 18(1):47–50, January/ June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gordon-Ross:2002:EFP

[GRCV02] A. Gordon-Ross, S. Cotterell, and F. Vahid. Exploiting fixed programs in embedded systems: a loop cache example. IEEE Computer Architecture Letters, 1(1):2, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gottscho:2017:MIM

[GSG^{+}17] Mark Gottscho, Mohammed Shoaib, Sriram Govindan, Bikash Sharma, Di Wang, and Puneet Gupta. Measuring the impact of memory errors on application performance. IEEE Computer Architecture Letters, 16(1): 51–55, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

GomezRequena:2008:BFT

[GVG^{+}08] C. Gomez Requena, F. Gilabert Villamon, M. Gomez, P. Lopez, and J. Duato. Beyond fat-tree: Unidirectional load–balanced multistage interconnection network. IEEE Computer Architecture Letters, 7(2):49–52, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic). See comment [Ant09].

Golander:2008:DDS

[GWR08] A. Golander, S. Weiss, and R. Ronen. DDMR: Dynamic and scalable dual modular redundancy with short validation intervals. IEEE Computer Architecture Letters, 7 (2):65–68, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Hafezan:2024:IEE

[HA24] Mohammad Hafezan and Ehsan Atoofian. Improving energy-efficiency of capsule networks on modern GPUs. IEEE Computer Architecture Letters, 23(1):49–52, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Hoang:2010:CAN

[HBL^{+}10] Giang Hoang, Chang Bae, John Lange, Lide Zhang, Peter Dinda, and Russ Joseph. A case for alternative nested paging models for virtualized systems. IEEE Computer Architecture Letters, 9(1):17–20, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Hossain:2023:SDA

[HBW^{+}23] Naorin Hossain, Alper Buyuktosunoglu, John-David Wellman, Pradip Bose, and Margaret Martonosi. SoCurity: a design approach for enhancing SoC security. IEEE Computer Architecture Letters, 22 (2):105–108, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Ham:2021:NDP

[HCK^{+}21] Hyungkyu Ham, Hyunuk Cho, Minjae Kim, Jueon Park, Jeongmin Hong, Hyojin Sung, Eunhyeok Park, Euicheol Lim, and Gwangsun Kim. Near-data processing in memory expander for DNN acceleration on GPUs. IEEE Computer Architecture Letters, 20(2):171–174, July/ December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Hong:2022:OMC

[HCK22] Jeongmin Hong, Sungjun Cho, and Gwangsun Kim. Overcoming memory capacity wall of GPUs with heterogeneous memory stack. IEEE Computer Architecture Letters, 21(2):61–64, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Hammoud:2010:DPA

[HCM10] Mohammad Hammoud, Sangyeun Cho, and Rami G. Melhem. A dynamic pressure-aware associative placement strategy for large scale chip multiprocessors. IEEE Computer Architecture Letters, 9(1):29–32, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Hadjilambrou:2018:SCV

[HDAS18] Zacharias Hadjilambrou, Shidhartha Das, Marco A. Antoniades, and Yiannakis Sazeides. Sensing CPU voltage noise through electromagnetic emanations. IEEE Computer Architecture Letters, 17(1): 68–71, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Heirman:2021:RRC

[HEDH21] Wim Heirman, Stijn Eyerman, Kristof Du Bois, and Ibrahim Hur. RIO: ROBcentric in-order modeling of out-of-order processors. IEEE Computer Architecture Letters, 20(1):78–81, January/ June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Huang:2022:EDC

[HH22] Jianming Huang and Yu Hua. Ensuring data confidentiality in eADR-Based NVM systems. IEEE Computer Architecture Letters, 21(2):153– 156, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Hameed:2022:DPA

[HKO^{+}22] Fazal Hameed, Asif Ali Khan, Sebastien Ollivier, Alex K. Jones, and Jeronimo Castrillon. DNA pre-alignment filter using processing near racetrack memory. IEEE Computer Architecture Letters, 21(2):53–56, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Hu:2016:TDM

[HLH16] Qi Hu, Peng Liu, and Michael C. Huang. Threads and data mapping: Affinity analysis for traffic reduction. IEEE Computer Architecture Letters, 15(2):133–136, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Hyun:2021:CAD

[HLR21] Bongjoon Hyun, Jiwon Lee, and Minsoo Rhu. Characterization and analysis of deep learning for 3D point cloud analytics. IEEE Computer Architecture Letters, 20 (2):106–109, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Hashemi:2016:EEB

[HMCP16] Milad Hashemi, Debbie Marr, Doug Carmean, and Yale N. Patt. Efficient execution of bursty applications. IEEE Computer Architecture Letters, 15(2):85–88, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Hoseinzadeh:2018:FBS

[Hos18] Morteza Hoseinzadeh. Flowbased simulation methodology. IEEE Computer Architecture Letters, 17(1):51–54, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Hastings:2023:ASR

[HPS23] Adam Hastings, Ryan Piersma, and Simha Sethumadhavan. Architectural security regulation. IEEE Computer Architecture Letters, 22(2):173– 176, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Hilton:2010:SDE

[HR10] Andrew Hilton and Amir Roth. SMT-Directory: Efficient load-load ordering for SMT. IEEE Computer Architecture Letters, 9(1):25–28, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Hernandez:2011:FTV

[HRF^{+}11] Carles Hernandez, Antoni Roca, Jose Flich, Federico Silla, and Jose Duato. Faulttolerant vertical link design for effective 3D stacking. IEEE Computer Architecture Letters, 10(2):41–44, July/ December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Holloway:2004:CPS

[HS04] A. L. Holloway and G. S.

Sohi. Characterization of problem stores. IEEE Computer Architecture Letters, 3 (1):9, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ho:2011:EIB

[HSUS11] Chen-Han Ho, Garret Staus, Aaron Ulmer, and Karthikeyan Sankaralingam. Exploring the interaction between device lifetime reliability and security vulnerabilities. IEEE Computer Architecture Letters, 10(2):37–40, July/ December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Hou:2022:CUE

[HXL^{+}22] Xiaofeng Hou, Cheng Xu, Jiacheng Liu, Xuehan Tang, Lingyu Sun, Chao Li, and Kwang-Ting Cheng. Characterizing and understanding end-to-end multi-modal neural networks on GPUs. IEEE Computer Architecture Letters, 21(2):125–128, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Im:2020:PBA

[IKW^{+}20] Junsu Im, Hanbyeol Kim, Yumin Won, Jiho Oh, Minjae Kim, and Sungjin Lee. Probability-based address translation for flash SSDs. IEEE Computer Architecture Letters, 19(2):97–100, July/ December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Iqbal:2010:POS

[ILG10] Syed Muhammad Zeeshan Iqbal, Yuchen Liang, and Hakan Grahn. ParMiBench — an open-source benchmark for embedded multiprocessor systems. IEEE Computer Architecture Letters, 9(2):45–48, July/December 2010. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ishii:2020:RIP

[ILNS20] Yasuo Ishii, Jaekyu Lee, Krishnendra Nathella, and Dam Sunwoo. Rebasing instruction prefetching: an industry perspective. IEEE Computer Architecture Letters, 19 (2):147–150, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Ipek:2018:BLL

[ILXY18a] Engin Ipek, Florian Longnos, Shihai Xiao, and Wei Yang. Bit-level load balancing: a new technique for improving the write throughput of deeply scaled STT-MRAM. IEEE Computer Architecture Letters, 17(2):139–142, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ipek:2018:VWC

[ILXY18b] Engin Ipek, Florian Longnos, Shihai Xiao, and Wei Yang. Vertical writes: Closing the throughput gap between deeply scaled STT-MRAM and DRAM. IEEE Computer Architecture Letters, 17 (2):151–154, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ilic:2014:CAR

[IPS14] Aleksandar Ilic, Frederico Pratas, and Leonel Sousa. Cache-aware roofline model: Upgrading the loft. IEEE Computer Architecture Letters, 13(1):21–24, January/ June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Iliakis:2018:DMS

[IXS18] Konstantinos Iliakis, Sotirios Xydis, and Dimitrios Soudris. Decoupled MapReduce for shared-memory multi-core architectures. IEEE Computer Architecture Letters, 17 (2):143–146, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Iliakis:2019:LIG

[IXS19] Konstantinos Iliakis, Sotirios Xydis, and Dimitrios Soudris. LOOG: Improving GPU efficiency with light-weight outof-order execution. IEEE Computer Architecture Letters, 18(2):166–169, July 2019. ISSN 1556-6064.

Jacob:2016:PPT

[Jac16a]  Bruce Jacob.  The 2

PetaFLOP, 3 petabyte, 9 TB/s, 90 kW cabinet: a system architecture for exascale and big data. IEEE Computer Architecture Letters, 15 (2):125–128, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jacob:2016:CVC

[Jac16b] Bruce Jacob. The case for VLIW–CMP as a building block for exascale. IEEE Computer Architecture Letters, 15(1):54–57, January/ June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Judd:2017:SBS

[JAM17] Patrick Judd, Jorge Albericio, and Andreas Moshovos. Stripes: Bit-serial deep neural network computing. IEEE Computer Architecture Letters, 16(1):80–83, January/ June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jeon:2017:CCA

[JC17] Dong-Ik Jeon and Ki-Seok Chung. CasHMC: a cycleaccurate simulator for hybrid memory cube. IEEE Computer Architecture Letters, 16(1):10–13, January/ June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Juang:2002:IDT

[JDK^{+}02] Philo Juang, P. Diodato, S. Kaxiras, K. Skadron, Zhigang Hu, M. Martonosi, and D. W. Clark. Implementing decay techniques using 4T quasi-static memory cells. IEEE Computer Architecture Letters, 1(1):10, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jalili:2022:MPD

[JE22] Majid Jalili and Mattan Erez.

Managing prefetchers with deep reinforcement learning. IEEE Computer Architecture Letters, 21(2):105–108, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Jeon:2019:LAG

[JEAG^{+}19] Hyeran Jeon, Hodjat Asghari Esfeden, Nael B. Abu-Ghazaleh, Daniel Wong, and Sindhuja Elango. Localityaware GPU register file. IEEE Computer Architecture Letters, 18(2):153–156, July 2019. ISSN 1556-6064.

Jin:2022:MPC

[JJP^{+}22] Hoyong Jin, Donghun Jeong, Taewon Park, Jong Hwan Ko, and Jungrae Kim. Multiprediction compression: an efficient and scalable memory compression framework for GP-GPU. IEEE Computer Architecture Letters, 21(2): 37–40, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Jang:2021:DPT

[JKK^{+}21] Yongjoo Jang, Sejin Kim, Daehoon Kim, Sungjin Lee, and Jaeha Kung. Deep partitioned training from nearstorage computing to DNN accelerators. IEEE Computer Architecture Letters, 20(1): 70–73, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Jung:2016:LPS

[JLA16] Daejin Jung, Sheng Li, and Jung Ho Ahn. Large pages on steroids: Small ideas to accelerate big memory applications. IEEE Computer Architecture Letters, 15 (2):101–104, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jeon:2023:HAR

[JLKK23] Kiseok Jeon, Junghee Lee, Bumsoo Kim, and James J. Kim. Hardware accelerated reusable Merkle tree generation for bitcoin blockchain headers. IEEE Computer Architecture Letters, 22(2): 69–72, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Jerger:2007:CSC

[JLP07] N. Enright Jerger, M. Lipasti, and L. Peh. Circuit-switched coherence. IEEE Computer Architecture Letters, 6(1):5–8, January 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jung:2018:PCU

[JLRA18] Daejin Jung, Sunjung Lee, Wonjong Rhee, and Jung Ho Ahn. Partitioning compute units in CNN acceleration for statistical memory traffic shaping. IEEE Computer Architecture Letters, 17(1):72– 75, January/June 2018. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jeong:2023:LLA

[JLS^{+}23] Ipoom Jeong, Jiaqi Lou, Yongseok Son, Yongjoo Park, Yifan Yuan, and Nam Sung Kim. LADIO: Leakage-aware direct I/O for I/O-intensive workloads. IEEE Computer Architecture Letters, 22(2): 77–80, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Joao:2007:DPI

[JMKP07] J. A. Joao, O. Mutlu, H. Kim, and Y. N. Patt. Dynamic predication of indirect jumps. IEEE Computer Architecture Letters, 6(2):25–28, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Joao:2008:DPI

[JMKP08] J. A. Joao, O. Mutlu, H. Kim, and Y. N. Patt. Dynamic predication of indirect jumps. IEEE Computer Architecture Letters, 7(1):1–4, January 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Joo:2013:HPS

[JP13] Yongsoo Joo and Sangsoo Park. A hybrid PRAM and STT–RAM cache architecture for extending the lifetime of PRAM caches. IEEE Computer Architecture Letters, 12(2):55–58, July/ December 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jeon:2018:HMP

[JPC18] Dong-Ik Jeon, Kyeong-Bin Park, and Ki-Seok Chung. HMC-MAC: Processing-in memory architecture for multiply–accumulate operations with hybrid memory cube. IEEE Computer Architecture Letters, 17(1):5–8, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jian:2013:HPE

[JSDK13] Xun Jian, John Sartori, Henry Duwe, and Rakesh Kumar. High performance, energy efficient chipkill correct memory with multidimensional parity. IEEE Computer Architecture Letters, 12(2): 39–42, July/December 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jahanshahi:2020:GNC

[JSLW20] Ali Jahanshahi, Hadi Zamani Sabzi, Chester Lau, and Daniel Wong. GPU-NEST: Characterizing energy efficiency of multi-GPU inference servers. IEEE Computer Architecture Letters, 19 (2):139–142, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Jimenez:2023:LLC

[JTG23] Daniel A. Jiménez, Elvira Teran, and Paul V. Gratz. Last-level cache insertion and promotion policy in the presence of aggressive prefetching. IEEE Computer Architecture Letters, 22(1):17–20, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Jung:2017:NIP

[Jun17] Myoungsoo Jung. NearZero:

an integration of phase change memory with multi-core coprocessor. IEEE Computer Architecture Letters, 16 (2):136–140, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jung:2018:SMS

[JZA^{+}18] Myoungsoo Jung, Jie Zhang, Ahmed Abulila, Miryeong Kwon, Narges Shahidi, John Shalf, Nam Sung Kim, and Mahmut Kandemir. SimpleSSD: Modeling solid state drives for holistic system simulation. IEEE Computer Architecture Letters, 17(1):37– 41, January/June 2018. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2007:FBT

[KBD07] J. Kim, J. Balfour, and W. J. Dally. Flattened butterfly topology for on-chip networks. IEEE Computer Architecture Letters, 6(2):37–40, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kumar:2020:PSM

[KCB^{+}20] Chanchal Kumar, Aayush Chaudhary, Shubham Bhawalkar, Utkarsh Mathur, Saransh Jain, Adith Vastrad, and Eric Rotenberg. Post-silicon microarchitecture. IEEE Computer Architecture Letters, 19(1):26–29, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2019:IGM

[KCP^{+}19] Jiho Kim, Jehee Cha, Jason Jong Kyu Park, Dongsuk Jeon, and Yongjun Park. Improving GPU multitasking efficiency using dynamic resource sharing. IEEE Computer Architecture Letters, 18 (1):1–5, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kaliorakis:2018:SAM

[KCPG18] Manolis Kaliorakis, Athanasios Chatzidimitriou, George Papadimitriou, and Dimitris Gizopoulos. Statistical analysis of multicore CPUs operation in scaled voltage conditions. IEEE Computer Architecture Letters, 17(2):109– 112, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Khoram:2023:EEB

[KDL23] Soroosh Khoram, Kyle Daruwalla, and Mikko Lipasti. Energyefficient Bayesian inference using bitstream computing. IEEE Computer Architecture Letters, 22(1):37–40, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kokkinis:2022:DOC

[KDS22] Argyris Kokkinis, Dionysios Diamantopoulos, and Kostas Siozios. Dynamic optimization of on-chip memories for HLS targeting many-accelerator platforms. IEEE Computer Architecture Letters, 21(2):41–44, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kumar:2003:PPR

[KFJ^{+}03] R. Kumar, K. Farkas, N. P. Jouppi, P. Ranganathan, and D. M. Tullsen. Processor power reduction via single-ISA heterogeneous multi-core architectures. IEEE Computer Architecture Letters, 2 (1):2, January 2003. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2010:LUC

[KG10] Hyungjun Kim and Paul V. Gratz. Leveraging unused cache block words to reduce power in CMP interconnect. IEEE Computer Architecture Letters, 9(1):33–36, January/ June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kondguli:2018:BUS

[KH18] Sushant Kondguli and Michael Huang. Bootstrapping: Using SMT hardware to improve single-thread performance. IEEE Computer Architecture Letters, 17(2):205– 208, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Khan:2019:RCA

[KHB^{+}19] Asif Ali Khan, Fazal Hameed, Robin Bläsing, Stuart Parkin, and Jeronimo Castrillon. RT-Sim: a cycle-accurate simulator for racetrack memories. IEEE Computer Architecture Letters, 18(1):43–46, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2024:ADR

[KHS^{+}24] Yang-Gon Kim, Yun-Ki Han, Jae-Kang Shin, Jun-Kyum Kim, and Lee-Sup Kim. Accelerating deep reinforcement learning via phase-level parallelism for robotics applications. IEEE Computer Architecture Letters, 23(1):41– 44, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2021:DSR

[KJK21] Jiho Kim, Myoungsoo Jung, and John Kim. Decoupled SSD: Reducing data movement on NAND-based flash SSD. IEEE Computer Architecture Letters, 20 (2):150–153, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2019:THA

[KJS^{+}19] S. Kim, H. Jung, W. Shin, H. Lee, and H. Lee. HAD-TWL: Hot address detectionbased wear leveling for phasechange memory systems with low latency. IEEE Computer Architecture Letters, 18 (2):107–110, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kasan:2021:CDB

[KK21] Hans Kasan and John Kim. The case for dynamic bias in global adaptive routing. IEEE Computer Architecture Letters, 20(1):38–41, January/ June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2014:VPT

[KKH14] Daehoon Kim, Hwanju Kim, and Jaehyuk Huh. vCache: Providing a transparent view of the LLC in virtualized environments. IEEE Computer Architecture Letters, 13 (2):109–112, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2022:SSE

[KKJ^{+}22] Sejin Kim, Jungwoo Kim, Yongjoo Jang, Jaeha Kung, and Sungjin Lee. SEMS: Scalable embedding memory system for accelerating embedding-based DNNs. IEEE Computer Architecture Letters, 21(2):157–160, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2013:CFC

[KKK13] Hanjoon Kim, Yonggon Kim, and John Kim. Clumsy flow control for high-throughput bufferless on-chip networks. IEEE Computer Architecture Letters, 12(2):47–50, July/ December 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2018:ZRV

[KKKH18] Seikwon Kim, Wonsang Kwak, Changdae Kim, and Jaehyuk Huh. Zebra refresh: Value transformation for zeroaware DRAM refresh reduction. IEEE Computer Architecture Letters, 17(2):130– 133, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kodakara:2007:CRM

[KKL^{+}07] S. Kodakara, J. Kim, D. Lilja, D. Hawkins, W. Hsu, and P. Yew. CIM: a reliable metric for evaluating program phase classifications. IEEE Computer Architecture Letters, 6 (1):9–12, January 2007. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2015:PEM

[KKL^{+}15] Seung Hun Kim, Dohoon Kim, Changmin Lee, Won Seob Jeong, Won Woo Ro, and Jean-Luc Gaudiot. A performance-energy model to evaluate single thread execution acceleration. IEEE Computer Architecture Letters, 14 (2):99–102, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2020:TSA

[KKL20] Minsub Kim, Jaeha Kung, and Sungjin Lee. Towards scalable analytics with inference-enabled solid-state drives. IEEE Computer Architecture Letters, 19(1):13– 17, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2023:HAC

[KKL^{+}23] Yonghae Kim, Anurag Kar, Jaewon Lee, Jaekyu Lee, and Hyesoon Kim. Hardwareassisted code-pointer tagging for forward-edge control-flow integrity. IEEE Computer Architecture Letters, 22(2):117– 120, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2022:CSD

[KKLL22] Minjae Kim, Bryan S. Kim, Eunji Lee, and Sungjin Lee. A case study of a DRAM-NVM hybrid memory allocator for key–value stores. IEEE Computer Architecture Letters, 21(2):81–84, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2018:SPM

[KKP^{+}18] Jihun Kim, Joonsung Kim, Pyeongsu Park, Jong Kim, and Jangwoo Kim. SSD performance modeling using bottleneck analysis. IEEE Computer Architecture Letters, 17(1):80–83, January/ June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

KleinOsowski:2002:MNS

[KL02] A. J. KleinOsowski and D. J.

Lilja. MinneSPEC: a new SPEC benchmark workload for simulation-based computer architecture research. IEEE Computer Architecture Letters, 1(1):7, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2018:HBP

[KL18] Chinam Kim and Hyukjun Lee. A high-bandwidth PCMbased memory system for highly available IP routing table lookup. IEEE Computer Architecture Letters, 17 (2):246–249, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2021:RSD

[KLCA21] Hweesoo Kim, Sunjung Lee, Jaewan Choi, and Jung Ho Ahn. Row-streaming dataflow using a chaining buffer and systolic array+ structure.

IEEE Computer Architecture Letters, 20(1):34–37, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2014:SGA

[KLKK14] Youngsok Kim, Jaewon Lee, Donggyu Kim, and Jangwoo Kim. ScaleGPU: GPU architecture for memory-unaware GPU programming. IEEE Computer Architecture Letters, 13(2):101–104, July/ December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2024:FAD

[KLR24] Hyeseong Kim, Yunjae Lee, and Minsoo Rhu. FPGAaccelerated data preprocessing for personalized recommendation systems. IEEE Computer Architecture Letters, 23(1):7–10, January/ June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Khan:2011:DDC

[KLSD11] Omer Khan, Mieszko Lis, Yildiz Sinangil, and Srinivas Devadas. DCC: a dependable cache coherence multicore architecture. IEEE Computer Architecture Letters, 10(1): 12–15, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2021:ZCS

[KLWJ21] Sunghwan Kim, Gyusun Lee, Jiwon Woo, and Jinkyu Jeong. Zero-copying I/O stack for low-latency SSDs. IEEE Computer Architecture Letters, 20(1):50–53, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kong:2012:ASF

[KLZ12] Ji Kong, Peilin Liu, and Yu Zhang. Atomic streaming: a framework of on-chip data supply system for taskparallel MPSoCs. IEEE Computer Architecture Letters, 11 (1):5–8, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kline:2018:CAR

[KMJ18] Donald Kline, Jr., Rami Melhem, and Alex K. Jones. Counter advance for reliable encryption in phase change memory. IEEE Computer Architecture Letters, 17(2):209– 212, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kvatinsky:2014:MBM

[KNE^{+}14] Shahar Kvatinsky, Yuval H.

Nacson, Yoav Etsion, Eby G. Friedman, Avinoam Kolodny, and Uri C. Weiser. Memristorbased multithreading. IEEE Computer Architecture Letters, 13(1):41–44, January/ June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Khatamifard:2018:MSD

[KNG^{+}18] S. Karen Khatamifard, M. Hassan Najafi, Ali Ghoreyshi, Ulya R. Karpuzcu, and David J. Lilja. On memory system design for stochastic computing. IEEE Computer Architecture Letters, 17 (2):117–121, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kang:2015:SRT

[KNGK15] Suk Chan Kang, Chrysostomos Nicopoulos, Ada Gavrilovska, and Jongman Kim. Subtleties of run-time virtual address stacks. IEEE Computer Architecture Letters, 14 (2):152–155, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2015:ASM

[KNQ15] Dae-Hyun Kim, Prashant J. Nair, and Moinuddin K. Qureshi. Architectural support for mitigating row hammering in DRAM memories. IEEE Computer Architecture Letters, 14(1):9–12, January/ June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kalani:2021:ICB

[KP21] Neelu Shivprakash Kalani and Biswabandan Panda. Instruction criticality based energyefficient hardware data prefetching. IEEE Computer Architecture Letters, 20(2):146– 149, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Krimer:2010:SNT

[KPEC10] Evgeni Krimer, Robert Pawlowski, Mattan Erez, and Patrick Chiang. Synctium: a nearthreshold stream processor for energy-constrained parallel applications. IEEE Computer Architecture Letters, 9(1):21– 24, January/June 2010. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kang:2020:NPP

[KPKK20] Ki-Dong Kang, Gyeongseo Park, Nam Sung Kim, and Daehoon Kim. Network packet processing modeaware power management for data center servers. IEEE Computer Architecture Letters, 19(1):1–4, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2021:TTR

[KPL^{+}21] Byeongho Kim, Jaehyun Park, Eojin Lee, Minsoo Rhu, and Jung Ho Ahn. TRiM: Tensor reduction in memory. IEEE Computer Architecture Letters, 20(1):5–8, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kwon:2021:FQM

[KPPK21] Hyoukjun Kwon, Michael Pellauer, Angshuman Parashar, and Tushar Krishna. Flexion: a quantitative metric for flexibility in DNN accelerators. IEEE Computer Architecture Letters, 20(1):1–4, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Kulkarni:2018:LAI

[KQD18] Neeraj Kulkarni, Feng Qi, and Christina Delimitrou. Leveraging approximation to improve datacenter resource efficiency. IEEE Computer Architecture Letters, 17(2):171– 174, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kannan:2016:EAP

[KQGS16] Sudarsun Kannan, Moinudin Qureshi, Ada Gavrilovska, and Karsten Schwan. Energy aware persistence: Reducing the energy overheads of persistent memory. IEEE Computer Architecture Letters, 15(2):89–92, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kwon:2018:CMC

[KR18] Youngeun Kwon and Minsoo Rhu. A case for memorycentric HPC system architecture for training deep neural networks. IEEE Computer Architecture Letters, 17 (2):134–138, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Karsli:2013:EDT

[KRB^{+}13] I. Burak Karsli, Pedro Reviriego, M. Fatih Balli, Oğuz Ergin, and J. A. Maestro. Enhanced duplication: a technique to correct soft errors in narrow values. IEEE Computer Architecture Letters, 12(1):13–16, January/ June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kumar:2019:HRA

[KSB19] Chanchal Kumar, Sidharth Singh, and Gregory T. Byrd. Hybrid remote access protocol. IEEE Computer Architecture Letters, 18(1):30–33, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kleanthous:2016:TML

[KSO^{+}16] Marios Kleanthous, Yiannakis Sazeides, Emre Ozer, Chrysostomos Nicopoulos, Panagiota Nikolaou, and Zacharias Hadjilambrou. Toward multi-layer holistic evaluation of system designs.

IEEE Computer Architecture Letters, 15(1):58–61, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Krishnan:2020:SLV

[KWB^{+}20] Srivatsan Krishnan, Zishen Wan, Kshitij Bhardwaj, Paul Whatmough, Aleksandra Faust, Gu-Yeon Wei, David Brooks, and Vijay Janapa Reddi. The sky is not the limit: a visual performance model for cyber-physical co-design in autonomous machines. IEEE Computer Architecture Letters, 19(1):38–42, January/ June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Khatamifard:2018:NCC

[KWKK18] S. Karen Khatamifard, Longfei Wang, Selcuk Köse, and Ulya R. Karpuzcu. A new class of covert channels exploiting power management vulnerabilities. IEEE Computer Architecture Letters, 17 (2):201–204, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kai:2013:GRP

[KWL13] Yi Kai, Yi Wang, and Bin Liu. GreenRouter: Reducing power by innovating router’s architecture. IEEE Computer Architecture Letters, 12(2): 51–54, July/December 2013.

CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Khan:2017:CMC

[KWL^{+}17] Samira Khan, Chris Wilkerson, Donghyuk Lee, Alaa R. Alameldeen, and Onur Mutlu. A case for memory contentbased detection and mitigation of data-dependent failures in DRAM. IEEE Computer Architecture Letters, 16(2):88–93, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2016:RFE

[KYM16] Yoongu Kim, Weikun Yang, and Onur Mutlu. Ramulator: a fast and extensible DRAM simulator. IEEE Computer Architecture Letters, 15(1):45–49, January/ June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kumar:2021:DDS

[KYP21] Pratik Kumar, Chavhan Sujeet Yashavant, and Biswabandan Panda. DAMARU: a denial-of-service attack on randomized last-level caches. IEEE Computer Architecture Letters, 20(2):138–141, July/ December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Khoram:2018:AAA

[KZL18] Soroosh Khoram, Yue Zha, and Jing Li. An alternative analytical approach to associative processing. IEEE Computer Architecture Letters, 17(2):113–116, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kommareddy:2019:CMS

[KZY^{+}19] Vamsee Reddy Kommareddy, Baogang Zhang, Fan Yao, Rickard Ewetz, and Amro Awad. Are crossbar memories secure? New security vulnerabilities in crossbar memories. IEEE Computer Architecture Letters, 18(2):174– 177, July 2019. ISSN 1556-6064.

Li:2016:ICL

[LA16] Dongdong Li and Tor M. Aamodt. Inter-core locality aware memory scheduling. IEEE Computer Architecture Letters, 15(1):25–28, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lavasani:2014:FBL

[LAC14] Maysam Lavasani, Hari Angepat, and Derek Chiou. An FPGA-based in-line accelerator for Memcached. IEEE Computer Architecture Letters, 13(2):57–60, July/ December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lucas:2022:LHI

[LAM^{+}22] Benjamin J. Lucas, Ali Alwan, Marion Murzello, Yazheng Tu, Pengzhou He, Andrew J. Schwartz, David Guevara, Ujjwal Guin, Kyle Juretus, and Jiafeng Xie. Lightweight hardware implementation of binary ring-LWE PQC accelerator. IEEE Computer Architecture Letters, 21(1):17–20, January/ June 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lenjani:2022:PAH

[LAS22] Marzieh Lenjani, Alif Ahmed, and Kevin Skadron. Pulley: an algorithm/hardware cooptimization for in-memory sorting. IEEE Computer Architecture Letters, 21(2):109– 112, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lazarev:2020:DTE

[LAX^{+}20] Nikita Lazarev, Neil Adit, Shaojie Xiang, Zhiru Zhang, and Christina Delimitrou. Dagger: Towards efficient RPCs in cloud microservices with near-memory reconfigurable NICs. IEEE Computer Architecture Letters, 19 (2):134–138, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Leng:2019:ARA

[LBB^{+}19] J. Leng, A. Buyuktosunoglu, R. Bertran, P. Bose, and V. J. Reddi. Asymmetric resilience for accelerator-rich systems. IEEE Computer Architecture Letters, 18(1):83– 86, January/June 2019. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lai:2020:TDB

[LCHL20] Bo-Cheng Lai, Chun-Yen Chen, Yi-Da Hsin, and Bo-Yen Lin. A two-directional BigData sorting architecture on FPGAs. IEEE Computer Architecture Letters, 19(1): 72–75, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2023:HFT

[LCKA23] Hailong Li, Jaewan Choi, Yongsuk Kwon, and Jung Ho Ahn. A hardware-friendly tiled singular-value decompositionbased matrix multiplication for transformer-based models. IEEE Computer Architecture Letters, 22(2):169–172, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2016:PHP

[LCW^{+}16] Minghua Li, Guancheng Chen, Qijun Wang, Yonghua Lin, Peter Hofstee, Per Stenstrom, and Dian Zhou. PA-Ter: a hardware prefetching automatic tuner on IBM POWER8 processor. IEEE Computer Architecture Letters, 15(1):37–40, January/ June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lachtar:2020:CSA

[LEBM20] Nada Lachtar, Abdulrahman Abu Elkhail, Anys Bacha, and Hafiz Malik. A cross-stack approach towards defending against cryptojacking. IEEE Computer Architecture Letters, 19(2):126– 129, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2020:CDE

[LG20] Congmiao Li and Jean-Luc Gaudiot. Challenges in detecting an evasive spectre. IEEE Computer Architecture Letters, 19(1):18–21, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2017:FFE

[LGLK17] Junghee Lee, Kalidas Ganesh, Hyuk-Jun Lee, and Youngjae Kim. FESSD: a fast encrypted SSD employing onchip access-control memory. IEEE Computer Architecture Letters, 16(2):115–118, July/ December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2022:MES

[LHCK22] Dusol Lee, Duwon Hong, Wonil Choi, and Jihong Kim. MQSim-E: an enterprise SSD simulator. IEEE Computer Architecture Letters, 21(1): 13–16, January/June 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Liu:2021:SMS

[LHE^{+}21] Wenjie Liu, Wim Heirman, Stijn Eyerman, Shoaib Akram, and Lieven Eeckhout. Scale-model simulation. IEEE Computer Architecture Letters, 20(2):175–178, July/ December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2023:HHF

[LHPR23] Seonho Lee, Ranggi Hwang, Jongse Park, and Minsoo Rhu. HAMMER: Hardwarefriendly approximate computing for self-attention with mean-redistribution and linearization. IEEE Computer Architecture Letters, 22(1): 13–16, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lyons:2010:ASF

[LHWB10] Michael J. Lyons, Mark Hempstead, Gu-Yeon Wei, and David Brooks. The accelerator store framework for high-performance, lowpower accelerator-based systems. IEEE Computer Architecture Letters, 9(2):53–56, July/December 2010. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Liu:2019:UFT

[LHZ19] He Liu, Jianhui Han, and Youhui Zhang. A unified framework for training, mapping and simulation of ReRAM-based convolutional neural network acceleration. IEEE Computer Architecture Letters, 18(1):63–66, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Luo:2004:EES

[LJ04] Yue Luo and L. K. John. Efficiently evaluating speedup using sampled processor simulation. IEEE Computer Architecture Letters, 3(1):6, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lou:2018:BSB

[LJ18] Qian Lou and Lei Jiang.

BRAWL: a spintronics-based portable basecalling-in-memory architecture for nanopore genome sequencing. IEEE Computer Architecture Letters, 17(2):239–242, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Liu:2014:PTE

[LJM^{+}14] Qixiao Liu, Victor Jimenez, Miquel Moreto, Jaume Abella, Francisco J. Cazorla, and Mateo Valero. Per-task energy accounting in computing systems. IEEE Computer Architecture Letters, 13(2):85– 88, July/December 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2015:RDA

[LKA15] Sungjin Lee, Jihong Kim, and Arvind. Refactored design of I/O architecture for flash storage. IEEE Computer Architecture Letters, 14(1):70– 74, January/June 2015. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2019:ELM

[LKK19] Seunghak Lee, Nam Sung Kim, and Daehoon Kim. Exploiting OS-level memory offlining for DRAM power management. IEEE Computer Architecture Letters, 18(2):141– 144, July 2019. ISSN 1556-6064.

Lee:2015:SSI

[LKKS15] Junghee Lee, Youngjae Kim, Jongman Kim, and Galen M. Shipman. Synchronous I/O scheduling of independent write caches for an array of SSDs. IEEE Computer Architecture Letters, 14(1):79–82, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2021:LPM

[LKL^{+}21] Hyeon Gyu Lee, Minwook Kim, Juwon Lee, Eunji Lee, Bryan S. Kim, Sungjin Lee, Yeseong Kim, Sang Lyul Min, and Jin-Soo Kim. Learned performance model for SSD. IEEE Computer Architecture Letters, 20(2):154–157, July/ December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2023:NPR

[LKP^{+}23] Seunghak Lee, Ki-Dong Kang, Gyeongseo Park, Nam Sung Kim, and Daehoon Kim. No-Hammer: Preventing row hammer with last-level cache management. IEEE Computer Architecture Letters, 22 (2):157–160, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2021:UIN

[LKR21] Yunjae Lee, Youngeun Kwon, and Minsoo Rhu. Understanding the implication of non-volatile memory for large-scale graph neural network training. IEEE Computer Architecture Letters, 20 (2):118–121, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2018:BSB

[LLD^{+}18] Zhaoshi Li, Leibo Liu, Yangdong Deng, Shouyi Yin, and Shaojun Wei. Breaking the synchronization bottleneck with reconfigurable transactional execution. IEEE Computer Architecture Letters, 17 (2):147–150, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2023:CDC

[LLJK23] Hwanjun Lee, Seunghak Lee, Yeji Jung, and Daehoon Kim. T-CAT: Dynamic cache allocation for tiered memory systems with memory interleaving. IEEE Computer Architecture Letters, 22(2):73–76, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lakshminarayana:2012:DSP

[LLKS12] Nagesh B. Lakshminarayana, Jaekyu Lee, Hyesoon Kim, and Jinwoo Shin. DRAM scheduling policy for GPGPU architectures based on a potential function. IEEE Computer Architecture Letters, 11(2):33–36, July/December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2006:ASC

[LLLM06] Moon-Sang Lee, Sang-Kwon Lee, Joonwon Lee, and Seung-Ryoul Maeng. Adopting system call based address translation into user-level communication. IEEE Computer Architecture Letters, 5(1):26–29, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2021:RRA

[LLM^{+}21] Qian Li, Bin Li, Pietro Mercati, Ramesh Illikkal, Charlie Tai, Michael Kishinevsky, and Christos Kozyrakis. RAMBO: Resource allocation for microservices using Bayesian optimization. IEEE Computer Architecture Letters, 20(1): 46–49, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lin:2019:DSE

[LLPC19] Ting-Ru Lin, Yunfan Li, Massoud Pedram, and Lizhong Chen. Design space exploration of memory controller placement in throughput processors with deep learning. IEEE Computer Architecture Letters, 18(1):51–54, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Liu:2015:LHP

[LLS^{+}15] Longjun Liu, Chao Li, Hongbin Sun, Yang Hu, Jingmin Xin, Nanning Zheng, and Tao Li. Leveraging heterogeneous power for improving datacenter efficiency and resiliency. IEEE Computer Architecture Letters, 14(1):41– 45, January/June 2015. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2018:TTW

[LLSA18] Eojin Lee, Sukhan Lee, G. Edward Suh, and Jung Ho Ahn. TWiCe: Time window counter based row refresh to prevent row-hammering.

IEEE Computer Architecture Letters, 17(1):96–99, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Luque:2009:CAC

[LMC^{+}09] Carlos Luque, Miquel Moreto, Francisco J. Cazorla, Roberto Gioiosa, Alper Buyuktosunoglu, and Mateo Valero. CPU accounting in CMP processors. IEEE Computer Architecture Letters, 8(1):17–20, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2012:LSS

[LMJ12] Yong Li, Rami Melhem, and Alex K. Jones. Leveraging sharing in second level translation-lookaside buffers for chip multiprocessors. IEEE Computer Architecture Letters, 11(2):49–52, July/ December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2006:PBH

[LMK06] W. Li, S. Mohanty, and K. Kavi. A page-based hybrid (software–hardware) dynamic memory allocator. IEEE Computer Architecture Letters, 5(2):13, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Leverich:2009:PMD

[LMT^{+}09] Jacob Leverich, Matteo Monchiero, Vanish Talwar, Partha Ranganathan, and Christos Kozyrakis. Power management of datacenter workloads using percore power gating. IEEE Computer Architecture Letters, 8(2):48–51, July/December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lai:2016:QMD

[LPK16] Bo-Cheng Charles Lai, Luis Garrido Platero, and Hsien-Kai Kuo. A quantitative method to data reuse patterns of SIMT applications. IEEE Computer Architecture Letters, 15(2):73–76, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lu:2023:FEA

[LQYF23] Lingfei Lu, Yudi Qiu, Shiyan Yi, and Yibo Fan. A flexible embedding-aware near memory processing architecture for recommendation system. IEEE Computer Architecture Letters, 22(2):165–168, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2019:PBP

[LSJ^{+}19] Chen Li, Yifan Sun, Lingling Jin, Lingjie Xu, Zheng Cao, Pengfei Fan, David Kaeli, Sheng Ma, Yang Guo, and Jun Yang. Priority-based PCIe scheduling for multitenant multi-GPU systems. IEEE Computer Architecture Letters, 18(2):157–160, July 2019. ISSN 1556-6064.

Lo:2023:LLV

[LTL23] Yun-Chen Lo, Yu-Chih Tsai, and Ren-Shuo Liu. LV: Latency-versatile floatingpoint engine for high-performance deep neural networks. IEEE Computer Architecture Letters, 22(2):125–128, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lyons:2013:SFF

[LWB13] Michael Lyons, Gu-Yeon Wei, and David Brooks. Shrink-Fit: a framework for flexible accelerator sizing. IEEE Computer Architecture Letters, 12(1):17–20, January/ June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Liu:2020:STA

[LWM20] Zhi-Gang Liu, Paul N. Whatmough, and Matthew Mattina. Systolic tensor array: an efficient structured-sparse GEMM accelerator for mobile CNN inference. IEEE Computer Architecture Letters, 19(1):34–37, January/ June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2008:PDD

[LX08] J. Lee and X. Xiao. A parallel deadlock detection algorithm with O(1) overall runtime complexity. IEEE Computer Architecture Letters, 7 (2):45–48, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Liang:2016:CGR

[LYL^{+}16] Shuang Liang, Shouyi Yin, Leibo Liu, Yike Guo, and Shaojun Wei. A coarsegrained reconfigurable architecture for compute-intensive MapReduce acceleration. IEEE Computer Architecture Letters, 15(2):69–72, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2020:DCA

[LYR^{+}20] Shang Li, Zhiyuan Yang, Dhiraj Reddy, Ankur Srivastava, and Bruce Jacob. DRAMsim3: a cycle-accurate, thermal-capable DRAM simulator. IEEE Computer Architecture Letters, 19(2):106– 109, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2021:HAG

[LYY^{+}21] Han Li, Mingyu Yan, Xiaocheng Yang, Lei Deng, Wenming Li, Xiaochun Ye, Dongrui Fan, and Yuan Xie. Hardware acceleration for GCNs via bidirectional fusion. IEEE Computer Architecture Letters, 20(1):66– 4, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Liu:2023:ILG

[LZD^{+}23] Jie Liu, Zhongyuan Zhao, Zijian Ding, Benjamin Brock, Hongbo Rong, and Zhiru Zhang. An intermediate language for general sparse format customization. IEEE Computer Architecture Letters, 22(2):153–156, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2020:SFA

[LZL^{+}20] Joo Hwan Lee, Hui Zhang, Veronica Lagrange, Praveen Krishnamoorthy, Xiaodong Zhao, and Yang Seok Ki. SmartSSD: FPGA accelerated near-storage data analytics on SSD. IEEE Computer Architecture Letters, 19 (2):110–113, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Liao:2015:AWL

[LZLX15] Jianwei Liao, Fengxiang Zhang, Li Li, and Guoqiang Xiao. Adaptive wear-leveling in flash-based memory. IEEE Computer Architecture Letters, 14(1):1–4, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2008:TAN

[LZS^{+}08] Z. Li, C. Zhu, L. Shang, R. Dick, and Y. Sun. Transaction-aware networkon-chip resource reservation. IEEE Computer Architecture Letters, 7(2):53–56, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Marinakis:2019:PFI

[MA19] Theodoros Marinakis and Iraklis Anagnostopoulos. Performance and fairness improvement on CMPs considering bandwidth and cache utilization. IEEE Computer Architecture Letters, 18(2):1–4, July 2019. ISSN 1556-6064.

Min:2018:SCD

[MAHK18] Seungwon Min, Mohammad Alian, Wen-Mei Hwu, and Nam Sung Kim. Semicoherent DMA: an alternative I/O coherency management for embedded systems. IEEE Computer Architecture Letters, 17(2):221–224, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Manohar:2015:CSD

[Man15] Rajit Manohar. Comparing stochastic and deterministic computing. IEEE Computer Architecture Letters, 14 (2):119–122, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Martinez:2013:MNE

[Mar13a] J. Martinez. A message from the new Editor-in-Chief and introduction of new Associate editors. IEEE Computer Architecture Letters, 12(1):2–4, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Martinez:2013:E

[Mar13b] J. F. Martinez. Editorial. IEEE Computer Architecture Letters, 12(2):37–38, July/ December 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mirhosseini:2017:SPB

[MAT17] Amirhossein Mirhosseini, Aditya Agrawal, and Josep Torrellas. Survive: Pointerbased in-DRAM incremental checkpointing for low-cost data persistence and rollbackrecovery. IEEE Computer Architecture Letters, 16(2):153– 157, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Manevich:2010:CAR

[MCKW10] Ran Manevich, Israel Cidon, Avinoam Kolodny, and Isask’har Walter. Centralized adaptive routing for NoCs. IEEE Computer Architecture Letters, 9(2):57–60, July/December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Maddah:2013:DDS

[MCM13] Rakan Maddah, Sangyeun Cho, and Rami Melhem. Data dependent sparing to manage better-than-bad blocks. IEEE Computer Architecture Letters, 12(2):43–46, July/ December 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

MoretoPlanas:2007:EDC

[MCRV07] M. Moreto Planas, F. Cazorla, A. Ramirez, and M. Valero. Explaining dynamic cache partitioning speed ups. IEEE Computer Architecture Letters, 6(1):1–4, January 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Meza:2012:EES

[MCY^{+}12] Justin Meza, Jichuan Chang, HanBin Yoon, Onur Mutlu, and Parthasarathy Ranganathan. Enabling efficient and scalable hybrid memories using fine-granularity DRAM cache management. IEEE Computer Architecture Letters, 11(2):61–64, July/ December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Moon:2023:AAD

[MDK^{+}23] Yaebin Moon, Wanju Doh, Kwanhee Kyung, Eojin Lee, and Jung Ho Ahn. ADT: Aggressive demotion and promotion for tiered memory. IEEE Computer Architecture Letters, 22(1):21–24, January/ June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Mason:2020:UPI

[MDSG20] Tony Mason, Thaleia Dimitra Doudali, Margo Seltzer, and Ada Gavrilovska. Unexpected performance of Intel Optane DC persistent memory. IEEE Computer Architecture Letters, 19(1):55–58, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Ma:2022:FBA

[MGH^{+}22] Rui Ma, Evangelos Georganas, Alexander Heinecke, Sergey Gribok, Andrew Boutros, and Eriko Nurvitadhi. FPGA-Based AI smart NICs for scalable distributed AI training systems. IEEE Computer Architecture Letters, 21(2):49–52, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Mirosanlou:2020:MED

[MGHP20] Reza Mirosanlou, Danlu Guo, Mohamed Hassan, and Rodolfo Pellizzoni. MCsim: an extensible DRAM memory controller simulator. IEEE Computer Architecture Letters, 19(2):105–109, July/ December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Martinsen:2014:HTL

[MGI14] Jan Kasper Martinsen, Hakan Grahn, and Anders Isberg. Heuristics for thread-level speculation in Web applications. IEEE Computer Architecture Letters, 13(2):77– 80, July/December 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mohammadi:2015:DDB

[MHAD15] Milad Mohammadi, Song Han, Tor M. Aamodt, and William J. Dally. On-demand dynamic branch prediction. IEEE Computer Architecture Letters, 14(1):50–53, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ma:2024:PFA

[MHM^{+}24] Rui Ma, Jia-Ching Hsu, Ali Mansoorshahi, Joseph Garvey, Michael Kinsner, Deshanand Singh, and Derek Chiou. Primate: a framework to automatically generate soft processors for network applications. IEEE Computer Architecture Letters, 23(1):57– 60, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Michaud:2013:DMT

[Mic13] Pierre Michaud. Demystifying multicore throughput metrics. IEEE Computer Architecture Letters, 12(2):63– 66, July/December 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Michaud:2020:ETT

[Mic20] Pierre Michaud. Exploiting thermal transients with deterministic turbo clock frequency. IEEE Computer Architecture Letters, 19(1):43– 46, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Michelogiannakis:2011:PCE

[MJBD11]  George  Michelogiannakis,

Nan Jiang, Daniel U. Becker, and William J. Dally. Packet chaining: Efficient singlecycle allocation for on-chip networks. IEEE Computer Architecture Letters, 10(2): 33–36, July/December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mars:2023:JPP

[MKD^{+}23] Jason Mars, Yiping Kang, Roland Daynauth, Baichuan Li, Ashish Mahendra, Krisztian Flautner, and Lingjia Tang. The Jaseci programming paradigm and runtime stack: Building scale-out production applications easy and fast. IEEE Computer Architecture Letters, 22(2):101– 104, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Marquez:2017:MCH

[MKM17] David Gonzalez Marquez, Adrian Cristal Kestelman, and Esteban Mocskos. Mth: Codesigned hardware/software support for fine grain threads. IEEE Computer Architecture Letters, 16(1):64–67, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mosquera:2023:GCC

[MKMJ23] Fernando Mosquera, Krishna Kavi, Gayatri Mehta, and Lizy John. Guard cache: Creating noisy side-channels. IEEE Computer Architecture Letters, 22(2):97–100, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Mishra:2024:ASA

[MKP^{+}24] Deepanjali Mishra, Konstantinos Kanellopoulos, Ashish Panwar, Akshitha Sriraman, Vivek Seshadri, Onur Mutlu, and Todd C. Mowry. Address scaling: Architectural support for fine-grained threadsafe metadata management. IEEE Computer Architecture Letters, 23(1):69–72, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Mutlu:2005:RRP

[MKSP05] O. Mutlu, Hyesoon Kim, J. Stark, and Y. N. Patt. On reusing the results of pre-executed instructions in a runahead execution processor. IEEE Computer Architecture Letters, 4(1):2, January 2005. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mahmoodi:2014:RCC

[MLA^{+}14] Hamid Mahmoodi, Sridevi Srinivasan Lakshmipuram, Manish Arora, Yashar Asgarieh, Houman Homayoun, Bill Lin, and Dean M. Tullsen. Resistive computation: a critique. IEEE Computer Architecture Letters, 13(2):89–92, July/ December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Manatunga:2015:HSS

[MLK15] Dilan Manatunga, Joo Hwan Lee, and Hyesoon Kim. Hardware support for safe execution of native client applications. IEEE Computer Architecture Letters, 14(1):37–40, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mallik:2006:UDF

[MLM^{+}06] A. Mallik, B. Lin, G. Memik, P. Dinda, and R. P. Dick. User-driven frequency scaling. IEEE Computer Architecture Letters, 5(2):16, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Milenkovic:2003:SBT

[MM03] A. Milenkovic and M. Milenkovic.

Stream-based trace compression. IEEE Computer Architecture Letters, 2(1):4, January 2003. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Munoz-Martinez:2021:SEC

[MMAAK21] Francisco Muñoz-Martínez, José L. Abellán, Manuel E. Acacio, and Tushar Krishna. STONNE: Enabling cyclelevel microarchitectural simulation for DNN inference accelerators. IEEE Computer Architecture Letters, 20 (2):122–125, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Martinez:2017:SII

[MMR17] Jorge A. Martínez, Juan Antonio Maestro, and Pedro Reviriego. A scheme to improve the intrinsic error detection of the instruction set architecture. IEEE Computer Architecture Letters, 16 (2):103–106, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Morad:2014:GMO

[MMY^{+}14] Amir Morad, Tomer Y.

Morad, Leonid Yavits, Ran Ginosar, and Uri Weiser. Generalized MultiAmdahl: Optimization of heterogeneous multi-accelerator SoC.

IEEE Computer Architecture Letters, 13(1):37–40, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mikhailenko:2020:ASA

[MNFI20] Darya Mikhailenko, Yujin Nakamoto, Ben Feinberg, and Engin Ipek. Adapting in situ accelerators for sparsity with granular matrix reordering. IEEE Computer Architecture Letters, 19(2):143–146, July/ December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Markovic:2015:TLS

[MNU^{+}15] Nikola Markovic, Daniel Nemirovsky, Osman Unsal, Mateo Valero, and Adrian Cristal. Thread lock sectionaware scheduling on asymmetric single-ISA multi-core. IEEE Computer Architecture Letters, 14(2):160–163, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Min:2018:AAB

[MPA^{+}18] Donghyun Min, Donggyu Park, Jinwoo Ahn, Ryan Walker, Junghee Lee, Sungyong Park, and Youngjae Kim. Amoeba: an autonomous backup and recovery SSD for ransomware attack defense. IEEE Computer Architecture Letters, 17(2):243–246, July/ December 2018. CODEN ????

ISSN 1556-6056 (print), 1556-6064 (electronic).

Manivannan:2017:RAG

[MPPS17] Madhavan Manivannan, Miquel Pericàs, Vassilis Papaefstathiou, and Per Stenström. Runtime-assisted global cache management for task-based parallel programs. IEEE Computer Architecture Letters, 16(2):145–148, July/ December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Maycock:2016:HES

[MS16] Matthew Maycock and Simha Sethumadhavan. Hardware enforced statistical privacy. IEEE Computer Architecture Letters, 15(1):21–24, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Matsuo:2019:IIF

[MSA19] Reoma Matsuo, Ryota Shioya, and Hideki Ando. Improving the instruction fetch throughput with dynamically configuring the fetch pipeline. IEEE Computer Architecture Letters, 18(2):170–173, July 2019. ISSN 1556-6064.

Morad:2017:ORO

[MSE^{+}17] Tomer Y. Morad, Gil Shomron, Mattan Erez, Avinoam Kolodny, and Uri C. Weiser. Optimizing read-once data flow in big-data applications. IEEE Computer Architecture Letters, 16(1):68–71, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mashimo:2018:VMS

[MSI18] Susumu Mashimo, Ryota Shioya, and Koji Inoue. VMOR: Microarchitectural support for operand access in an interpreter. IEEE Computer Architecture Letters, 17(2):217– 220, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mars:2011:HHW

[MTH11] Jason Mars, Lingjia Tang, and Robert Hundt. Heterogeneity in “Homogeneous” warehouse-scale computers: a performance opportunity. IEEE Computer Architecture Letters, 10(2):29–32, July/ December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mouris:2018:TSB

[MTM18] Dimitris Mouris, Nektarios Georgios Tsoutsos, and Michail Maniatakos. TER-Minator suite: Benchmarking privacy-preserving architectures. IEEE Computer Architecture Letters, 17(2):122– 125, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Miller:2012:MEP

[MTT12] Timothy N. Miller, Renji Thomas, and Radu Teodorescu. Mitigating the effects of process variation in ultra-low voltage chip multiprocessors using dual supply voltages and half-speed units. IEEE Computer Architecture Letters, 11(2):45– 48, July/December 2012. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Musoll:2009:PVA

[Mus09] Enric Musoll. A processvariation aware technique for tile-based, massive multicore processors. IEEE Computer Architecture Letters, 8(2):52– 55, July/December 2009. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mittal:2015:ATE

[MV15] Sparsh Mittal and Jeffrey S.

Vetter. AYUSH: a technique for extending lifetime of SRAM–NVM hybrid caches. IEEE Computer Architecture Letters, 14(2):115–118, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mittal:2017:ARD

[MVJ17] Sparsh Mittal, Jeffrey S. Vetter, and Lei Jiang. Addressing read-disturbance issue in STT–RAM by data compression and selective duplication.

IEEE Computer Architecture Letters, 16(2):94–98, July/ December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Morad:2006:PPE

[MWK^{+}06] T. Y. Morad, U. C. Weiser, A. Kolodnyt, M. Valero, and E. Ayguade. Performance, power efficiency and scalability of asymmetric cluster chip multiprocessors. IEEE Computer Architecture Letters, 5 (1):14–17, January 2006. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Masouros:2019:RRS

[MXS19] D. Masouros, S. Xydis, and D. Soudris. Rusty: Runtime system predictability leveraging LSTM neural networks. IEEE Computer Architecture Letters, 18(2):103–106, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Naghibijouybari:2017:CCG

[NAG17] Hoda Naghibijouybari and Nael Abu-Ghazaleh. Covert channels on GPGPUs. IEEE Computer Architecture Letters, 16(1):22–25, January/ June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Nagabhiru:2024:AFP

[NB24] Mahita Nagabhiru and Gregory T. Byrd. Achieving forward progress guarantee in small hardware transactions. IEEE Computer Architecture Letters, 23(1):53– 56, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Nilakantan:2013:MES

[NBH13] Siddharth Nilakantan, Steven Battle, and Mark Hempstead. Metrics for early-stage modeling of many-accelerator architectures. IEEE Computer Architecture Letters, 12(1):25– 28, January/June 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Nam:2023:XRD

[NBW^{+}23] Hwayong Nam, Seungmin Baek, Minbok Wi, Michael Jaemin Kim, Jaehyun Park, Chihun Song, Nam Sung Kim, and Jung Ho Ahn. X-ray: Discovering DRAM internal structure and error characteristics by issuing memory commands. IEEE Computer Architecture Letters, 22(2): 89–92, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Naithani:2019:PRE

[NFAE19] Ajeya Naithani, Josue Feliu, Almutaz Adileh, and Lieven Eeckhout. Precise runahead execution. IEEE Computer Architecture Letters, 18(1): 71–74, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Nowatzki:2015:GBP

[NGS15] Tony Nowatzki, Venkatraman Govindaraju, and Karthikeyan Sankaralingam. A graphbased program representation for analyzing hardware specialization approaches. IEEE Computer Architecture Letters, 14(2):94–98, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Nabavinejad:2019:CDP

[NHKR19] Seyed Morteza Nabavinejad, Hassan Hafez-Kolahi, and Sherief Reda. Coordinated DVFS and precision control for deep neural networks. IEEE Computer Architecture Letters, 18(2):136–140, July 2019. ISSN 1556-6064.

Nye:2022:SSS

[NK22] Jared Nye and Omer Khan. SSE: Security service engines to accelerate enclave performance in secure multicore processors. IEEE Computer Architecture Letters, 21 (2):129–132, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Nandakumar:2014:OKS

[NMS14] Vivek S. Nandakumar and Małgorzata Marek-Sadowska. On optimal kernel size for integrated CPU–GPUs — a case study. IEEE Computer Architecture Letters, 13(2): 81–84, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Nematallah:2023:ELS

[NPBS23] Ahmed Nematallah, Chang Hyun Park, and David Black-Schaffer. Exploring the latency sensitivity of cache replacement policies. IEEE Computer Architecture Letters, 22(2):93–96, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Navarro:2021:HSS

[NPS21] Marta Navarro, Lucia Pons, and Julio Sahuquillo. Hy-Sched: a simple hyperthreadingaware thread to core allocation strategy. IEEE Computer Architecture Letters, 20(1): 26–29, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Nabavinejad:2021:BLB

[NR21] Seyed Morteza Nabavinejad and Sherief Reda. BayesTuner: Leveraging Bayesian optimization for DNN inference configuration selection. IEEE Computer Architecture Letters, 20(2):166–170, July/ December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Nathan:2015:AGC

[NS15] Ralph Nathan and Daniel J. Sorin. Argus-G: Comprehensive, low-cost error detection for GPGPU cores. IEEE Computer Architecture Letters, 14(1):13–16, January/ June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Newton:2020:PGP

[NSC20] Newton, Virendra Singh, and Trevor E. Carlson. PIM-GraphSCC: PIM-based graph processing using graph’s community structures. IEEE Computer Architecture Letters, 19(2):151–154, July/ December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Nematollahi:2018:NSD

[NSF^{+}18] Negin Nematollahi, Mohammad Sadrosadati, Hajar Falahati, Marzieh Barkhordar, and Hamid Sarbazi-Azad. Neda: Supporting direct inter-core neighbor data exchange in GPUs. IEEE Computer Architecture Letters, 17 (2):225–229, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Omar:2018:MRI

[ODKK18] Hamza Omar, Halit Dogan, Brian Kahne, and Omer Khan. Multicore resource isolation for deterministic, resilient and secure concurrent execution of safety-critical applications. IEEE Computer Architecture Letters, 17 (2):230–234, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Omori:2022:OSH

[OK22] Yu Omori and Keiji Kimura.

Open-source hardware memory protection engine integrated with NVMM simulator. IEEE Computer Architecture Letters, 21(2):77–80, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

O:2015:CCI

[OKS^{+}15] Seongil O, Sanghyuk Kwon, Young Hoon Son, Yujin Park, and Jung Ho Ahn. CIDR: a cache inspired area-efficient DRAM resilience architecture against permanent faults. IEEE Computer Architecture Letters, 14(1):17–20, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ottoni:2006:SPC

[ORS^{+}06] G. Ottoni, R. Rangan, A. Stoler, M. J. Bridges, and D. I. August. From sequential programs to concurrent threads. IEEE Computer Architecture Letters, 5(1):6–9, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Olson:2016:SIT

[OSH16] Lena E. Olson, Simha Sethumadhavan, and Mark D. Hill. Security implications of thirdparty accelerators. IEEE Computer Architecture Letters, 15(1):50–53, January/ June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Panda:2016:EPP

[PB16] Biswabandan Panda and Shankar Balachandran. Expert prefetch prediction: an expert predicting the usefulness of hardware prefetchers. IEEE Computer Architecture Letters, 15(1):13–16, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Pekhimenko:2015:TAC

[PBO^{+}15] Gennady Pekhimenko, Evgeny Bolotin, Mike O’Connor, Onur Mutlu, Todd C. Mowry, and Stephen W. Keckler. Toggle-aware compression for GPUs. IEEE Computer Architecture Letters, 14(2):164– 168, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Pu:2016:NIP

[PDGV16] Libei Pu, Kshitij Doshi, Ellis Giles, and Peter Varman. Non-intrusive persistence with a backend NVM controller. IEEE Computer Architecture Letters, 15(1): 29–32, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Perais:2021:CSS

[Per21] Arthur Perais. A case for speculative strength reduction. IEEE Computer Architecture Letters, 20(1):22– 25, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Piccolboni:2022:ASS

[PGC22] Luca Piccolboni, Davide Giri, and Luca P. Carloni. Accelerators & security: The socket approach. IEEE Computer Architecture Letters, 21(2): 65–68, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Panda:2012:BFB

[PGJ12] Reena Panda, Paul V. Gratz, and Daniel A. Jiménez. B-Fetch: Branch prediction directed prefetching for in-order processors. IEEE Computer Architecture Letters, 11(2): 41–44, July/December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Patel:2023:TIP

[PGR^{+}23] Pratyush Patel, Zibo Gong, Syeda Rizvi, Esha Choukse, Pulkit Misra, Thomas Anderson, and Akshitha Sriraman. Towards improved power management in cloud GPUs. IEEE Computer Architecture Letters, 22(2):141– 144, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Pham:2018:TSM

[PHBC18] Binh Pham, Derek Hower, Abhishek Bhattacharjee, and Trey Cain. TLB shootdown mitigation for lowpower many-core servers with L1 virtual caches. IEEE Computer Architecture Letters, 17(1):17–20, January/ June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Power:2015:GGH

[PHO^{+}15] Jason Power, Joel Hestness, Marc S. Orr, Mark D. Hill, and David A. Wood. gem5gpu: a heterogeneous CPU– GPU simulator. IEEE Computer Architecture Letters, 14(1):34–36, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Park:2022:SML

[PJ22] Jongwon Park and Jinkyu Jeong. Speculative multilevel access in LSM treebased KV store. IEEE Computer Architecture Letters, 21 (2):145–148, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Park:2023:CEE

[PKKK23] Gyeongseo Park, Ki-Dong Kang, Minho Kim, and Daehoon Kim. CoreNap: Energy efficient core allocation for latency-critical workloads. IEEE Computer Architecture Letters, 22(1):1–4, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Patil:2010:URT

[PL10] Shruti Patil and David J.

Lilja. Using resampling techniques to compute confidence intervals for the harmonic mean of rate-based performance metrics. IEEE Computer Architecture Letters, 9 (1):1–4, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Poluri:2015:SET

[PL15] Pavan Poluri and Ahmed Louri. A soft error tolerant network-on-chip router pipeline for multi-core systems. IEEE Computer Architecture Letters, 14(2):107– 110, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Park:2023:DHP

[PLK^{+}23] Chanyoung Park, Chun-Yi Liu, Kyungtae Kang, Mahmut Kandemir, and Wonil Choi. Design of a highperformance, high-endurance key-value SSD for large-key workloads. IEEE Computer Architecture Letters, 22 (2):149–152, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Pao:2008:PAM

[PLL08] D. Pao, W. Lin, and B. Liu. Pipelined architecture for multi-string matching. IEEE Computer Architecture Letters, 7(2):33–36, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Piscitelli:2012:HLP

[PP12] Roberta Piscitelli and Andy D. Pimentel. A high-level power model for MPSoC on FPGA. IEEE Computer Architecture Letters, 11(1):13–16, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Prieto:2011:MCM

[PPG11] Pablo Prieto, Valentin Puente, and Jose-Angel Gregorio. Multilevel cache modeling for chip-multiprocessor systems. IEEE Computer Architecture Letters, 10(2):49–52, July/ December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Pinto:2017:TTA

[PPG^{+}17] Sandro Pinto, Jorge Pereira, Tiago Gomes, Mongkol Ekpanyapong, and Adriano Tavares. Towards a TrustZoneassisted hypervisor for realtime embedded systems. IEEE Computer Architecture Letters, 16(2):158–161, July/ December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Perais:2017:SFM

[PS17] Arthur Perais and Andre Seznec. Storage-free memory dependency prediction. IEEE Computer Architecture Letters, 16(2):149–152, July/ December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Peltekis:2024:DDM

[PTND24] Christodoulos Peltekis, Vasileios Titopoulos, Chrysostomos Nicopoulos, and Giorgos Dimitrakopoulos. DeMM: a decoupled matrix multiplication engine supporting relaxed structured sparsity. IEEE Computer Architecture Letters, 23(1):17–20, January/ June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Price:2006:CCT

[PV06] G. Price and M. Vachharajani. A case for compressing traces with BDDs. IEEE Computer Architecture Letters, 5(2):18, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Poremba:2015:NUF

[PZX15] Matthew Poremba, Tao Zhang, and Yuan Xie. NVMain 2.0: a user-friendly memory simulator to model (non-) volatile memory systems. IEEE Computer Architecture Letters, 14(2):140– 143, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ringlein:2023:ACD

[RAD^{+}23] Burkhard Ringlein, Francois Abel, Dionysios Diamantopoulos, Beat Weiss, Christoph Hagleitner, and Dietmar Fey. Advancing compilation of DNNs for FPGAs using operation set architectures. IEEE Computer Architecture Letters, 22(1):9–12, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Rao:2019:ATC

[RADZ19] J. Rao, T. Ao, K. Dai, and X. Zou. ARCE: Towards code pointer integrity on embedded processors using architectureassisted run-time metadata management. IEEE Computer Architecture Letters, 18 (2):115–118, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Ranganath:2019:SCC

[RASW19] Kiran Ranganath, AmirAli Abdolrashidi, Shuaiwen Leon Song, and Daniel Wong. Speeding up collective communications through inter-GPU re-routing. IEEE Computer Architecture Letters, 18 (2):128–131, July 2019. ISSN 1556-6064.

Rotem:2014:BUI

[RB14] Nadav Rotem and Yosi Ben Asher. Block unification IFconversion for high performance architectures. IEEE Computer Architecture Letters, 13(1):17–20, January/ June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Rosenfeld:2011:DCA

[RCBJ11] Paul Rosenfeld, Elliott Cooper-Balis, and Bruce Jacob. DRAMSim2: a cycle accurate memory system simulator. IEEE Computer Architecture Letters, 10(1):16–19, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Resch:2021:CPC

[RCK21] Salonik Resch, Husrev Cilasun, and Ulya R. Karpuzcu. Cryogenic PIM: Challenges opportunities. IEEE Computer Architecture Letters, 20(1):74–77, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Rodopoulos:2015:TPV

[RCS15] Dimitrios Rodopoulos, Francky Catthoor, and Dimitrios Soudris. Tackling performance variability due to RAS mechanisms with PIDcontrolled DVFS. IEEE Computer Architecture Letters, 14 (2):156–159, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ros:2020:EIP

[RJ20] Alberto Ros and Alexandra Jimborean. The entangling instruction prefetcher. IEEE Computer Architecture Letters, 19(2):84–87, July/ December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Resch:2022:VSQ

[RK22] Salonik Resch and Ulya Karpuzcu. On variable strength quantum ECC. IEEE Computer Architecture Letters, 21(2):93–96, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Ramanujam:2008:RPM

[RL08] R. Sunkam Ramanujam and B. Lin. Randomized partiallyminimal routing on threedimensional mesh networks. IEEE Computer Architecture Letters, 7(2):37–40, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ramanujam:2009:WRR

[RL09] Rohit Sunkam Ramanujam and Bill Lin. Weighted random routing on torus networks. IEEE Computer Architecture Letters, 8(1):1–4, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ravi:2017:TSM

[RL17] Gokul Subramanian Ravi and Mikko Lipasti. Timing speculation in multi-cycle data paths. IEEE Computer Architecture Letters, 16(1):84–87, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Rakshit:2018:LLO

[RM18] Joydeep Rakshit and Kartik Mohanram. LEO: Low overhead encryption ORAM for non-volatile memories. IEEE Computer Architecture Letters, 17(2):100–104, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Rezaei:2020:NNM

[RMA^{+}20] Seyyed Hossein SeyyedAghaei Rezaei, Mehdi Modarressi, Rachata Ausavarungnirun, Mohammad Sadrosadati, Onur Mutlu, and Masoud Daneshtalab. NoM: Network-on-memory for inter-bank data transfer in highly-banked memories. IEEE Computer Architecture Letters, 19(1):

80–83, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Rezaei:2016:DRS

[RMMLK16] Seyyed Hossein Seyyedaghaei Rezaei, Abbas Mazloumi, Mehdi Modarressi, and Pejman Lotfi-Kamran. Dynamic resource sharing for highperformance 3-D networkson-chip. IEEE Computer Architecture Letters, 15(1):5–8, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Roth:2008:PRR

[Rot08] A. Roth. Physical register reference counting. IEEE Computer Architecture Letters, 7 (1):9–12, January 2008. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Rucker:2021:CTB

[RSO21] Alexander Rucker, Muhammad Shahbaz, and Kunle Olukotun. Chopping off the tail: Bounded nondeterminism for real-time accelerators. IEEE Computer Architecture Letters, 20 (2):110–113, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Rogers:2019:SLB

[RSRT19] Samuel Rogers, Joshua Slycord, Ronak Raheja, and Hamed Tabkhi. Scalable LLVM-based accelerator modeling in gem5. IEEE Computer Architecture Letters, 18(1):18–21, January/ June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Resch:2021:DLQ

[RTKQ21] Salonik Resch, Swamit Tannu, Ulya R. Karpuzcu, and Moinuddin Qureshi. A day in the life of a quantum error. IEEE Computer Architecture Letters, 20(1):13–16, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Robatmili:2004:TSI

[RYSN04] B. Robatmili, N. Yazdani, S. Sardashti, and M. Nourani. Thread-sensitive instruction issue for SMT processors. IEEE Computer Architecture Letters, 3(1):5, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Riley:2006:PCU

[RZ06] N. Riley and C. Zilles. Probabilistic counter updates for predictor hysteresis and bias. IEEE Computer Architecture Letters, 5(1):18–21, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sim:2023:CCM

[SAA^{+}23] Joonseop Sim, Soohong Ahn, Taeyoung Ahn, Seungyong Lee, Myunghyun Rhee, Jooyoung Kim, Kwangsik Shin, Donguk Moon, Euiseok Kim, and Kyoung Park. Computational CXL-Memory solution for accelerating memoryintensive applications. IEEE Computer Architecture Letters, 22(1):5–8, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Shwartz:2018:DMI

[SB18] Ofir Shwartz and Yitzhak Birk. Distributed memory integrity trees. IEEE Computer Architecture Letters, 17 (2):159–162, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shoghi:2021:SSQ

[SBQK21] Nima Shoghi, Andrei Bersatti, Moinuddin Qureshi, and Hyesoon Kim. SmaQ: Smart quantization for DNN training by exploiting value clustering. IEEE Computer Architecture Letters, 20(2):126– 129, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Sasaki:2017:MPC

[SBVB17] Hiroshi Sasaki, Alper Buyuktosunoglu, Augusto Vega, and Pradip Bose. Mitigating power contention: a scheduling based approach. IEEE Computer Architecture Letters, 16(1):60–63, January/ June 2017. CODEN ????

ISSN 1556-6056 (print), 1556-6064 (electronic).

Sutradhar:2020:PPP

[SCB^{+}20] Purab Ranjan Sutradhar, Mark Connolly, Sathwika Bavikadi, Sai Manoj Pudukotai Dinakarrao, Mark A. Indovina, and Amlan Ganguly. pPIM: a programmable processor-in-memory architecture with precision-scaling for deep learning. IEEE Computer Architecture Letters, 19 (2):118–121, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Stine:2004:CAR

[SCF04] J. M. Stine, N. P. Carter, and J. Flich. Comparing adaptive routing and dynamic voltage scaling for link power reduction. IEEE Computer Architecture Letters, 3(1):4, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Srinivasan:2006:PMU

[SCL06] R. Srinivasan, J. Cook, and O. Lubeck. Performance modeling using Monte Carlo simulation. IEEE Computer Architecture Letters, 5(1):38–41, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sun:2013:NWC

[SCL13] Guang Sun, Chia-Wei Chang, and Bill Lin. A new worstcase throughput bound for oblivious routing in odd radix mesh network. IEEE Computer Architecture Letters, 12 (1):9–12, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Son:2017:SAS

[SCR^{+}17] Young Hoon Son, Hyunyoon Cho, Yuhwan Ro, Jae W. Lee, and Jung Ho Ahn. SALAD: Achieving symmetric access latency with asymmetric DRAM architecture. IEEE Computer Architecture Letters, 16(1):76–79, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shaw:2002:MSC

[SD02] K. A. Shaw and W. J. Dally. Migration in single chip multiprocessors. IEEE Computer Architecture Letters, 1(1):12, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Singh:2004:BDB

[SD04] A. Singh and W. J. Dally. Buffer and delay bounds in high radix interconnection networks. IEEE Computer Architecture Letters, 3(1):8, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Singh:2004:GAL

[SDTG04] A. Singh, W. J. Dally, B. Towles, and A. K.

Gupta. Globally adaptive load-balanced routing on tori. IEEE Computer Architecture Letters, 3(1):2, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Seznec:2010:PCM

[Sez10]  Andre Seznec.  A phase

change memory as a secure main memory. IEEE Computer Architecture Letters, 9 (1):5–8, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sendag:2003:ACE

[SfCL03] R. Sendag, Peng fei Chuang, and D. J. Lilja. Address correlation: Exceeding the limits of locality. IEEE Computer Architecture Letters, 2 (1):3, January 2003. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Stow:2019:PPM

[SFFG^{+}19] Dylan Stow, Amin Farmahini-Farahani, Sudhanva Gurumurthi, Michael Ignatowski, and Yuan Xie. Power profiling of modern die-stacked memory. IEEE Computer Architecture Letters, 18(2):132– 135, July 2019. ISSN 1556-6064.

Sankar:2014:SFL

[SG14] Sriram Sankar and Sudhanva Gurumurthi. Soft failures in large datacenters. IEEE Computer Architecture Letters, 13(2):105–108, July/ December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

SanMiguel:2018:EMA

[SGBE18] Joshua San Miguel, Karthik Ganesan, Mario Badr, and Natalie Enright Jerger. The EH model: Analytical exploration of energy-harvesting architectures. IEEE Computer Architecture Letters, 17(1):76–79, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Seshadri:2015:FBB

[SHB^{+}15] Vivek Seshadri, Kevin Hsieh, Amirali Boroum, Donghyuk Lee, Michael A. Kozuch, Onur Mutlu, Phillip B. Gibbons, and Todd C. Mowry. Fast bulk bitwise AND and OR in DRAM. IEEE Computer Architecture Letters, 14 (2):127–131, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Su:2021:EPA

[SHJW21] Jiya Su, Linfeng He, Peng Jiang, and Rujia Wang. Exploring PIM architecture for high-performance graph pattern mining. IEEE Computer Architecture Letters, 20 (2):114–117, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Shi:2015:CLM

[SHK15] Qingchuan Shi, Henry Hoffmann, and Omer Khan. A cross-layer multicore architecture to tradeoff program accuracy and resilience overheads. IEEE Computer Architecture Letters, 14(2):85– 89, July/December 2015. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shomron:2019:SSS

[SHW19] G. Shomron, T. Horowitz, and U. Weiser. SMT-SA: Simultaneous multithreading in systolic arrays. IEEE Computer Architecture Letters, 18 (2):99–102, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Salvesen:2022:LAR

[SJ22] Peter Salvesen and Magnus Jahre. LMT: Accurate and resource-scalable slowdown prediction. IEEE Computer Architecture Letters, 21(2): 97–100, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Song:2017:EPU

[SJA^{+}17] Wonjun Song, Hyung-Joon Jung, Jung Ho Ahn, Jae W. Lee, and John Kim. Evaluation of performance unfairness in NUMA system architecture. IEEE Computer Architecture Letters, 16(1):26– 29, January/June 2017. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sohn:2002:RRE

[SJM02] YoungChul Sohn, NaiHoon Jung, and Seungryoul Maeng. Request reordering to enhance the performance of strict consistency models. IEEE Computer Architecture Letters, 1(1):11, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Seyedzadeh:2017:CBT

[SJM17] Seyed Mohammad Seyedzadeh, Alex K. Jones, and Rami Melhem. Counter-based tree structure for row hammering mitigation in DRAM. IEEE Computer Architecture Letters, 16(1):18–21, January/ June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shan:2021:ACP

[SK21] Mohsin Shan and Omer Khan. Accelerating concurrent priority scheduling using adaptive in-hardware task distribution in multicores. IEEE Computer Architecture Letters, 20(1):17– 21, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Skadron:2009:LE

[Ska09a] K. Skadron. Letter from the Editor. IEEE Computer Architecture Letters, 8(2):39, July/December 2009. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Skadron:2009:U

[Ska09b] Kevin Skadron. Untitled.

IEEE Computer Architecture Letters, 8(2):39, July/ December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Skadron:2010:ELE

[Ska10a] K. Skadron. Editorial: Letter from the Editor-in-Chief. IEEE Computer Architecture Letters, 9(2):37–44, July/ December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Skadron:2010:U

[Ska10b] Kevin Skadron. Untitled.

IEEE Computer Architecture Letters, 9(2):37–44, July/ December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Skadron:2011:ELE

[Ska11a] K. Skadron. Editorial: Letter from the Editor-in-Chief. IEEE Computer Architecture Letters, 10(1):1–3, January/ June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Skadron:2011:U

[Ska11b] Kevin Skadron. Untitled.

IEEE Computer Architecture Letters, 10(1):1–3, January/ June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Skadron:2013:INE

[Ska13] Kevin Skadron. Introducing the new Editor-in-Chief of the IEEE Computer Architecture Letters. IEEE Computer Architecture Letters, 12(1):1, January/June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sartor:2020:HHL

[SKA^{+}20] Anderson L. Sartor, Anish Krishnakumar, Samet E. Arda, Umit Y. Ogras, and Radu Marculescu. HiLITE: Hierarchical and lightweight imitation learning for power management of embedded SoCs. IEEE Computer Architecture Letters, 19(1):63– 67, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Sudarsanam:2009:PPD

[SKD09] Arvind Sudarsanam, Ramachandra Kallam, and Aravind Dasu. PRR–PRR dynamic relocation. IEEE Computer Architecture Letters, 8(2):44–47, July/December 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shin:2022:OOS

[SKK22] Gyeongcheol Shin, Junsoo Kim, and Joo-Young Kim.

OpenMDS: an open-source shell generation framework for high-performance design on Xilinx multi-die FPGAs. IEEE Computer Architecture Letters, 21(2):101–104, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Seongil:2015:CCI

[SKS^{+}15] O. Seongil, Sanghyuk Kwon, Young Hoon Son, Yujin Park, and Jung Ho Ahn. CIDR: a cache inspired area-efficient DRAM resilience architecture against permanent faults. IEEE Computer Architecture Letters, 14(1):17–20, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sazeides:2005:DIB

[SKTC05] Y. Sazeides, R. Kumar, D. M.

Tullsen, and T. Constantinou. The danger of intervalbased power efficiency metrics: When worst is best. IEEE Computer Architecture Letters, 4(1):1, January 2005. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sihn:2003:SCS

[SLC03] K.-H. Sihn, Joonwon Lee, and Jung-Wan Cho. A speculative coherence scheme using decoupling synchronization for multiprocessor systems. IEEE Computer Architecture Letters, 2(1):1, January 2003. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shim:2014:TMP

[SLKD14] Keun Sup Shim, Mieszko Lis, Omer Khan, and Srinivas Devadas. Thread migration prediction for distributed shared caches. IEEE Computer Architecture Letters, 13(1):53– 56, January/June 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Swami:2018:AAS

[SM18] Shivam Swami and Kartik Mohanram. ARSENAL: Architecture for secure nonvolatile memories. IEEE Computer Architecture Letters, 17(2):192–196, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Seo:2015:DDF

[SMLS15] Bon-Keun Seo, Seungryoul Maeng, Joonwon Lee, and Euiseong Seo. DRACO: a deduplicating FTL for tangible extra capacity. IEEE Computer Architecture Letters, 14(2):123–126, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Song:2015:ARL

[SMY15] William Song, Saibal Mukhopadhyay, and Sudhakar Yalamanchili. Architectural reliability: Lifetime reliability characterization and management of many-core processors. IEEE Computer Architecture Letters, 14(2):103–106, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Scionti:2018:EMM

[SMZ18] Alberto Scionti, Somnath Mazumdar, and Stephane Zuckerman. Enabling massive multi-threading with fast hashing. IEEE Computer Architecture Letters, 17(1):1–4, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Subramoni:2010:ISI

[SPAP10] Hari Subramoni, Fabrizio Petrini, Virat Agarwal, and Davide Pasetto. Intra-socket and inter-socket communication in multi-core systems. IEEE Computer Architecture Letters, 9(1):13–16, January/ June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shin:2022:RSA

[SPHS22] Yongwon Shin, Juseong Park, Jeongmin Hong, and Hyojin Sung. Runtime support for accelerating CNN models on digital DRAM processing-in-memory hardware. IEEE Computer Architecture Letters, 21(2):33–36, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Shang:2002:PEI

[SPJ02] Li Shang, L. Peh, and N. K. Jha. Power-efficient interconnection networks: Dynamic voltage scaling with links. IEEE Computer Architecture Letters, 1(1):6, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Saileshwar:2023:MBM

[SQ23] Gururaj Saileshwar and Moinuddin Qureshi. The mirage of breaking MIRAGE: Analyzing the modeling pitfalls in emerging attacks on MI-RAGE. IEEE Computer Architecture Letters, 22(2):121– 124, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Sharifi:2020:AAC

[SRH20] Ferdous Sharifi, Nezam Rohbani, and Shaahin Hessabi. Aging-aware context switching in multicore processors based on workload classification. IEEE Computer Architecture Letters, 19 (2):159–162, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Singh:2020:VLB

[SRLM20] Rahul Singh, Gokul Subramanian Ravi, Mikko Lipasti, and Joshua San Miguel. Value locality based approximation with ODIN. IEEE Computer Architecture Letters, 19(2): 88–91, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Soteriou:2009:HTD

[SRLP09] Vassos Soteriou, Rohit Sunkam Ramanujam, Bill Lin, and Li-Shiuan Peh. A highthroughput distributed sharedbuffer NoC router. IEEE Computer Architecture Letters, 8(1):21–24, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Siozios:2011:SRT

[SRS11] Kostas Siozios, Dimitrios Rodopoulos, and Dimitrios Soudris. On supporting rapid thermal analysis. IEEE Computer Architecture Letters, 10(2):53–56, July/December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sadredini:2020:ESP

[SRS20] Elaheh Sadredini, Reza Rahimi, and Kevin Skadron. Enabling in-SRAM pattern processing with low-overhead reporting architecture. IEEE Computer Architecture Letters, 19 (2):167–170, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Sethumadhavan:2012:CHD

[SRT12] Simha Sethumadhavan, Ryan Roberts, and Yannis Tsividis. A case for hybrid discrete-continuous architectures. IEEE Computer Architecture Letters, 11(1):1–4, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sadredini:2019:SEM

[SRV^{+}19] E. Sadredini, R. Rahimi, V. Verma, M. Stan, and K. Skadron. A scalable and efficient in-memory interconnect architecture for automata processing. IEEE Computer Architecture Letters, 18(2):87–90, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Shah:2021:TDS

[SSS^{+}21] Parth Shah, Ranjal Gautham Shenoy, Vaidyanathan Srinivasan, Pradip Bose, and Alper Buyuktosunoglu. TokenSmart: Distributed, scalable power management in the many-core era. IEEE Computer Architecture Letters, 20(1):42–45, January/ June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Sahoo:2018:RRD

[SSSM18] Debiprasanna Sahoo, Swaraj Sha, Manoranjan Satpathy, and Madhu Mutyam. ReDRAM: a reconfigurable DRAM cache for GPGPUs. IEEE Computer Architecture Letters, 17(2):213–216, July/ December 2018. CODEN ????

ISSN 1556-6056 (print), 1556-6064 (electronic).

Sasaki:2017:HTP

[SSTS17] Hiroshi Sasaki, Fang-Hsiang Su, Teruo Tanimoto, and Simha Sethumadhavan. Heavy tails in program structure. IEEE Computer Architecture Letters, 16(1):34–37, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sarkar:2021:DDA

[SSVS21] Arindam Sarkar, Newton Singh, Varun Venkitaraman, and Virendra Singh. DAM: Deadblock aware migration techniques for STT-RAM-based hybrid caches. IEEE Computer Architecture Letters, 20(1):62–4, January/ June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Samara:2020:CDS

[ST20] Ahmed Samara and James Tuck. The case for domainspecialized branch predictors for graph-processing. IEEE Computer Architecture Letters, 19(2):101–104, July/ December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Sen:2016:GFM

[SW16] Rathijit Sen and David A.

Wood. GPGPU footprint models to estimate per-core power. IEEE Computer Architecture Letters, 15(2): 97–100, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shomron:2019:SCV

[SW19] Gil Shomron and Uri Weiser. Spatial correlation and value prediction in convolutional neural networks. IEEE Computer Architecture Letters, 18(1):10–13, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sendag:2007:BMP

[SYC07] R. Sendag, J. Yi, and P. Chuang. Branch misprediction prediction: Complementary branch predictors. IEEE Computer Architecture Letters, 6(2):49–52, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Song:2014:AFB

[SYC14] Xiang Song, Jian Yang, and Haibo Chen. Architecting flash-based solid-state drive for high-performance I/O virtualization. IEEE Computer Architecture Letters, 13(2): 61–64, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tavakkol:2013:NSS

[TASA13] Arash Tavakkol, Mohammad Arjomand, and Hamid Sarbazi-Azad. Network-on-SSD: a scalable and highperformance communication design paradigm for SSDs. IEEE Computer Architecture Letters, 12(1):5–8, January/ June 2013. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Towles:2002:WCT

[TD02] B. Towles and W. J. Dally.

Worst-case traffic for oblivious routing functions. IEEE Computer Architecture Letters, 1(1):4, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tomusk:2016:DDG

[TDO16] Erik Tomusk, Christophe Dubach, and Michael O’Boyle. Diversity: a design goal for heterogeneous processors. IEEE Computer Architecture Letters, 15(2):81–84, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tang:2011:PEM

[TLG^{+}11] Jie Tang, Shaoshan Liu, Zhimin Gu, Chen Liu, and Jean-Luc Gaudiot. Prefetching in embedded mobile systems can be energy-efficient. IEEE Computer Architecture Letters, 10(1):8–11, January/ June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tovletoglou:2019:SIH

[TMNK19] Konstantinos Tovletoglou, Lev Mukhanov, Dimitrios S. Nikolopoulos, and Georgios Karakonstantis. Shimmer: Implementing a heterogeneousreliability DRAM framework on a commodity server. IEEE Computer Architecture Letters, 18(1):26–29, January/ June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tavakkol:2016:TTB

[TMSA16] Arash Tavakkol, Pooyan Mehrvarzy, and Hamid Sarbazi-Azad. TBM: Twin block management policy to enhance the utilization of plane-level parallelism in SSDs. IEEE Computer Architecture Letters, 15 (2):121–124, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tzenetopoulos:2023:DLD

[TMSX23] Achilleas Tzenetopoulos, Dimosthenis Masouros, Dimitrios Soudris, and Sotirios Xydis. DVFaaS: Leveraging DVFS for FaaS workflows. IEEE Computer Architecture Letters, 22(2):85–88, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Tan:2019:DWO

[TNC19] Tian Tan, Eriko Nurvitadhi, and Derek Chiou. Dark wires and the opportunities for reconfigurable logic. IEEE Computer Architecture Letters, 18(1):67–70, January/ June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tan:2021:FQF

[TND^{+}21] Tian Tan, Eriko Nurvitadhi, Aravind Dasu, Martin Langhammer, and Derek Chiou. FlexScore: Quantifying flexibility. IEEE Computer Architecture Letters, 20(1):58– 4, January/June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Tanimoto:2017:EDG

[TOIS17] Teruo Tanimoto, Takatsugu Ono, Koji Inoue, and Hiroshi Sasaki. Enhanced dependence graph model for critical path analysis on modern out-of-order processors. IEEE Computer Architecture Letters, 16 (2):111–114, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tambat:2002:PLB

[TV02] S. Tambat and S. Vajapeyam.

Page-level behavior of cache contention. IEEE Computer Architecture Letters, 1(1):9, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Tembey:2013:SSS

[TVB^{+}13] Priyanka Tembey, Augusto Vega, Alper Buyuktosunoglu, Dilma Da Silva, and Pradip Bose. SMT switch: Software mechanisms for power shifting. IEEE Computer Architecture Letters, 12(2):67–70, July/December 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Thomas:2024:BMT

[TWI^{+}24] Samuel Thomas, Kidus Workneh, Ange-Thierry Ishimwe, Zack McKevitt, Phaedra Curlin, R. Iris Bahar, Joseph Izraelevitz, and Tamara Lehman. Baobab Merkle tree for efficient secure memory. IEEE Computer Architecture Letters, 23(1):33–36, January/ June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Trochatos:2023:QCT

[TXD^{+}23] Theodoros Trochatos, Chuanqi Xu, Sanjay Deshpande, Yao Lu, Yongshan Ding, and Jakub Szefer. A quantum computer trusted execution environment. IEEE Computer Architecture Letters, 22 (2):177–180, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Unsal:2002:CFC

[UKM02] O. S. Unsal, C. M. Krishna, and C. A. Mositz. Cool-Fetch: Compiler-enabled power-aware fetch throttling. IEEE Computer Architecture Letters, 1(1):5, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ueno:2024:ITB

[UTT^{+}24] Yosuke Ueno, Yuna Tomida, Teruo Tanimoto, Masamitsu Tanaka, Yutaka Tabuchi, Koji Inoue, and Hiroshi Nakamura. Inter-temperature bandwidth reduction in cryogenic QAOA machines. IEEE Computer Architecture Letters, 23(1):9–12, January/ June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Vandierendonck:2002:ATC

[VD02] H. Vandierendonck and K. De Bosschere. An address transformation combining blockand word-interleaving. IEEE Computer Architecture Letters, 1(1):8, January 2002. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

VandenSteen:2018:MSP

[VE18] Sam Van den Steen and Lieven Eeckhout. Modeling superscalar processor memory-level parallelism. IEEE Computer Architecture Letters, 17(1):9–12, January/ June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Vakil-Ghahani:2018:CRP

[VGMSLN^{+}18] Armin Vakil-Ghahani, Sara Mahdizadeh-Shahri, Mohammad-Reza Lotfi-Namin, Mohammad Bakhshalipour, Pejman Lotfi-Kamran, and Hamid Sarbazi-Azad. Cache replacement policy based on expected hit count. IEEE Computer Architecture Letters, 17(1):64–67, January/ June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Vandierendonck:2015:EEB

[VHN15] Hans Vandierendonck, Ahmad Hassan, and Dimitrios S. Nikolopoulos. On the energyefficiency of byte-addressable non-volatile memory. IEEE Computer Architecture Letters, 14(2):144–147, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Valero:2016:ELD

[VMP^{+}16] Alejandro Valero, Negar Miralaei, Salvador Petit, Julio Sahuquillo, and Timothy M. Jones. Enhancing the L1 data cache design to mitigate HCI. IEEE Computer Architecture Letters, 15(2):93– 96, July/December 2016. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Verner:2017:EAL

[VMS17] Uri Verner, Avi Mendelson, and Assaf Schuster. Extending Amdahl’s Law for multicores with turbo boost. IEEE Computer Architecture Letters, 16(1):30–33, January/ June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Volos:2021:CRA

[Vol21] Haris Volos. The case for replication-aware memoryerror protection in disaggregated memory. IEEE Computer Architecture Letters, 20 (2):130–133, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Verdu:2016:PSA

[VP16] Javier Verdu and Alex Pajuelo. Performance scalability analysis of JavaScript applications with Web workers. IEEE Computer Architecture Letters, 15(2):105–108, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Vieira:2024:GAP

[VRFT24] João Vieira, Nuno Roma, Gabriel Falcao, and Pedro Tomás. gem5-accel: a pre-RTL simulation toolchain for accelerator architecture validation. IEEE Computer Architecture Letters, 23(1):1–4, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Vijayaraghavan:2018:MBA

[VRS18] Thiruvengadam Vijayaraghavan, Amit Rajesh, and Karthikeyan Sankaralingam. MPU–BWM: Accelerating sequence alignment. IEEE Computer Architecture Letters, 17(2):179–182, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Vandierendonck:2011:FMM

[VS11] Hans Vandierendonck and Andre Seznec. Fairness metrics for multi-threaded processors. IEEE Computer Architecture Letters, 10(1):4–7, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wolff:2014:RUR

[WB14] Sonya R. Wolff and Ronald D. Barnes. Revisiting using the results of pre-executed instructions in runahead processors. IEEE Computer Architecture Letters, 13(2): 97–100, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wingbermuehle:2014:OAS

[WCC14] Joseph G. Wingbermuehle, Ron K. Cytron, and Roger D. Chamberlain. Optimization of application-specific memories. IEEE Computer Architecture Letters, 13(1):45–48, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Walter:2008:BBE

[WCK08] I. Walter, I. Cidon, and A. Kolodny. BENoC: a busenhanced network on-chip for a power efficient CMP. IEEE Computer Architecture Letters, 7(2):61–64, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2009:PST

[WCYC09] Po-Han Wang, Yen-Ming Chen, Chia-Lin Yang, and Yu-Jung Cheng. A predictive shutdown technique for GPU shader processors. IEEE Computer Architecture Letters, 8(1):9–12, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2012:ISA

[WCZ^{+}12] Yaohua Wang, Shuming Chen, Kai Zhang, Jianghua Wan, Xiaowen Chen, Hu Chen, and Haibo Wang. Instruction shuffle: Achieving MIMD-like performance on SIMD architectures. IEEE Computer Architecture Letters, 11(2): 37–40, July/December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2019:MEM

[WJA^{+}19] L. Wang, M. Jahre, A. Adileh, Z. Wang, and L. Eeckhout.

Modeling emerging memorydivergent GPU applications. IEEE Computer Architecture Letters, 18(2):95–98, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2011:GGC

[WJFH11] Xiaoqun Wang, Zhenzhou Ji, Chen Fu, and Mingzeng Hu. GCMS: a global contention management scheme in hardware transactional memory. IEEE Computer Architecture Letters, 10(1):24–27, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wu:2012:CID

[WKE12] Lisa Wu, Martha A. Kim, and Stephen A. Edwards. Cache impacts of datatype acceleration. IEEE Computer Architecture Letters, 11(1):21–24, January/June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wu:2016:MCN

[WL16] Wo-Tak Wu and Ahmed Louri. A methodology for cognitive NoC design. IEEE Computer Architecture Letters, 15(1):1–4, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Weng:2019:DMC

[WLDN19] Jian Weng, Sihao Liu, Vidushi Dadu, and Tony Nowatzki. DAEGEN: a modular compiler for exploring decoupled spatial accelerators. IEEE Computer Architecture Letters, 18(2):161–165, July 2019. ISSN 1556-6064.

Wu:2017:CSB

[WLL17] Hao Wu, Fangfei Liu, and Ruby B. Lee. Cloud server benchmark suite for evaluating new hardware architectures. IEEE Computer Architecture Letters, 16(1):14–17, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2022:CIR

[WLL^{+}22] Yinshen Wang, Wenming Li, Tianyu Liu, Liangjiang Zhou, Bingnan Wang, Zhihua Fan, Xiaochun Ye, Dongrui Fan, and Chibiao Ding. Characterization and implementation of radar system applications on a reconfigurable dataflow architecture. IEEE Computer Architecture Letters, 21 (2):121–124, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2022:ISE

[WLN22] Zhengrong Wang, Christopher Liu, and Tony Nowatzki. Infinity Stream: Enabling transparent and automated in-memory computing. IEEE Computer Architecture Letters, 21(2):85–88, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2019:DDD

[WLWZ19] Yicheng Wang, Yang Liu, Peiyun Wu, and Zhao Zhang. Detect DRAM disturbance error by using disturbance bin counters. IEEE Computer Architecture Letters, 18(1):34– 37, January/June 2019. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wu:2023:RAI

[WLZZ23] Peiyun Wu, Trung Le, Zhichun Zhu, and Zhao Zhang. Redundant array of independent memory devices. IEEE Computer Architecture Letters, 22(2):181–184, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Weston:2023:SLI

[WMJM23] Kevin Weston, Farabi Mahmud, Vahid Janfaza, and Abdullah Muzahid. SmartIndex: Learning to index caches to improve performance. IEEE Computer Architecture Letters, 22(1):33–36, January/ June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2017:DAS

[WMZY17] Rujia Wang, Sparsh Mittal, Youtao Zhang, and Jun Yang. Decongest: Accelerating super-dense PCM under write disturbance by hot page remapping. IEEE Computer Architecture Letters, 16 (2):107–110, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wu:2022:DCG

[WSVS22] Lingxi Wu, Rasool Sharifi, Ashish Venkat, and Kevin Skadron. DRAM-CAM: General-purpose bitserial exact pattern matching. IEEE Computer Architecture Letters, 21(2):89–92, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wu:2021:GOD

[WTSW21] Benjamin Wu, Trishita Tiwari, G. Edward Suh, and Aaron B. Wagner. Guessing outputs of dynamically pruned CNNs using memory access patterns. IEEE Computer Architecture Letters, 20 (2):98–101, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wu:2014:ATE

[Wu14] Carole-Jean Wu. Architectural thermal energy harvesting opportunities for sustainable computing. IEEE Computer Architecture Letters, 13(2):65–68, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2021:WWP

[WXZ^{+}21] Lei Wang, Xingwang Xiong, Jianfeng Zhan, Wanling Gao, Xu Wen, Guoxin Kang, and Fei Tang. WPC: Wholepicture workload characterization across intermediate representation, ISA, and microarchitecture. IEEE Computer Architecture Letters, 20(2):86–89, July/December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2015:PTM

[WYL^{+}15] Zhaoguo Wang, Han Yi, Ran Liu, Mingkai Dong, and Haibo Chen. Persistent transactional memory. IEEE Computer Architecture Letters, 14(1):58–61, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2016:SMF

[WYM^{+}16] Zhenning Wang, Jun Yang, Rami Melhem, Bruce Childers, Youtao Zhang, and Minyi Guo. Simultaneous multikernel: Fine-grained sharing of GPUs. IEEE Computer Architecture Letters, 15 (2):113–116, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Wu:2023:CUD

[WYY^{+}23] Meng Wu, Mingyu Yan, Xiaocheng Yang, Wenming Li, Zhimin Zhang, Xiaochun Ye, and Dongrui Fan. Characterizing and understanding defense methods for GNNs on GPUs. IEEE Computer Architecture Letters, 22 (2):137–140, July/December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wang:2015:LNV

[WZLQ15] Rui Wang, Wangyuan Zhang, Tao Li, and Depei Qian. Leveraging non-volatile storage to achieve versatile cache optimizations. IEEE Computer Architecture Letters, 14(1):46–49, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Xu:2019:PFD

[XCW^{+}19] Sheng Xu, Xiaoming Chen, Ying Wang, Yinhe Han, Xuehai Qian, and Xiaowei Li. PIMSim: a flexible and detailed processing-in-memory simulator. IEEE Computer Architecture Letters, 18(1):6– 9, January/June 2019. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Xie:2022:MSS

[XGH^{+}22] Xinfeng Xie, Peng Gu, Jiayi Huang, Yufei Ding, and Yuan Xie. MPU-Sim: a simulator for in-DRAM near-bank processing architectures. IEEE Computer Architecture Letters, 21(1):1–4, January/June 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Xie:2019:NXB

[XHG^{+}19] Xinfeng Xie, Xing Hu, Peng Gu, Shuangchen Li, Yu Ji, and Yuan Xie. NNBench-X: Benchmarking and understanding neural network workloads for accelerator designs. IEEE Computer Architecture Letters, 18(1):38–42, January/June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Xin:2009:ELI

[XJ09] Jing Xin and Russ Joseph. Exploiting locality to improve circuit-level timing speculation. IEEE Computer Architecture Letters, 8(2):40–43, July/December 2009. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Xiao:2007:NPD

[XL07] X. Xiao and J. Lee. A novel parallel deadlock detection algorithm and hardware for multiprocessor system-on-a-chip. IEEE Computer Architecture Letters, 6(2):41–44, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Xu:2014:STM

[XWG^{+}14] Yunlong Xu, Rui Wang, Nilanjan Goswami, Tao Li, and Depei Qian. Software transactional memory for GPU architectures. IEEE Computer Architecture Letters, 13(1):49– 52, January/June 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Xiao:2016:TAC

[XYMY16] He Xiao, Wen Yueh, Saibal Mukhopadhyay, and Sudhakar Yalamanchili. Thermally adaptive cache access mechanisms for 3D many-core architectures. IEEE Computer Architecture Letters, 15 (2):129–132, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Xiao:2015:SCD

[XYZ15] Canwen Xiao, Yue Yang, and Jianwen Zhu. A sufficient condition for deadlock-free adaptive routing in mesh networks. IEEE Computer Architecture Letters, 14(2):111–114, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yavits:2024:DCD

[Yav24] L. Yavits. DRAMA: Commodity DRAM based content addressable memory. IEEE Computer Architecture Letters, 23(1):65–68, January/ June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

You:2015:QSA

[YC15] Daecheol You and Ki-Seok Chung. Quality of serviceaware dynamic voltage and frequency scaling for embedded GPUs. IEEE Computer Architecture Letters, 14(1): 66–69, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yan:2020:CUG

[YCD^{+}20] Mingyu Yan, Zhaodong Chen, Lei Deng, Xiaochun Ye, Zhimin Zhang, Dongrui Fan, and Yuan Xie. Characterizing and understanding GCNs on GPU. IEEE Computer Architecture Letters, 19(1):22– 25, January/June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Yang:2024:JIJ

[YCH24] Yuxin Yang, Xiaoming Chen, and Yinhe Han. JANM-IK: Jacobian argumented Nelder– Mead algorithm for inverse kinematics and its hardware acceleration. IEEE Computer Architecture Letters, 23(1): 45–48, January/June 2024. ISSN 1556-6056 (print), 1556-6064 (electronic).

Yalcin:2007:UTM

[YE07] G. Yalcin and O. Ergin. Using tag-match comparators for detecting soft errors. IEEE Computer Architecture Letters, 6(2):53–56, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yavits:2018:ASM

[YG18] Leonid Yavits and Ran Ginosar. Accelerator for sparse machine learning. IEEE Computer Architecture Letters, 17(1):21–24, January/ June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yasoubi:2017:PEA

[YHM17] Ali Yasoubi, Reza Hojabr, and Mehdi Modarressi. Power-efficient accelerator design for neural networks using computation reuse. IEEE Computer Architecture Letters, 16(1):72–75, January/ June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yang:2022:SEP

[YHY^{+}22] Ling Yang, Libo Huang, Run Yan, Nong Xiao, Sheng Ma, Li Shen, and Weixia Xu. Stride equality prediction for value speculation. IEEE Computer Architecture Letters, 21(2):57–60, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Yuan:2015:SGR

[YJZ15] Fengkai Yuan, Zhenzhou Ji, and Suxia Zhu. Set-granular regional distributed cooperative caching. IEEE Computer Architecture Letters, 14(1): 75–78, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yavits:2015:RAP

[YKMG15] Leonid Yavits, Shahar Kvatinsky, Amir Morad, and Ran Ginosar. Resistive associative processor. IEEE Computer Architecture Letters, 14 (2):148–151, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yun:2022:GND

[YKP^{+}22] Sungmin Yun, Byeongho Kim, Jaehyun Park, Hwayong Nam, Jung Ho Ahn, and Eojin Lee. GraNDe: Near-data processing architecture with adaptive matrix mapping for graph convolutional networks. IEEE Computer Architecture Letters, 21(2):45–48, July/ December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Yu:2021:MDC

[YLK21] Chao Yu, Sihang Liu, and Samira Khan. MultiPIM: a detailed and configurable multi-stack processing-in-memory simulator. IEEE Computer Architecture Letters, 20(1):54–57, January/ June 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Yasin:2019:TPM

[YMBA19] A. Yasin, A. Mendelson, and Y. Ben-Asher. Tuning performance via metrics with expectations. IEEE Computer Architecture Letters, 18(2):91– 94, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Yavits:2014:CHO

[YMG14] Leonid Yavits, Amir Morad, and Ran Ginosar. Cache hierarchy optimization. IEEE Computer Architecture Letters, 13(2):69–72, July/ December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yoon:2008:CHP

[YNS^{+}08] J. H. Yoon, E. H. Nam, Y. J.

Seong, H. Kim, B. Kim, S. L. Min, and Y. Cho. Chameleon: a high performance flash/ FRAM hybrid solid state disk architecture. IEEE Computer Architecture Letters, 7(1):17– 20, January 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yun:2023:FPP

[YP23] Yugyoung Yun and Eunhyeok Park. Fast performance prediction for efficient distributed DNN training. IEEE Computer Architecture Letters, 22(2):133–136, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Yazdanshenas:2014:CLL

[YPFP14] Sadegh Yazdanshenas, Marzieh Ranjbar Pirbasti, Mahdi Fazeli, and Ahmad Patooghy. Coding last level STT-RAM cache for high endurance and low power. IEEE Computer Architecture Letters, 13(2):73– 76, July/December 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Park:2010:EIP

[yPSS^{+}10] Seon yeong Park, Euiseong Seo, Ji-Yong Shin, Seungryoul Maeng, and Joonwon Lee. Exploiting internal parallelism of flash-based SSDs. IEEE Computer Architecture Letters, 9 (1):9–12, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yoo:2021:MBU

[YSL^{+}21] Mingi Yoo, Jaeyong Song, Jounghoo Lee, Namhyung Kim, Youngsok Kim, and Jinho Lee. Making a better use of caches for GCN accelerators with feature slicing and automatic tile morphing. IEEE Computer Architecture Letters, 20(2):102–105, July/ December 2021. ISSN 1556-6056 (print), 1556-6064 (electronic).

Yavits:2017:RAD

[YWG17] Leonid Yavits, Uri Weiser, and Ran Ginosar. Resistive address decoder. IEEE Computer Architecture Letters, 16 (2):141–144, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yun:2018:RPP

[YYK^{+}18] Ji-Tae Yun, Su-Kyung Yoon, Jeong-Geun Kim, Bernd Burgstaller, and Shin-Dug Kim. Regression prefetcher with preprocessing for DRAM– PCM hybrid main memory. IEEE Computer Architecture Letters, 17(2):163–166, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yan:2022:CUH

[YZY^{+}22] Mingyu Yan, Mo Zou, Xiaocheng Yang, Wenming Li, Xiaochun Ye, Dongrui Fan, and Yuan Xie. Characterizing and understanding HGNNs on GPUs. IEEE Computer Architecture Letters, 21(2): 69–72, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhan:2017:CCS

[ZAK^{+}17] Xin Zhan, Reza Azimi, Svilen Kanev, David Brooks, and Sherief Reda. CARB: a Cstate power management arbiter for latency-critical workloads. IEEE Computer Architecture Letters, 16(1):6–9, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhou:2019:QCD

[ZB19] H. Zhou and G. T. Byrd.

Quantum circuits for dynamic runtime assertions in quantum computation. IEEE Computer Architecture Letters, 18(2):111–114, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhu:2020:HIR

[ZBA^{+}20] Lingjun Zhu, Lennart Bamberg, Anthony Agnesina, Francky Catthoor, Dragomir Milojevic, Manu Komalan, Julien Ryckaert, Alberto Garcia-Ortiz, and Sung Kyu Lim. Heterogeneous 3D integration for a RISC-V system with STT-MRAM. IEEE Computer Architecture Letters, 19(1):51–54, January/ June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhao:2018:KOA

[ZCG18] Wenyi Zhao, Quan Chen, and Minyi Guo. KSM: Online application-level performance slowdown prediction for spatial multitasking GPGPU. IEEE Computer Architecture Letters, 17 (2):187–191, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2006:BIC

[Zha06] Chuanjun Zhang. Balanced instruction cache: reducing conflict misses of directmapped caches through balanced subarray accesses. IEEE Computer Architecture Letters, 5(1):2–5, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhou:2006:CFT

[Zho06] Huiyang Zhou. A case for fault tolerance and performance enhancement using chip multi-processors. IEEE Computer Architecture Letters, 5(1):22–25, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2018:RFA

[ZKF^{+}18] Jiangwei Zhang, Donald Kline, Jr., Long Fang, Rami Melhem, and Alex K. Jones. RETROFIT: Faultaware wear leveling. IEEE Computer Architecture Letters, 17(2):167–170, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2020:FRP

[ZKH^{+}20] Jie Zhang, Miryeong Kwon, Sanghyun Han, Nam Sung Kim, Mahmut Kandemir, and Myoungsoo Jung. FastDrain: Removing page victimization overheads in NVMe storage stack. IEEE Computer Architecture Letters, 19(2):92–96, July/December 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zidenberg:2012:MHS

[ZKW12] Tsahee Zidenberg, Isaac Keslassy, and Uri Weiser. MultiAmdahl: How should I divide my heterogeneous chip? IEEE Computer Architecture Letters, 11(2):65– 68, July/December 2012. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zha:2017:IFM

[ZL17] Yue Zha and Jing Li. IMEC: a fully morphable in-memory computing fabric enabled by resistive crossbar. IEEE Computer Architecture Letters, 16(2):123–126, July/ December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zha:2018:CRC

[ZL18a] Yue Zha and Jing Li. CMA: a reconfigurable complex matching accelerator for wirespeed network intrusion detection. IEEE Computer Architecture Letters, 17(1):33– 36, January/June 2018. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zheng:2018:EPE

[ZL18b] Hao Zheng and Ahmed Louri. EZ-Pass: an energy & performance-efficient powergating router architecture for scalable NoCs. IEEE Computer Architecture Letters, 17(1):88–91, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhao:2017:LIC

[ZLAE17] Xia Zhao, Yuxi Liu, Almutaz Adileh, and Lieven Eeckhout. LA-LLC: Inter-core localityaware last-level cache to exploit many-to-many traffic in GPGPUs. IEEE Computer Architecture Letters, 16(1): 42–45, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2020:AIG

[ZLM^{+}20] Zhihui Zhang, Jingwen Leng, Lingxiao Ma, Youshan Miao, Chao Li, and Minyi Guo. Architectural implications of graph neural networks. IEEE Computer Architecture Letters, 19(1):59–62, January/ June 2020. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2010:FCA

[ZLS10] Meng Zhang, Alvin R.

Lebeck, and Daniel J. Sorin. Fractal consistency: Architecting the memory system to facilitate verification. IEEE Computer Architecture Letters, 9(2):61–64, July/ December 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zebchuk:2007:BBC

[ZM07] J. Zebchuk and A. Moshovos. A building block for coarsegrain optimizations in the onchip memory hierarchy. IEEE Computer Architecture Letters, 6(2):33–36, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2017:WDP

[ZMC17] Dan Zhang, Xiaoyu Ma, and Derek Chiou. Worklistdirected prefetching. IEEE Computer Architecture Letters, 16(2):170–173, July/ December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2023:BPA

[ZNTJE23] Shiqing Zhang, Mahmood Naderan-Tahan, Magnus Jahre, and Lieven Eeckhout. Balancing performance against cost and sustainability in multichip-module GPUs. IEEE Computer Architecture Letters, 22(2):145–148, July/ December 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2018:LHC

[ZS18] Guowei Zhang and Daniel Sanchez. Leveraging hardware caches for memoization. IEEE Computer Architecture Letters, 17(1):59–63, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhu:2014:EWC

[ZSLR14] Yuhao Zhu, Aditya Srikanth, Jingwen Leng, and Vijay Janapa Reddi. Exploiting webpage characteristics for energy-efficient mobile Web browsing. IEEE Computer Architecture Letters, 13(1): 33–36, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhou:2022:LPL

[ZTRA22] Ranyang Zhou, Sepehr Tabrizchi, Arman Roohi, and Shaahin Angizi. LT-PIM: an LUT-Based Processing-in-DRAM architecture with RowHammer self-tracking. IEEE Computer Architecture Letters, 21 (2):141–144, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2016:SIW

[ZTS16] Chulian Zhang, Hamed Tabkhi, and Gunar Schirner. Studying inter-warp divergence aware execution on GPUs. IEEE Computer Architecture Letters, 15(2):117–120, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhang:2003:WHC

[ZVYW03] Chuanjun Zhang, F. Vahid, Jun Yang, and W. Walid. A way-halting cache for lowenergy high-performance systems. IEEE Computer Architecture Letters, 2(1):5, January 2003. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zheng:2015:ACC

[ZWL15] Zhong Zheng, Zhiying Wang, and Mikko Lipasti. Adaptive cache and concurrency allocation on GPGPUs. IEEE Computer Architecture Letters, 14(2):90–93, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhu:2022:RBP

[ZWT22] Yongye Zhu, Shijia Wei, and Mohit Tiwari. Revisiting browser performance benchmarking from an architectural perspective. IEEE Computer Architecture Letters, 21 (2):113–116, July/December 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zokaee:2018:APM

[ZZJ18] Farzaneh Zokaee, Hamid R. Zarandi, and Lei Jiang. AligneR: a process-in-memory architecture for short read alignment in ReRAMs. IEEE Computer Architecture Letters, 17(2):235–238, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zou:2022:AGP

[ZZW^{+}22] Mo Zou, Mingzhe Zhang, Rujia Wang, Xian-He Sun, Xiaochun Ye, Dongrui Fan, and Zhimin Tang. Accelerating graph processing with lightweight learningbased data reordering. IEEE Computer Architecture Letters, 21(1):5–8, January/June 2022. ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhao:2023:RAL

[ZZW^{+}23] Xia Zhao, Guangda Zhang, Lu Wang, Yangmei Li, and Yongjun Zhang. RouteReplies: Alleviating long latency in many-chip-module GPUs. IEEE Computer Architecture Letters, 22(1):29–32, January/June 2023. ISSN 1556-6056 (print), 1556-6064 (electronic).